### v3

# Thresholded Lasso for high dimensional variable selection

Abstract
Given $n$ noisy samples with $p$ dimensions, where $n\ll p$, we show that the multi-step thresholding procedure based on the Lasso – we call it the Thresholded Lasso, can accurately estimate a sparse vector $\beta\in\mathbb{R}^{p}$ in a linear model $Y=X\beta+\epsilon$, where $X_{n\times p}$ is a design matrix normalized to have column $\ell_{2}$-norm $\sqrt{n}$ and $\epsilon\sim N(0,\sigma^{2}I_{n})$. Here $I_{n}$ denotes the identity matrix. We show that under the restricted eigenvalue (RE) condition, it is possible to achieve the $\ell_{2}$ loss within a logarithmic factor of the ideal mean square error one would achieve with an oracle while selecting a sufficiently sparse model – hence achieving sparse oracle inequalities; the oracle would supply perfect information about which coordinates are non-zero and which are above the noise level. We also show for the Gauss-Dantzig selector (Candès-Tao 07), if $X$ obeys a uniform uncertainty principle, one will achieve the sparse oracle inequalities as above, while allowing at most $s_{0}$ irrelevant variables in the model in the worst case, where $0<s_{0}\leq s$ is the smallest integer such that for $\lambda=\sqrt{2\log p/n}$, $\sum_{i=1}^{p}\min(\beta_{i}^{2},\lambda^{2}\sigma^{2})\leq s_{0}\lambda^{2}\sigma^{2}$. Our simulation results on the Thresholded Lasso match our theoretical analysis excellently.
1 Introduction
In a typical high dimensional setting, the number of variables $p$ is much larger than the number of observations $n$. This challenging setting appears in linear regression, signal recovery, covariance selection in graphical modeling, and sparse approximation. In this paper, we consider recovering a vector $\beta\in\mathbb{R}^{p}$ in the following linear model:
$Y=X\beta+\epsilon.$ (1)
Here $X$ is an $n\times p$ design matrix, $Y$ is a vector of noisy observations, and $\epsilon$ is the noise term. We assume throughout this paper that $p\geq n$ (i.e. high-dimensional) and $\epsilon$ is a vector of i.i.d. normal $N(0,\sigma^{2})$ random variables. Denote by $p]=\{1,\ldots,p\}$. The notation $\left\lVert\beta\right\rVert_{2}=(\sum_{j}\beta_{j}^{2})^{1/2}$ stands for the $\ell_{2}$ norm of $\beta$. Given such a linear model, two key tasks are: (1) to select the relevant set of variables and (2) to estimate $\beta$ with bounded $\ell_{2}$ loss. In particular, recovery of the sparsity pattern $S=\mathop{\text{\rm supp}\kern 0.86108pt}(\beta):=\left\{j\,:\,\beta_{j}\neq 0\right\}$, also known as variable (model) selection, refers to the task of correctly identifying the support set, or a subset of “significant” coefficients in $\beta$, based on the noisy observations. Here and in the sequel, we assume that each column vector $X_{j}\in\mathbb{R}^{n},j\in[p]$ of the fixed design matrix $X$ has $\ell_{2}$-length $\left\lVert X_{j}\right\rVert_{2}=\sqrt{n}$.
Even in the noiseless case, recovering $\beta$ (or its support) from $(X,Y)$ seems impossible when $n\ll p$ given that we have more variables than observations. Over the past two decades, a line of research shows that when $\beta$ is sparse, that is, when it has a relatively small number of nonzero coefficients, and when design matrix $X$ behaves sufficiently nicely in a sense that it satisfies certain incoherence conditions, it becomes possible to reconstruct $\beta$ [12, 8, 9].
Throughout this paper, we refer to a vector $\beta\in\mathbb{R}^{p}$ with at most $s$ non-zero entries, where $0\leq s\leq p$, as an $s$-sparse vector. Consider now the linear regression model in (1). For a chosen penalization parameter $\lambda_{n}\geq 0$, regularized estimation with the $\ell_{1}$-norm penalty, also known as the Lasso [34] refers to the following convex optimization problem
$\;\;\hat{\beta}=\arg\min_{\beta\in\mathbb{R}^{p}}\frac{1}{2n}\|Y-X\beta\|_{2}^{2}+\lambda_{n}\|\beta\|_{1},$ (2)
where the scaling factor $1/(2n)$ is chosen by convenience and $\|\beta\|_{1}=\sum_{i=1}^{p}\left\lvert\beta_{i}\right\rvert$. In the present work, we explore model selection beyond focusing on the notion of exact recovery of the support set $\mathop{\text{\rm supp}\kern 0.86108pt}(\beta)$ which crucially depends on the so called $\beta_{\min}$ condition as well as the Neighborhood stability or irrepresentability condition [25, 45, 38, 37]. One can not hope that such incoherence conditions always hold in reality. As pointed out by [36], the irrepresentability condition which is essentially a necessary condition for exact recovery of the non-zero coefficients (for which a $\beta_{\min}$ condition needs to hold) by the Lasso, is much too restrictive in comparison to the Restricted Eigenvalue condition [3]; cf. (3). For some integer $s\in[p]$ and a positive number $k_{0}>0$, we say $\textnormal{{RE}}(s,k_{0},X)$ holds with $K(s,k_{0})$ if for all $\upsilon\not=0$,
$\frac{1}{K(s,k_{0})}\stackrel{{\scriptstyle\triangle}}{{=}}\min_{J\subseteq[p],|J|\leq s}\min_{\left\lVert\upsilon_{J^{c}}\right\rVert_{1}\leq k_{0}\left\lVert\upsilon_{J}\right\rVert_{1}}\;\;\frac{\left\lVert X\upsilon\right\rVert_{2}}{\sqrt{n}\left\lVert\upsilon_{J}\right\rVert_{2}}>0$ (3)
where $\upsilon_{J}$ represents the subvector of $\upsilon\in\mathbb{R}^{p}$ confined to a subset $J$ of $[p]$. It is clear that as $k_{0}$ and $s_{0}$ become smaller, this condition is easier to satisfy. To be clear, RE conditions alone are not sufficient for the Lasso to recover the model $S$ exactly. Moreover, to ensure variable selection consistency, a
$\text{ $\beta_{\min}$ condition:}\quad\min_{j\in S}\left\lvert\beta_{j}\right\rvert\geq C\sigma\sqrt{{2\log p}/{n}},$ (4)
is imposed for some constant $C>1/2$, and shown to be crucial to recover the support of $\beta$ in the information theoretic limit by [37] and [42]. Such a $\beta_{\min}$ condition and the corresponding signal-to-noise ratio (SNR) defined as $\beta_{\min}^{2}/\sigma^{2}$, rather than the typical $\lVert\beta\rVert_{2}^{2}/\sigma^{2}$, is shown to be the key quantity that controls subset selection [37].
Ideally, we aim to remove or relax the $\beta_{\min}$ condition, which is rather unnatural for many applications. Toward this end, we define sparse oracle inequalities as the new criteria for model selection consistency when the irrepresentability condition or related mutual incoherence conditions are replaced with the Restricted Eigenvalue (RE) type of conditions. Roughly speaking, the new criteria ask one to identify a sparse model such that the corresponding least-squares (OLS) estimator based on the selected model achieves an oracle inequality in terms of the $\ell_{2}$ loss while keeping the selection set small. We deem the bound on the $\ell_{2}$-loss as a natural criterion for evaluating a sparse model especially when it is not exactly $S$. We achieve this goal by controlling the false positive selection through thresholding initial estimates of $\beta$ obtained via the Lasso (or the Dantzig selector) at the critical threshold level.
Contributions. Our contributions in this work are twofold. From a methodological point of view, we propose to study the Thresholded Lasso estimator with the aforementioned goals in mind:
  • Step 1
    First, we obtain an initial estimator $\beta_{\text{\rm init}}$ using the Lasso (2) with $\lambda_{n}=d_{0}\sigma\sqrt{2\log p/n}$, for some constant $d_{0}>0$, which is allowed to depend on sparse and restricted eigenvalue parameters;
  • Step 2
    Threshold the initial Lasso estimator $\beta_{\text{\rm init}}$ with $t_{0}$, with the general goal such that, we get a set $I$ with cardinality at most $2s$; in general, we also have $|I\cup S|\leq 2s$, where $I=\left\{j\in[p]:\left\lvert\beta_{j,\text{\rm init}}\right\rvert\geq t_{0}\right\}$ for some $t_{0}\asymp\sigma\sqrt{2\log p/n}$ with hidden constant to be specified;
  • Step 3
    Feed $(Y,X_{I})$ to the ordinary least squares (OLS) estimator to obtain $\hat{\beta}$, where we set $\hat{\beta}_{I}=(X_{I}^{T}X_{I})^{-1}X_{I}^{T}Y$ and the other coordinates to zero.
In Theorem 2.1, we show that the critical threshold level for estimating a high dimensional sparse vector $\beta\in\mathbb{R}^{p}$ should be set at the level $t_{0}\asymp\lambda\sigma$ for $\lambda=\sqrt{2\log p/n}$, to retain signals in $\beta_{\text{\rm init}}$ at or above the level of $\lambda\sigma$, where $\beta_{\text{\rm init}}$ is the solution to the Lasso estimator (2) obtained in Step 1. Moreover, we show that $d_{0}$ and $t_{0}$ are allowed to depend on sparse and restricted eigenvalue parameters of the design matrix; cf. Section 3.1. From a theoretical point of view, the framework for our analysis is set upon the Restricted Eigenvalue type of condition and an upper sparse eigenvalue (USE) condition, namely,
$\text{(USE)}\quad\Lambda_{\max}(2s)\;\stackrel{{\scriptstyle\triangle}}{{=}}\;\max_{\upsilon\not=0;2s-\text{sparse}}\;\left\lVert X\upsilon\right\rVert_{2}^{2}/(n\left\lVert\upsilon\right\rVert_{2}^{2})<\infty.$ (5)
These are among the most general assumptions on the design matrix, guaranteeing sparse recovery and oracle inequalities in the $\ell_{2}$ loss for the Lasso estimator as well as the Dantzig selector: $\textnormal{{RE}}(s,k_{0},X)$ is shown to be a relaxation of the restricted isometry property (RIP) under suitable choices of parameters involved in each condition [9, 3]. Part of this work was presented in a conference paper by [47]. Importantly, we present significant and novel extensions in both theory and numerical simulations, with regards to the Thresholded Lasso and the Lasso under the RE and sparse eigenvalue conditions.
While the crucial theoretical and methodological ideas presented here originate from [47], the current work significantly expands the original ideas and show new results on the sparse oracle inequalities in Theorems 2.1 and 2.4. Compared with the original paper [47], we further study the behavior of the Thresholded Lasso in several challenging situations in Sections 45, and the supplementary Section I. We find that our estimator is robust and adaptive to the overwhelming presence of weak signals in $\mathop{\text{\rm supp}\kern 0.86108pt}(\beta)$ in both theoretical and practical senses. We show that the Thresholded Lasso tradeoffs false positives and false negatives nicely in this case: its advantage in terms of model selection over the Lasso and adaptive Lasso [52, 51] is clearly evident by examining their ROC curves empirically. Our numerical simulations in Section I show that the rates for exact recovery of the support rise sharply for a few types of random matrices once the number of samples passes a certain threshold, using the Thresholded Lasso estimator.
Notation and definitions. Let $T\subseteq[p]$ be a fixed subset of indices. Let $X_{T}$ be the $n\times\left\lvert T\right\rvert$ submatrix obtained by extracting columns of $X$ indexed by $T$. We use $\upsilon_{T}$ to represent the subvector of $\upsilon\in\mathbb{R}^{p}$ confined to a subset $T$ of $[p]$. Let $\lVert\upsilon\rVert_{2}^{2}=\sum_{j=1}^{p}\upsilon_{j}^{2}$. Occasionally, we use $\upsilon_{T}\in\mathbb{R}^{\left\lvert T\right\rvert}$ to also represent its 0-extended version $\upsilon^{\prime}\in\mathbb{R}^{p}$ such that $\upsilon^{\prime}_{T^{c}}=0$ and $\upsilon^{\prime}_{T}=\upsilon_{T}$. We will use $\beta^{\mathop{\natural}\kern 0.60275pt}(T)\in\mathbb{R}^{p}$ to represent the 0-extended version of $\beta_{T}\in\mathbb{R}^{\left\lvert T\right\rvert}$ such that $\beta_{T}^{\mathop{\natural}\kern 0.60275pt}({T})=\beta_{T}$ and $\beta_{T^{c}}^{\mathop{\natural}\kern 0.60275pt}({T})=0$. For a matrix $A$, let $\Lambda_{\min}(A)$ and $\Lambda_{\max}(A)$ denote the smallest and the largest eigenvalues respectively. Let $s=|S|$. We assume the lower sparse eigenvalue (LSE) condition, namely,
$\text{(LSE)}\quad\Lambda_{\min}(2s)\;\stackrel{{\scriptstyle\triangle}}{{=}}\;\min_{\upsilon\not=0;2s-\text{sparse}}\;\;\left\lVert X\upsilon\right\rVert_{2}^{2}/(n\left\lVert\upsilon\right\rVert_{2}^{2})>0,$ (6)
where $n\geq 2s$ is necessary, as any submatrix with more than $n$ columns must be singular. Also relevant is the $(s,s^{\prime})$-restricted orthogonality constant $\theta_{s,s^{\prime}}$ [9], which is defined to be the smallest quantity such that for all disjoint sets $T,T^{\prime}\subseteq[p]$ of cardinality $|T|\leq s$ and $|T^{\prime}|\leq s^{\prime}$:
${\left\lvert{\;\langle{\,X_{T}\upsilon,X_{T^{\prime}}{\upsilon^{\prime}}\,}\rangle\;}\right\rvert}/{n}\leq\theta_{s,s^{\prime}}\left\lVert\upsilon\right\rVert_{2}\left\lVert\upsilon^{\prime}\right\rVert_{2},\;\text{ where }s+s^{\prime}\leq p,$ (7)
$\text{ and }\;\theta_{s,s^{\prime}}\leq(\Lambda_{\max}(s)\Lambda_{\max}(s^{\prime}))^{1/2}\;\text{ by the Cauchy Schwarz inequality.}$ (8)
Note that small values of $\theta_{s,s^{\prime}}$ indicate that disjoint subsets of covariates in $X_{T}$ and $X_{T^{\prime}}$ span nearly orthogonal subspaces. Moreover, we have $\theta_{s,s^{\prime}}\leq{(\Lambda_{\max}(s+s^{\prime})-\Lambda_{\min}(s+s^{\prime}))}/{2}$; cf. Lemma 2.8. Technically speaking, each of the entities defined above, namely, $1/\Lambda_{\min}(2s)$, $\Lambda_{\max}(2s)$, $\theta_{s,s^{\prime}}$, and $K(s,k_{0})$ as introduced in (3), is a non-decreasing function of $s$, $s^{\prime}$, and $k_{0}$. Nonetheless, we crudely consider these as constants following how they are typically treated in the literature as it is to be understood that they grow very slowly with $s$ and $s^{\prime}$; see for example [8, 9][26], and [3]. For two numbers $a,b$, $a\wedge b:=\min(a,b)$, and $a\vee b:=\max(a,b)$. We write $a\asymp b$ if $ca\leq b\leq Ca$ for some positive absolute constants $c,C$ which are independent of $n,p$, and $\gamma$. We write $f=O(h)$ or $f\ll h$ if $\left\lvert f\right\rvert\leq Ch$ for some absolute constant $C<\infty$ and $f=\Omega(h)$ or $f\gg h$ if $h=O(f)$. We write $f=o(h)$ if $f/h\to 0$ as $n\to\infty$.
1.1 Sparse oracle inequalities
In this section, we define sparse oracle inequalities as the new criteria for model selection consistency when some of the signals in $\beta$ are relatively weak, for example, well below the information theoretic detection limit (4) for high dimensional sparse recovery. While the idea of thresholding and refitting is widely used in statistical theory and applications in various contexts, we quantify the threshold level based on the oracle $\ell_{2}$ loss for the Lasso (and Dantzig selector respectively) in the present work, with the following goals.
[…]
Formally, we evaluate the selection set through the following criterion. Consider the least squares estimators $\hat{\beta}_{I}=(X_{I}^{T}X_{I})^{-1}X_{I}^{T}Y$, where $I\subset[p]$ and |I|s|I|\ $\leq$ s. Here and in the sequel, let $\hat{\beta}_{I}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)=\hat{\beta}_{I}$ and $\hat{\beta}_{I^{c}}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)=0$. Consider the ideal least-squares estimator $\beta^{\diamond}$ based on a subset $I$ of size at most $s$, which minimizes the mean squared error:
$\beta^{\diamond}=\operatorname{argmin}_{I\subseteq[p],\;|I|\leq s}{\textbf{E}}\left\lVert\beta-\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)\right\rVert_{2}^{2}.$ (9)
It follows from the analysis by [9] that for $\Lambda_{\max}(s)<\infty$,
${\textbf{E}}\left\lVert\beta-\beta^{\diamond}\right\rVert_{2}^{2}$ $\geq$ $\min\left(1,{1}/{\Lambda_{\max}(s)}\right)\sum_{i=1}^{p}\min(\beta_{i}^{2},\sigma^{2}/n),\text{ where }$ (10)
$\sum_{i=1}^{p}\min(\beta_{i}^{2},\sigma^{2}/n)$ $=$ $\min_{I\subseteq[p]}\left\lVert\beta-\beta^{\mathop{\natural}\kern 0.60275pt}({I})\right\rVert_{2}^{2}+{|I|\sigma^{2}}/{n}$
represents the squared bias and variance. Now we check if (11)
$\lVert\hat{\beta}-\beta\rVert_{2}^{2}$ $=$ $O(\lambda^{2}\sigma^{2}+\sum_{i=1}^{p}\min(\beta_{i}^{2},\lambda^{2}\sigma^{2}))$ (11)
holds with high probability; If so, we claim the following holds:
$\lVert\hat{\beta}-\beta\rVert_{2}^{2}=O_{P}(\log p\max(1,\Lambda_{\max}(s)){\textbf{E}}\left\lVert\beta^{\diamond}-\beta\right\rVert_{2}^{2}),$ (12)
in view of (10); cf. the supplementary Section A. Here the $\ell_{2}$ loss in (11) is optimal up to a $\log p$ factor. We note that (12) is not the tightest upper bound that we could derive due to a relaxation we have on the lower bound as stated in (10). Nevertheless, we use it for its simplicity.
Essential sparsity and objectives. The current paper answers the following question: Is there a good thresholding rule that enables us to obtain a sufficiently sparse estimator $\hat{\beta}$ that satisfies an oracle inequality in the sense of (11), when some components of $\beta_{S}$ are well below $\sigma/\sqrt{n}$? Such oracle results are accomplished without any knowledge of the significant coordinates or parameter values of $\beta$. Both Theorem 2.1 and the supplementary Theorem G.3 answer this question positively, where we elaborate upon the sparse recovery properties of the Lasso and Dantzig selector in combination with thresholding and refitting.
For a given pair of $(n,p)$ values, the essential sparsity parameter $s_{0}$ characterizes more accurately than $s$, the number of significant coefficients of $\beta$ with respect to the noise level $\sigma$ that we should (could) try to recover. Denote by $s_{0}$ the smallest integer such that the following holds [9]:
$\sum_{i=1}^{p}\min(\beta_{i}^{2},\lambda^{2}\sigma^{2})\leq s_{0}\lambda^{2}\sigma^{2},\text{ where }\lambda=\sqrt{2\log p/n}.$ (13)
The parameter $s_{0}$ is relevant especially when we do not wish to impose any lower bound on $\beta_{\min}$. This is the focus of Theorem 2.1. To make this statement precise, we have as a consequence of the definition in (13),
$|\beta_{j}|<\lambda\sigma\;\;\;\text{ for all }j>s_{0},\;\text{ if we order }\;|\beta_{1}|\geq|\beta_{2}|...\geq|\beta_{p}|;$ (14)
cf. Remark 2.2. For simplicity of presentation, we set $|I|<2s_{0}$ as our first goal while achieving the oracle inequality as in (11). One could aim to bound $|I|<cs_{0}$ for some other constant $c>0$. Moreover, to put the bound of $\left|I\right|\leq 2s_{0}$ in perspective, we show in Proposition 4.1 (by setting $c^{\prime}=1$) that the number of variables in $\beta$ that are larger than or equal to $\sigma\sqrt{\log p/n}$ in magnitude is bounded by $2s_{0}$. Roughly speaking, we wish to include most of them by taking $2s_{0}$ as the upper bound on the model size $I$.
The Thresholded Lasso algorithm is constructive in that it relies neither on the unknown parameters $\left\lvert S\right\rvert$ or $\beta_{\min}:=\min_{j\in S}|\beta_{j}|$, nor the exact knowledge of those that characterize the incoherence conditions on $X$. Instead, our choices of penalty $\lambda_{n}$ (Step 1) and threshold $t_{0}$ (Step 2) only depend on $\sigma,n$, and $p$, and some crude estimation of certain sparse eigenvalue parameters; cf. Section 3.1. In practical settings, one can choose $\lambda_{n}$ using cross-validation; See for example the subsequent work by [50], where we use cross-validation to choose both penalty and threshold parameters in the context of covariance selection based on Gaussian graphical models.
In Section 4, we briefly discuss possibilities of recovering a subset of strong signals via thresholding, despite the existence of (or influence from) other relatively weaker signals. Let $T_{0}$ denote the largest $s_{0}$ coordinates of $\beta$ in absolute values. As a consequence of the definition in (13), we have $T_{0}=\{1,\ldots,s_{0}\}$ and $|\beta_{j}|<\lambda\sigma$ for all $j\in T_{0}^{c}$; cf. (14). More precisely, we decompose $T_{0}=\{1,\ldots,s_{0}\}$ into two sets: $A_{0}$ and $T_{0}\setminus A_{0}$, where $A_{0}$ contains the set of coefficients of $\beta$ strictly above $\lambda\sigma$, for which we define a constant $\beta_{\min,A_{0}}$:
$\beta_{\min,A_{0}}:=\min_{j\in A_{0}}|\beta_{j}|>\lambda\sigma$ where $A_{0}=\{j:|\beta_{j}|>\lambda\sigma\}.$ (15)
The goal is to demonstrate the remarkable properties of the Lasso and the Thresholded Lasso estimators: while exact recovery of all non-zero variables requires very stringent incoherence and $\beta_{\min}$ conditions, we can significantly relax both conditions when we only require a subset $A_{0}$ of active variables to be included in our selection set. We loosely refer to $A_{0}$ or its superset $T_{0}\supseteq A_{0}$ which we aim to identify as an active set throughout this work. When $\beta_{\min,A_{0}}$ (15) is sufficiently large, we have $A_{0}\subset I$ while achieving the sparse oracle inequalities in the sense of (11); cf. (4.3) and Theorem 4.4.
One of the reviewers brought to our attention that [43] provide bounds similar to the Dantzig selector under the upper and lower sparse Riesz condition (SRC), which are similar to the upper and lower bounds in (31) in the present work. Both models allow potentially many small coefficients in the true $\beta$. Specifically, we design a new set of experiments to evaluate the impact of sparsity $s$ and the $\beta_{\min,A_{0}}$ (e.g., $C_{a}\lambda\sigma$ in (69)) condition on the recovery of the first $s_{0}$ components in $\beta$, where we set $\beta_{T_{0}^{c}}$ (with support size $s-s_{0}$) to have a fixed $\ell_{2}$ norm but potentially many small coordinates with magnitude $\asymp\sigma/\sqrt{n}$. See Section 5 for the setup of numerical simulations.
Not included in the present work are Theorem 2.1 and the Iterative Procedure by [47], where we show conditions under which one can recover a sparse subset of strong signals when $\beta_{\min}:=\min_{j\in S}\left\lvert\beta_{j}\right\rvert\geq C\sigma\sqrt{{2s\log p}/{n}}$, where $S=\mathop{\text{\rm supp}\kern 0.86108pt}(\beta)$, $s=\left\lvert S\right\rvert$ and $C$ depends on the restricted eigenvalue parameter; cf. Theorems 3.1 in [48]. When $\beta_{\min}$ is sufficiently large, the range of thresholding parameters is even more flexible, which we elaborate in [47] and [48], cf. Theorem 3.1, and hence details are omitted from the current paper. We do show numerical examples for which the Thresholded Lasso recovers the support $S$ exactly with high probability, using a small number of samples per non-zero component in $\beta$, for which the Lasso would certainly have failed, as predicted by the work of [38, 37]. These results have been presented in part in an earlier version of the present paper [48] and the conference paper by [47].
Finally, we define a quantity $\lambda_{\sigma,a,p}$, which bounds the maximum correlation between the noise and covariates of $X$; For each $a\geq 0$, let
${{\mathcal{T}}_{a}}:=\biggl\{\epsilon:\left\lVert{X^{T}\epsilon}/{n}\right\rVert_{\infty}\leq\lambda_{\sigma,a,p},\text{ where }\lambda_{\sigma,a,p}=\sigma\sqrt{1+a}\sqrt{{2\log p}/{n}}\biggr\}.$ (16)
Then, we have ${\mathbb{P}}\left({\mathcal{T}}_{a}\right)\geq 1-(\sqrt{\pi\log p}p^{a})^{-1}$ when $X$ has column $\ell_{2}$ norms bounded by $\sqrt{n}$.
Organization of the paper. The rest of the paper is organized as follows. In Section 2, we describe a thresholding framework for the general setting, and highlight the role thresholding plays in terms of recovering the best subset of variables; we present the main Theorem 2.1 and oracle results for the Lasso estimator. We discuss related work in Section 2.3. Section 3 provides the proof sketch of Theorems 2.1 and 2.4, and the proof of Theorem 2.1 appears in Section 6.1. Section 4 discusses Type II errors and the $\ell_{1}$- and $\ell_{2}$-loss. Section 5 and the supplementary Section I include simulation results. We prove Lemmas 2.52.74.2, and Lemma 4.3 in Sections 6.4 to 6.7. We give a sketch proof of Theorem 3 [43] in the supplementary Section 7. We conclude in Section 8. Additional technical proofs are included in the supplement.
2 The Thresholded Lasso estimator
Theorem 2.1 states that sparse oracle inequalities as elaborated in Section 1.1 hold for the Thresholded Lasso under no restriction on $\beta_{\min}$. We do not optimize constants in this paper. Theorem 2.1 is the key contribution of this paper and is proved in Section 6.1.
Theorem 2.1.
(Ideal model selection for the Thresholded Lasso) Suppose $\beta\in\mathbb{R}^{p}$ is $s$-sparse. Let $s_{0}$ be as in (13). Let $Y=X\beta+\epsilon$, where $\epsilon=(\epsilon_{1},\ldots,\epsilon_{n})^{T}$ is a vector containing independent and identically distributed (i.i.d.) noise with $\epsilon_{i}\sim N(0,\sigma^{2})$ for all $i\in[n]$. Suppose the columns of $X$ are normalized to have $\ell_{2}$ norm $\sqrt{n}$. Suppose $\textnormal{{RE}}(s_{0},4,X)$ holds with $K(s_{0},4)$, and the sparse eigenvalue conditions (5) and (6) hold. Let $\beta_{\text{\rm init}}$ be an optimal solution to the Lasso (2) with $\lambda_{n}=d_{0}\sqrt{2\log p/n}\sigma$ $\geq 2\lambda_{\sigma,a,p}$, where $d_{0}\geq 2\sqrt{1+a}$ for $a\geq 0$. Set $t_{0}=C_{4}\lambda\sigma\geq 2\sqrt{1+a}\lambda\sigma$, for some constant $C_{4}\geq D_{1}$ for $D_{1}$ as in (27). Set $I=\left\{j\in[p]:\left\lvert\beta_{j,\text{\rm init}}\right\rvert\geq t_{0}\right\}$. Set $\hat{\beta}_{I}=(X_{I}^{T}X_{I})^{-1}X_{I}^{T}Y$ and $\hat{\beta}_{I^{c}}=0$. Then with probability at least $1-{\mathbb{P}}\left({\mathcal{T}}_{a}^{c}\right)$, we obtain
$|I|\leq s_{0}(1+D_{1}/C_{4})<2s_{0},\quad|I\cup S|\leq s+s_{0}\;\text{ and }$
[…]
$D_{4}^{2}\leq((D^{\prime}_{0}+C_{4})^{2}+1)\left(\frac{3}{2}+\frac{(\Lambda_{\max}(2s)-\Lambda_{\min}(2s))^{2}}{2\Lambda_{\min}^{2}(2s_{0})}\right).$ (18)
Theorem 2.1 relies on new oracle results that we prove for the Lasso estimator under $\textnormal{{RE}}(s_{0},4,X)$ and the upper sparse eigenvalue condition (5) in Theorem 2.4. The Lasso estimator achieves essentially the same bound in terms of $\ell_{2}$ loss as stated in (11), which adapts nearly ideally not only to the uncertainty in the support set $S$ but also to the size of the “significant” set; cf. Section 3.1. It is clear from our analysis in [47] that, showing an oracle inequality as in (11) for the initial Lasso estimator $\beta_{\text{\rm init}}$, cf. Theorem 2.4, as well as applying new techniques for analyzing algorithms involving thresholding followed by OLS refitting will be crucial in proving sparse oracle inequalities for the Thresholded Lasso. We discuss the initial estimator $\beta_{\text{\rm init}}$ and the set $I$ in Section 2.1. Moreover, (17) implies that (11) holds for the Thresholded Lasso estimator.
Remark 2.2.
To see (14), we have by definition of $s_{0}$, where $0\leq s_{0}\leq s<p$,
$s_{0}\lambda^{2}\sigma^{2}$ $\leq$ $\lambda^{2}\sigma^{2}+\sum_{i=1}^{p}\min(\beta_{i}^{2},\lambda^{2}\sigma^{2})$
$\leq$ $2\log p\big({\sigma^{2}}/{n}+\sum_{i=1}^{p}\min\big(\beta_{i}^{2},{\sigma^{2}}/{n}\big)\big)$
$\text{ and}\;\;s_{0}\lambda^{2}\sigma^{2}$ $\geq$ $\sum_{j=1}^{s_{0}+1}\min(\beta_{j}^{2},\lambda^{2}\sigma^{2})\geq(s_{0}+1)\min(\beta_{s_{0}+1}^{2},\lambda^{2}\sigma^{2}),$
which immediately implies that $\min(\beta_{s_{0}+1}^{2},\lambda^{2}\sigma^{2})<\lambda^{2}\sigma^{2}$ and hence (14) holds.
2.1 The thresholding rules
Consider the linear regression model in (1). Let $T_{0}$ denote the largest $s_{0}$ coordinates of $\beta$ in absolute values, for $s_{0}$ as in (13). Suppose we aim to select the set of variables of size at least $\sigma\sqrt{2\log p/n}$ while controlling the bias. Lemma 2.3 is a deterministic result toward this goal. We prove Lemma 2.3 in Section 6.2.
Lemma 2.3.
(A deterministic result.) Let $\beta_{\text{\rm init}}$ be an initial estimator of the $s$-sparse $\beta$ in the linear model (1), where $\epsilon\sim N_{n}(0,\sigma^{2}I)$ and $\left\lVert X_{j}\right\rVert_{2}=\sqrt{n},j\in[p]$. Let $T_{0}$ denote the largest $s_{0}$ coordinates of $\beta$ in absolute values. Let $\beta_{T_{0}}$ be the restriction of $\beta$ to the set $T_{0}$. Let $\beta^{\mathop{\natural}\kern 0.60275pt}({T_{0}})\in\mathbb{R}^{p}$ denote its 0-extended version such that $\beta_{T_{0}^{c}}^{\mathop{\natural}\kern 0.60275pt}({T_{0}})=0$ and $\beta_{T_{0}}^{\mathop{\natural}\kern 0.60275pt}({T_{0}})=\beta_{T_{0}}$. Let $h=\beta_{\text{\rm init}}-\beta^{\mathop{\natural}\kern 0.60275pt}({T_{0}})$. Suppose for $\lambda:=\sqrt{{2\log p}/{n}}$,
$\left\lVert h_{T_{0}}\right\rVert_{2}$ $\leq$ $D^{\prime}_{0}\lambda\sigma\sqrt{s_{0}}\;\text{ and }\;\left\lVert h_{T_{0}^{c}}\right\rVert_{1}\leq D_{1}\lambda\sigma s_{0}.$ (19)
Set $t_{0}=C_{4}\lambda\sigma$ for some positive constant $C_{4}$. Let $I=\{j:\left\lvert\beta_{j,\text{\rm init}}\right\rvert\geq t_{0}\}$ and ${\mathcal{D}}:=[p]\setminus I$. Then the set $I$ satisfies
$|I|$ $\leq$ $s_{0}(1+D_{1}/C_{4}),\quad|I\cup S|\leq s+s_{0}D_{1}/C_{4},\quad\text{ and }$ (20)
[…]
Then, a tighter bound on $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}:=\left\lVert\beta_{\text{\rm init},T_{0}^{c}}\right\rVert_{1}$ or $\lVert\beta_{\text{\rm init},T_{0}^{c}}\rVert_{2}$ will decrease the threshold $t_{0}$ while a tighter bound on $\lVert h_{T_{0}}\rVert_{2}:=\lVert(\beta-\beta_{\text{\rm init}})_{T_{0}}\rVert_{2}\leq D_{0}^{\prime}\lambda\sigma\sqrt{s_{0}}$ will tighten the upper bound in (21) on the bias component through the triangle inequality. In general, we allow $t_{0}$ to be chosen from a reasonably wide range, where we tradeoff the width of the range with the tightness of the upper bound on the $\ell_{2}$ loss for $\hat{\beta}$ in Step 3. This saves us from having to estimate incoherence parameters in a refined (and tedious) manner. Theorem 2.4 may be of independent interest. We give a proof sketch in Section 3.1.
Theorem 2.4.
(Oracle inequalities of the Lasso) Suppose $\beta\in\mathbb{R}^{p}$ is $s$-sparse. Let $s_{0}$ be as in (13). Suppose $\textnormal{{RE}}(s_{0},4,X)$ and the upper sparse eigenvalue condition (5) hold. Under the settings of Lemma 2.3, let $\beta_{\text{\rm init}}$ be an optimal solution to the Lasso (2) with $\lambda_{n}=d_{0}\lambda\sigma\geq 2\lambda_{\sigma,a,p}$, where $a\geq 0$ and $d_{0}\geq 2\sqrt{1+a}$. Let $h=\beta_{\text{\rm init}}-\beta^{\mathop{\natural}\kern 0.60275pt}({T_{0}})$ be defined as in Lemma 2.3. Then on ${\mathcal{T}}_{a}$,
$\left\lVert\beta_{\text{\rm init}}-\beta\right\rVert_{2}$ $\leq$ $\lambda\sigma\sqrt{s_{0}}(\sqrt{D_{0}^{2}+D_{1}^{2}}+1),$
$\left\lVert h_{T_{0}}\right\rVert_{1}+\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ $=$ $\left\lVert\beta_{T_{0}}-\beta_{\text{\rm init},T_{0}}\right\rVert_{1}+\left\lVert\beta_{\text{\rm init},T_{0}^{c}}\right\rVert_{1}\leq D_{2}\lambda\sigma s_{0},\quad\text{and}$
$\left\lVert X\beta_{\text{\rm init}}-X\beta\right\rVert_{2}/\sqrt{n}$ $\leq$ $D_{3}\lambda\sigma\sqrt{s_{0}},$
where $D_{0},\ldots,D_{3}$ are defined in (23) to (29). Moreover, for any subset $I_{0}\subset S$, by assuming that $\textnormal{{RE}}(|I_{0}|,4,X)$ holds with $K(|I_{0}|,4)$, we have
$\left\lVert X\beta_{\text{\rm init}}-X\beta\right\rVert_{2}/\sqrt{n}\leq\left\lVert X\beta-X_{I_{0}}\beta_{I_{0}}\right\rVert_{2}/\sqrt{n}+\frac{3}{2}K(|I_{0}|,4)\lambda_{n}\sqrt{|I_{0}|}.$ (22)
We compare Theorem 2.4 with the $\ell_{p}$ error bounds by [3] (cf. Theorem 7.2) in Section 3.1. The sparse oracle properties of the Thresholded Lasso in terms of variable selection, $\ell_{2}$ loss, and prediction error then follow from Theorem 2.4, Lemmas 2.3 and 2.7; cf. Section 6.1. The full proof of Theorem 2.4 appears in the supplementary Section D, which yields the following: cf. (19),
$D_{0}$ $=$ $D\vee\sqrt{2}\big(d_{0}K^{2}(s_{0},4)+K(s_{0},3)\sqrt{\Lambda_{\max}(s-s_{0})}+2d_{0}K^{2}(s_{0},3)\big),$ (23)
$D^{\prime}_{0}$ $=$ $D\vee(d_{0}K^{2}(s_{0},4))\vee(K(s_{0},3)\sqrt{\Lambda_{\max}(s-s_{0})}+3d_{0}K^{2}(s_{0},3)),$ (26)
$\text{ where }\;D=\sqrt{\frac{\Lambda_{\max}(s-s_{0})}{\Lambda_{\min}(2s_{0})}}\big(1+\frac{3\ell(s_{0})\sqrt{\Lambda_{\max}(s-s_{0})}}{d_{0}}\big)\;\text{ and }$
$\ell(s_{0})=({\theta_{s_{0},2s_{0}}}/{\sqrt{\Lambda_{\min}(2s_{0})}})\wedge\sqrt{\Lambda_{\max}(s_{0})},$
$D_{1}$ $=$ $d_{0}[\big(\frac{\Lambda_{\max}(s-s_{0})}{d^{2}_{0}}+\frac{9}{4}K^{2}(s_{0},3)\big)\vee 4K(s_{0},4)^{2}\vee\frac{3\Lambda_{\max}(s-s_{0})}{d^{2}_{0}}],$ (27)
$D_{2}$ $=$ $d_{0}[\big(\frac{\Lambda_{\max}(s-s_{0})}{d^{2}_{0}}+4K^{2}(s_{0},3)\big)\vee 5K(s_{0},4)^{2}\vee\frac{4\Lambda_{\max}(s-s_{0})}{d^{2}_{0}}],$ (29)
$\text{ and }\quad D_{3}=\sqrt{\Lambda_{\max}(s-s_{0})}+d_{0}K(s_{0},4)/2+d_{0}K(s_{0},3).$
Hence
$D^{\prime}_{2}$ $=$ $(\frac{1}{d_{0}}\Lambda_{\max}(s-s_{0})+4K(s_{0},3)^{2}d_{0})\vee(\frac{4}{d_{0}}\Lambda_{\max}(s-s_{0}))\vee 5d_{0}K(s_{0},4)^{2}$
$\leq$ $5\big(\Lambda_{\max}(s-s_{0})/d_{0}\vee(d_{0}K(s_{0},4)^{2})\big).$
The bounds in (30) and (36) ensure that $\theta_{s_{0},2s_{0}}<\infty$ under (5) and $1/\Lambda_{\min}(2s_{0})<2K^{2}(s_{0},1)\leq 2K^{2}(s_{0},4)<\infty$ under $\textnormal{{RE}}(s_{0},4,X)$. In fact, we can obtain an upper bound on $\theta_{s_{0},2s_{0}}$ in two ways. Let disjoint sets $J,J^{\prime}\subset[p]$ satisfy $\left\lvert J\right\rvert\leq s_{0}$ and $\left\lvert J^{\prime}\right\rvert\leq 2s_{0}$ and vectors $\upsilon,\upsilon^{\prime}$ satisfy $\left\lVert\upsilon\right\rVert_{2}=\left\lVert\upsilon^{\prime}\right\rVert_{2}=1$. We have by the Cauchy–Schwarz inequality and the parallelogram identity:
${\left\lvert{\;\langle{\,X_{J}\upsilon,X_{J^{\prime}}{\upsilon^{\prime}}\,}\rangle\;}\right\rvert}\leq\left\lVert X_{J}\upsilon\right\rVert_{2}\left\lVert X_{J^{\prime}}{\upsilon^{\prime}}\right\rVert_{2}\leq n\sqrt{\Lambda_{\max}(s_{0})\Lambda_{\max}(2s_{0})}\leq n\Lambda_{\max}(2s_{0})$
and moreover, cf. the proof of Lemma 2.8,
${\left\lvert{\;\langle{\,X_{J}\upsilon,X_{J^{\prime}}{\upsilon^{\prime}}\,}\rangle\;}\right\rvert}/{n}\leq(\Lambda_{\max}(3s_{0})-\Lambda_{\min}(3s_{0}))/2,\text{ and hence},$
$\theta_{s_{0},2s_{0}}\leq\sqrt{\Lambda_{\max}(s_{0})\Lambda_{\max}(2s_{0})}\wedge\big(\Lambda_{\max}(3s_{0})-\Lambda_{\min}(3s_{0})\big)/2<\infty.$ (30)
To help build intuition, we first state in Lemma 2.5 a general result on the $\ell_{2}$ loss for the OLS estimator, when a subset of relevant variables is missing from the fixed model $I$. Lemma 2.5 is also an important technical contribution of this paper, which may be of independent interest. The assumption on $I$ being fixed is then relaxed in Lemma 2.7. Given $\hat{\beta}_{I}=(X_{I}^{T}X_{I})^{-1}X_{I}^{T}Y$, let $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)$ be the 0-extended version of $\hat{\beta}_{I}$ such that $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}_{I^{c}}=0$ and $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}_{I}=\hat{\beta}_{I}$ in both lemmas. Let $c_{0},c>0$ be some absolute constants.
Lemma 2.5.
(OLS estimator with missing variables) Suppose sparse eigenvalue conditions (5) and (6) hold. Given a deterministic set $I\subset[p]$, set ${\mathcal{D}}:=[p]\setminus I$ and $S_{\mathcal{D}}={\mathcal{D}}\cap S=S\setminus I$. Let $\left|I\right|=m\leq(c_{0}s_{0})\wedge s$. Suppose $|I\cup S|\leq 2s$. Let $\hat{\beta}_{I}=(X_{I}^{T}X_{I})^{-1}X_{I}^{T}Y$. Then, with probability at least $1-2\exp(-3m/64)$,
$\left\lVert\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)-\beta\right\rVert_{2}^{2}\leq\big(\frac{2\theta^{2}_{|I|,|{S_{{\mathcal{D}}}}|}}{\Lambda_{\min}^{2}(m)}+1\big)\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}^{2}+\frac{3m\sigma^{2}}{n\Lambda_{\min}(m)}.$
Remark 2.6.
Lemmas 2.5 and 2.7 apply to $X$ so long as sparse eigenvalue conditions (5) and (6) hold. As a consequence, for any subset $I$ such that $|I|\leq 2s$,
$\infty>\Lambda_{\max}(2s)$ $\geq$ $\Lambda_{\max}(|I|)\geq\Lambda_{\max}\left({X_{I}^{T}X_{I}}/{n}\right)$ (31)
$\geq$ $\Lambda_{\min}\left({X_{I}^{T}X_{I}}/{n}\right)\geq\Lambda_{\min}(|I|)\geq\Lambda_{\min}(2s)>0.$
For disjoint sets $I$ and $S_{{\mathcal{D}}}=S\setminus I$ as in Lemmas 2.5 and 2.7, we have
$|I|+|S_{{\mathcal{D}}}|=|I\cup S|\leq 2s,\;\text{and}\;\theta_{|I|,|S_{{\mathcal{D}}}|}\leq{(\Lambda_{\max}(2s)-\Lambda_{\min}(2s))}/{2},$ (32)
where it is understood that $\Lambda_{\min}(2s)=0$ is also permitted; cf. Lemma 2.8 and its proof in the supplementary Section B.
Lemma 2.5 implies that even if we miss some columns of $X$ in $S$, we can still hope to get the $\ell_{2}$ loss bounded as in Theorem 2.1 so long as $\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}$ is bounded by $O_{P}(\lambda\sigma\sqrt{s_{0}})$ while $\left\lvert I\right\rvert$ is sufficiently small. Both conditions are guaranteed to hold by our choices of the thresholding parameters as shown in Lemma 2.3. Although the tight analysis of Lemma 2.5 depends on the fact that the selection set $I$ is deterministic, a simple variation of the statement makes it work well with the thresholded estimators as considered in the present paper, with $\ell_{2}$ error bounded essentially at the same order of magnitude as in (11) so long as $\left|I\right|=O(s_{0})$ and $|I\cup S|\leq 2s$. Lemma 2.7 is presented by [47].
Lemma 2.7.
[47] Suppose sparse eigenvalue conditions (5) and (6) hold. Given an arbitrary set $I\subset[p]$, possibly random, set ${\mathcal{D}}:=[p]\setminus I$ and $S_{\mathcal{D}}={\mathcal{D}}\cap S$. Suppose on event ${\mathcal{T}}_{a}$, $\left|I\right|\leq(c_{0}s_{0})\wedge s$ and $|I\cup S|\leq 2s$. Then, for $\hat{\beta}_{I}=(X_{I}^{T}X_{I})^{-1}X_{I}^{T}Y$ and $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)$ as defined in Lemma 2.5, it holds on ${\mathcal{T}}_{a}$ that
$\left\lVert\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)-\beta\right\rVert_{2}^{2}$ $\leq$ $\big(\frac{2\theta^{2}_{|I|,|{S_{{\mathcal{D}}}}|}}{\Lambda^{2}_{\min}(|I|)}+1\big)\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}^{2}+\frac{2|I|(1+a)\sigma^{2}\lambda^{2}}{\Lambda^{2}_{\min}(|I|)}.$ (33)
The proofs for Lemmas 2.5 and 2.7 are deferred to Sections 6.4 and 6.5. We now state Lemma 2.8, which follows from [8].
Lemma 2.8.
[Lemma 1.2 [8]] Suppose sparse eigenvalue conditions (5) and (6) hold. The following statements hold: (a) For all disjoint sets $I,S_{{\mathcal{D}}}\subseteq[p]$ of cardinality $|S_{{\mathcal{D}}}|<s$ and $|I|+|S_{{\mathcal{D}}}|\leq 2s$,
$\theta_{|I|,|S_{{\mathcal{D}}}|}\leq(\Lambda_{\max}(2s)-\Lambda_{\min}(2s))/{2};$
(b) Without the lower sparse eigenvalue (LSE) condition (6), we still obtain
$\theta_{|I|,|S_{{\mathcal{D}}}|}\leq\Lambda_{\max}(s)\wedge(\Lambda_{\max}(2s)/2)<\infty\;\text{ if }\;\left\lvert I\right\rvert\vee|S_{{\mathcal{D}}}|\leq s,$ (34)
under the USE condition (5), following the same arguments leading to (30).
Remark 2.9.
(I.) In Lemma 2.7, we have $\left\lvert I\right\rvert\vee|S_{{\mathcal{D}}}|\leq s$. Hence, the LSE condition (6) can be replaced with the following relaxed lower sparse eigenvalue condition:
$\Lambda_{\min}((cs_{0})\wedge s)>0,\text{ in case }\;|I|\leq(cs_{0})\wedge s\text{ for some $c\geq 2$}$ (35)
so that (33) holds on ${\mathcal{T}}_{a}$, with $\theta_{|I|,|{S_{{\mathcal{D}}}}|}\leq\Lambda_{\max}(s)\wedge(\Lambda_{\max}(2s)/2)$ as in (34).
(II.) We note that if $\textnormal{{RE}}(s_{0},k_{0},X)$ as defined in (3) is satisfied with $k_{0}\geq 1$ and $1\leq s_{0}<p$ then $1/{\Lambda_{\min}(2s_{0})}\leq 2K^{2}(s_{0},1)$. Consider $2s_{0}$ sparse vector $\upsilon$. Let $T_{0}$ denote the locations of the $s_{0}$ largest coefficients of $\upsilon$ in absolute values. Then $\left\lVert\upsilon\right\rVert_{2}^{2}\leq 2\left\lVert\upsilon_{T_{0}}\right\rVert_{2}^{2}$, since $\left\lVert\upsilon_{T_{0}^{c}}\right\rVert_{2}\leq\left\lVert\upsilon_{T_{0}}\right\rVert_{2}.$ Thus we have for any $2s_{0}$-sparse vector $\upsilon$, by $\textnormal{{RE}}(s_{0},4,X)$ (3),
$\frac{\left\lVert X\upsilon\right\rVert_{2}^{2}}{n\left\lVert\upsilon\right\rVert_{2}^{2}}\geq\frac{\left\lVert X\upsilon\right\rVert_{2}^{2}}{2n\left\lVert\upsilon_{T_{0}}\right\rVert_{2}^{2}}\geq 1/{(2K^{2}(s_{0},1))},\text{ where }\left\lVert\upsilon_{T_{0}^{c}}\right\rVert_{1}\leq\left\lVert\upsilon_{T_{0}}\right\rVert_{1}.$
Hence (35) holds for $c=2$, when $2s_{0}<s$, since
$\Lambda_{\min}(2s_{0})\;\stackrel{{\scriptstyle\triangle}}{{=}}\;\min_{\upsilon\not=0;2s_{0}-\text{sparse}}\;{\left\lVert X\upsilon\right\rVert_{2}^{2}}/{(n\left\lVert\upsilon\right\rVert_{2}^{2})}\geq 1/(2K^{2}(s_{0},1)).$ (36)
Remark 2.10.
One notion of the incoherence condition which has been formulated in the sparse reconstruction literature bears the name of restricted isometry property (RIP) [8, 9]. For each integer $s=1,2,\ldots$ such that $s<p$, the $s$-restricted isometry constant $\delta_{s}$ of $X$ is defined to be the smallest quantity such that
$(1-\delta_{s})\left\lVert\upsilon\right\rVert_{2}^{2}\leq\left\lVert X_{T}\upsilon\right\rVert_{2}^{2}/n\leq(1+\delta_{s})\left\lVert\upsilon\right\rVert_{2}^{2}$ (37)
for all $T\subset[p]$ with $|T|\leq s$ and coefficients sequences $(\upsilon_{j})_{j\in T}$ [8]; Hence the upper and lower $s$-sparse eigenvalues of design matrix $X$ satisfy $1+\delta_{s}\geq\Lambda_{\max}(s)\geq\Lambda_{\min}(s)>1-\delta_{s}$.
2.2 Discussions
The tight analysis in Theorem 2.1 for the Thresholded Lasso estimator is motivated by the sparse oracle inequalities on the Gauss-Dantzig selector under the UUP, which is originally shown to hold in a conference paper by [47]. Previously, it has been shown that (11) holds with high probability for the Dantzig selector under the condition of a Uniform Uncertainty Principle (UUP), where the UUP states that for all $s$-sparse sets $J$, the columns of $X$ corresponding to $J$ are almost orthogonal [9].
section 2.11.
(A Uniform Uncertainly Principle) For some integer $1\leq s<n/3$, assume $\delta_{2s}+\theta_{s,2s}<1-\tau$ for some $\tau>0$.
Then under the settings of Lemma 2.7, we have $\theta_{|I|,|{S_{{\mathcal{D}}}}|}\leq\delta_{2s}<1$, where $|I|+|S_{{\mathcal{D}}}|\leq 2s$; cf. (32). The sparse eigenvalue conditions in the present work are considerably relaxed from the incoherence condition (UUP) in Definition G.1: The UUP condition ensures that the $\textnormal{{RE}}(s,1,X)$ condition as in (3) holds with
$K(s,1)={\sqrt{\Lambda_{\min}(2s)}}/{(\Lambda_{\min}(2s)-\theta_{s,2s})}\leq{\sqrt{\Lambda_{\min}(2s)}}/{\tau};$
cf. [3]. Besides the sparse eigenvalue conditions, Theorem 2.1 requires $\textnormal{{RE}}(s_{0},4,X)$ to be satisfied, which depends on the essential sparsity $s_{0}$ rather than $s=\left\lvert\mathop{\text{\rm supp}\kern 0.86108pt}(\beta)\right\rvert$; cf. Section 3 for detailed discussions. In the supplementary Theorem G.3, we show the corresponding result for the Gauss-Dantzig selector under the UUP for completeness.
In summary, Lemmas 2.3 and 2.7 ensure that the general thresholding rules with threshold level at about $\lambda\sigma$ achieve the following property: although we cannot guarantee the presence of variables indexed by $S_{R}=\{j:|\beta_{j}|<\sigma\sqrt{\log p/n}\}$ to be included due to their lack of strength, we will include in $I$ most variables in $S\setminus S_{R}$ such that the OLS estimator based on model $I$ achieves the oracle bound (11). This goal is accomplished despite some variables from the support set $S$ are missing from the model $I$, since their overall $\ell_{2}$-norm $\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}$ is bounded in (21).
As mentioned, Proposition 4.1 (by setting $c^{\prime}=1$) shows that the number of variables in $\beta$ that are larger than or equal to $\sqrt{\log p/n}\sigma$ in magnitude is bounded by $2s_{0}$. In hindsight, it is clear that we wish to retain most of them by keeping $2s_{0}$ variables in the model $I$. Indeed, suppose $D_{1}s_{0}<(c_{0}s_{0})\wedge(s-s_{0})$. Then, by choosing $t_{0}\asymp\lambda\sigma$ on event ${\mathcal{T}}_{a}$, we are guaranteed to obtain under the settings of Theorem 2.1 (and Lemma 2.3),
$|I|$ $\leq$ $s_{0}(1+D_{1})<s,\;\;|I\cup S|\leq 2s\;\text{ and }\lVert\hat{\beta}-\beta\rVert_{2}^{2}\leq D_{4}^{2}s_{0}\lambda^{2}\sigma^{2}.$
Here, we assume that $D_{1}$ as in (27) will grow only mildly with the parameters $s_{0},s$, under conditions (3) and (5), and it is not necessary to set $C_{4}>D_{1}$. The set of missing variables in $\mathcal{D}$} is the price we pay in order to obtain a sparse model when some coordinates in the support $\mathop{\text{\rm supp}\kern 0.86108pt}(\beta)$ are well below $\sigma\sqrt{\log p/n}$. Note that when we allow the model size to increase by lowering $t_{0}$, the variance term $\propto|I|\lambda^{2}/\Lambda_{\min}(|I|)$ becomes correspondingly larger. Since the larger model $I$ may not include more true variables, the size of ${S_{{\mathcal{D}}}}=S\setminus I$ may remain invariant; If so, the overall interaction term $\theta\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}/\Lambda_{\min}(|I|)$ can still increase due to the increased orthogonality coefficient $\theta:=\theta_{|I|,|S_{{\mathcal{D}}}|}$, even though $\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}$ is a non-increasing function of the model size.
This argument favors the selection of a small (yet sufficient) model as stated in Theorem 2.1, rather than blindly including extraneous variables in the set $I$. We mention in passing that by setting an upper bound on the desired model size $\left\lvert I\right\rvert\leq 2s_{0}$, we are able to make some interesting connections between the thresholded estimators as studied in the present paper and the $\ell_{0}$ penalized least squares estimators. In particular, we show that the prediction error, $\lVert X\beta_{\text{\rm init}}-X\beta\rVert_{2}$, and a complexity-based penalty term $\sigma\sqrt{\left\lvert I\right\rvert\log p}$ on the chosen model $I$ are both bounded by $O_{P}(\sigma\sqrt{s_{0}\log p})$ in case $|I|\asymp 2s_{0}$ for the thresholded estimators by [48].
2.3 Background and related work
In this section, we briefly discuss related work. Lasso and the Dantzig selector are both computational efficient and shown with provable nice statistical properties; see for example [25, 17, 38, 45, 9, 6, 7, 22, 21, 43, 26, 3]. We refer to the books for a comprehensive survey of related results [5, 39].
Prior to our work, a similar two-step procedure, namely, the Gauss-Dantzig selector, has been proposed and empirically studied by [9]. This paper builds upon the methodology originally developed in a conference paper by the present author [47]. Inspired by [9][47] obtains oracle bounds in the spirit of Theorem 2.1 for the Gauss-Dantzig selector, under the stronger restricted isometry type of conditions as originally proposed by [9]; cf. the supplementary Theorem G.3.
Under variants of the RIP conditions, the exact recovery or approximate reconstruction of a sparse $\beta$ using the basis pursuit program [10] has been shown in a series of results [12, 8, 9, 30]. We refer to the book by [39] and the paper by [35] for a complete exposition. The sparse recovery problem under arbitrary noise is also well studied, see for example [28] and [27], where they require $s$ to be part of the input. See [16],[14][50][18] for further references and applications of the essential sparsity.
For the Lasso, [26] has also shown in theoretical analysis that thresholding is effective in obtaining a two-step estimator $\hat{\beta}$ that is consistent in its support with $\beta$ when $\beta_{\min}$ is sufficiently large. A weakening of the incoherence condition by [26] is still sufficient for (3) to hold [3]. See also [24][53] and [44]. A more general framework on multi-step variable selection was explored by [41]. They control the probability of false positives at the price of false negatives, similar to what we aim for here; their analysis is constrained to the case when $s$ is a constant.
Subsequent development. This choice of the threshold parameter identified in [47, 48] and the current paper has deep connection with the classic and current literature on model selection [16, 14, 41, 42, 39].  [42] proves the minimax concave penalty (MCP) procedure is selection consistent under a sparse Riesz condition and an information requirement in the sense of (4). Nonconvexity of the minimization problem causes computational and analytical difficulties; Compared to the elegant yet complex method in MCP, the Thresholded Lasso procedures [47, 48] provide a much simpler framework, which is desirable from the practical point of view, with overall good performance. This is confirmed in a subsequent study by [40].
While the focus of the present paper is on variable selection and oracle inequalities for the $\ell_{2}$ loss, prediction errors are also explicitly derived by [48].  [36] revisit the adaptive Lasso [52, 19, 51] as well as the Thresholded Lasso with refitting [47, 48], in a high-dimensional linear model, and study prediction error and bound the number of false positive selections. We refer to [16][1][4] and [33] for related work on complexity regularization criteria. In a subsequent work, [50] develop error bounds based on an earlier version of the present paper and applied these to obtain fast rates of convergence for covariance estimation based on a multivariate Gaussian graphical model. There we show comprehensive numerical results involving cross-validation to choose penalty $\lambda_{n}$ and thresholding $t_{0}$ parameters. We mention that a series of recent papers [29, 31, 23, 32, 49] show that RE condition holds for a broader class of random matrices with complex row/columnwise (or both) dependencies once the sample size is sufficiently large.
3 Proof sketch for the main result
We will now describe the main ideas of our analysis in this section. Combining Theorem 2.4 with Lemmas 2.5 and 2.3 allows us to prove Theorem 2.1, which we will elaborate in more details in Section 6.1. For now, we highlight the important differences between our results and a previous result on the Lasso by [3] (cf. Theorem 7.2), which we refer to as the BRT results. While a bound of $O_{P}(\lambda\sigma\sqrt{s})$ on the $\ell_{2}$ loss as obtained by [3] makes sense when all signals are strong, significant improvements are needed for the general case where $\beta_{\min}$ is not bounded from below.
In the present work, the goal is to investigate sufficient conditions under which we could achieve a bound of $O_{P}(\lambda\sigma\sqrt{s_{0}})$ on the $\ell_{2}$ loss for both the Lasso and the Thresholded Lasso. Given such error bound for the Lasso, thresholding of an initial estimator $\beta_{\text{\rm init}}$ at the level of $\asymp\sigma\sqrt{2\log p/n}$ will select nearly the best subset of variables in the spirit of Theorem 2.1. Some more comments.
(a) As stated in Theorem 2.1, we use $\textnormal{{RE}}(s_{0},4,X)$, for which we fix the sparsity level at $s_{0}$ and $k_{0}=4$, and sparse eigenvalue conditions (6) and (5). While the constants in association with the BRT results depend on $K^{2}(s,3)$, the constants in association with the Lasso and the Thresholded Lasso crucially depend on $K^{2}(s_{0},4)$, $\Lambda_{\max}(2s)$, $\Lambda_{\min}(2s_{0})$, and $\theta_{s_{0},2s_{0}}$ (cf. (30)).
(b) We note that the lower sparse eigenvalue condition $\Lambda_{\min}(2s)>0$ (6) is implied by, and hence is weaker than the $\textnormal{{RE}}(s,3,X)$ condition. Moreover, it is possible to prove Theorem 2.1 even if we leave condition (6) out. In particular, we note that so long as $|I|\leq 2s_{0}$, then $\textnormal{{RE}}(s_{0},4,X)$ already implies that (35) holds for $c=2$ [3]; cf. Remark 2.9. We will not pursue such optimizations in the present work.
(c) We note that in the $\textnormal{{RE}}(s,3,X)$ condition as required by [3] to achieve the $\ell_{2}$ loss of $O_{P}(\lambda\sigma\sqrt{s})$: while $k_{0}=3$ is chosen, they fix sparsity at $s$ instead of $s_{0}$, which is not ideal when $s_{0}$ is much smaller than $s$. We emphasize that in the $\textnormal{{RE}}(s_{0},4,X)$ condition that we impose, $k_{0}=4$ is rather arbitrarily chosen; in principle, it can be replaced by any number that is strictly larger than $3$. In the context of compressed sensing, RE conditions can also be taken as a way to guarantee recovery for anisotropic measurements [31]. Results by [31] reveal that for RE conditions with a smaller $s_{0}$, we need correspondingly smaller sample size $n$ in order for the random design matrix $X$ of dimension $n\times p$ to satisfy such a condition, when the independent row vectors of $X_{i},i=1,\ldots,n$ have covariance $\Sigma(X_{i})={\textbf{E}}X_{i}\otimes X_{i}={\textbf{E}}X_{i}X_{i}^{T}$ satisfying $\textnormal{{RE}}(s_{0},(1+\varepsilon)k_{0},\Sigma^{1/2})$ condition in the sense that for any $\varepsilon>0$, for all $\upsilon\not=0$,
$\frac{1}{K(s_{0},(1+\varepsilon)k_{0},\Sigma^{1/2})}\stackrel{{\scriptstyle\triangle}}{{=}}\min_{\stackrel{{\scriptstyle J_{0}\subseteq[p]}}{{|J_{0}|\leq s_{0}}}}\min_{\left\lVert\upsilon_{J_{0}^{c}}\right\rVert_{1}\leq k_{0}\left\lVert\upsilon_{J_{0}}\right\rVert_{1}}\;\;{\left\lVert\Sigma^{1/2}\upsilon\right\rVert_{2}}/{\left\lVert\upsilon_{J_{0}}\right\rVert_{2}}>0.$
(d) We impose an explicit upper bound on $\Lambda_{\max}(2s)$, which is absent from the paper by [3], in order to obtain the tighter bounds in the present work for both the Lasso and the Thresholded Lasso. This condition is required by our OLS refitting procedure as stated in Lemmas 2.7 and 2.8. This is consistent with the fact that known oracle inequalities for the Dantzig and Gauss-Dantzig selectors are proved under the UUP which impose tighter upper and lower sparse eigenvalue bounds in the sense that $\theta_{s,2s}+\delta_{2s}<1$. See also [43] and the exposition in Section 4.3.
3.1 Proof sketch of Theorems 2.1 and 2.4
We specify the initial Lasso estimator $\beta_{\text{\rm init}}$, and the parameters $\lambda_{n}$, $t_{0}$ in Theorem 2.4 and Lemma 2.3 respectively. Throughout this section, let $h=\beta_{\text{\rm init}}-\beta^{\mathop{\natural}\kern 0.60275pt}({T_{0}})$ be as in Lemma 2.3, where $T_{0}$ denotes the largest $s_{0}$ coordinates of $\beta$ in absolute values. Let $T_{1}$ be the $s_{0}$ largest positions of $h$ outside of $T_{0}$. Denote by $T_{01}=T_{0}\cup T_{1}$. In this section, we elaborate on the $\ell_{p},p=1,2$, loss on $h_{T_{0}}$ and $h_{T_{0}^{c}}$, and their implications on variable selection using thresholding. As mentioned, improving the bounds on each component of $h$ will result in a tighter upper bound on controlling the bias; cf. Lemma 2.3. Specifically, we will show in the supplementary Section D that on ${\mathcal{T}}_{a}$, (19) holds for $h=\beta_{\text{\rm init}}-\beta^{\mathop{\natural}\kern 0.60275pt}({T_{0}})$, where $D^{\prime}_{0}$ and $D_{1}$ are as in (26) and (27).
These bounds ensure that under the RE and sparse eigenvalue conditions, both the $\ell_{2}$ loss on the set $T_{0}$ of significant coefficients, namely, $\left\lVert h_{T_{0}}\right\rVert_{2}=\left\lVert(\beta_{\text{\rm init}}-\beta)_{T_{0}}\right\rVert_{2}$, and the $\ell_{1}$ (and the $\ell_{2}$) norm of the estimated coefficients on $T_{0}^{c}$, namely, $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}=\left\lVert\beta_{\text{\rm init},T_{0}^{c}}\right\rVert_{1}$ are tightly bounded with respect to $\left\lvert T_{0}\right\rvert$ for the Lasso; hence we obtain an oracle inequality on the $\ell_{2}$ loss in the sense of (11).
First, by definition of $h$ and $T_{0}$, we aim to keep variables in $T_{0}$, while for variables not in $T_{0}$, we may trim these off. It is also clear by Lemma 2.3 that we cannot cut too many “significant” variables in $T_{0}$ by following the thresholding rules in our proposal; for example, for those that are $\geq\lambda\sigma\sqrt{s_{0}}$, we can cut at most a constant number of them. Now, what do we do with those in $T_{1}$? The fate of these variables is pretty much up to the choice of the threshold for a given $\beta_{\text{\rm init}}$, knowing these are the largest in magnitude in $h$ (and $\beta_{\text{\rm init}}$) outside of $T_{0}$ and hence most likely to be included in model $I$. Moreover, even if we were able to retain all variables in $T_{0}$, we will include at least some variables in $T_{1}$ when the selection set has size $\left\lvert I\right\rvert>s_{0}$; in fact, we will include nearly all variables in $T_{1}$ when the set $I$ has size close to $2s_{0}$. In particular,
$\left|I\cap T^{c}_{0}\right|$ $\leq$ ${\left\lVert h_{T_{0}^{c}}\right\rVert_{1}}/{t_{0}}\leq D_{1}\lambda_{n}s_{0}/(f_{0}\lambda_{n})\leq(D_{1}/f_{0})s_{0}\quad\text{in case}\quad t_{0}=f_{0}\lambda_{n}$
for some $f_{0}>0$ by (19). Before we continue, we state Lemma 3.1, which is the same (up to normalization) as Lemma 3.1 [9], to illuminate the roles of sets $T_{0},T_{1}$ in the overall bounds on $\left\lVert h\right\rVert_{2}$. We note that in their original statement, the UUP condition is assumed; a careful examination of their proof shows that it is a sufficient but not necessary condition; indeed we only need to assume that sparse eigenvalues are bounded, namely, $\Lambda_{\min}(2s_{0})>0$ and $\Lambda_{\max}(2s_{0})<\infty$. We also state Lemma 3.2.
Lemma 3.1.
Suppose $\Lambda_{\min}(2s_{0})>0$ and $\Lambda_{\max}(2s_{0})<\infty$. Then
$\left\lVert h_{T_{01}}\right\rVert_{2}$ $\leq$ $\frac{1}{\sqrt{\Lambda_{\min}(2s_{0})}}\left\lVert Xh\right\rVert_{2}/\sqrt{n}+\frac{\theta_{s_{0},2s_{0}}}{\Lambda_{\min}(2s_{0})}\left\lVert h_{T_{0}^{c}}\right\rVert_{1}/\sqrt{s_{0}},$ (38)
$\left\lVert h_{T_{01}^{c}}\right\rVert_{2}^{2}$ $\leq$ $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}^{2}\sum_{k\geq s_{0}+1}1/k^{2}\;\leq\;\left\lVert h_{T_{0}^{c}}\right\rVert_{1}^{2}/s_{0}\;\text{ and thus}$ (39)
$\left\lVert h\right\rVert_{2}^{2}$ $\leq$ $\left\lVert h_{T_{01}}\right\rVert_{2}^{2}+s_{0}^{-1}\left\lVert h_{T_{0}^{c}}\right\rVert^{2}_{1}.$
Lemma 3.2.
[31] Denote by
${\mathcal{C}}(s_{0},k_{0}):=\left\{x\in\mathbb{R}^{p},\exists{J\subseteq[p],|J|=s_{0}}\text{ s.t.}\left\lVert x_{J^{c}}\right\rVert_{1}\leq k_{0}\left\lVert x_{J}\right\rVert_{1}\right\}.$ (40)
Let $T_{0}$ denote the locations of the largest coefficients of $x$ in absolute values. Then $\left\lVert x\right\rVert_{2}\leq\sqrt{1+k_{0}}\left\lVert x_{T_{0}}\right\rVert_{2}$ for $x\in{\mathcal{C}}(s_{0},k_{0})$.
Now we provide a proof sketch for Theorem 2.4. For $u=(u_{1},\ldots,u_{n})\in\mathbb{R}^{n}$, define the empirical norm of $u$ by $\left\lVert u\right\rVert_{n}^{2}=\left\lVert u\right\rVert_{2}^{2}/n.$ The proof of the original Theorem 5.1 in [48] draws upon techniques from a concurrent work by [36] and uses an elementary inequality for the Lasso; cf. the supplementary (D.1). Denote by $\beta_{0}:=\beta^{\mathop{\natural}\kern 0.60275pt}({T_{0}})$. The current proof replaces the supplementary (D.1) for the Lasso with the following updated inequality:
$\left\lVert X(\beta_{\text{\rm init}}-\beta)\right\rVert^{2}_{n}+\left\lVert X(\beta_{\text{\rm init}}-\beta_{0})\right\rVert^{2}_{n}\leq\left\lVert X(\beta-\beta_{0})\right\rVert^{2}_{n}$ (41)
$+\frac{2}{n}\epsilon^{T}X(\beta_{\text{\rm init}}-\beta_{0})+2\lambda_{n}(\left\lVert\beta_{0}\right\rVert_{1}-\lVert\beta_{\text{\rm init}}\rVert_{1}).$
See Eq(20) [11], where we set $\bar{\beta}=\beta_{0}=\beta^{\mathop{\natural}\kern 0.60275pt}({T_{0}})$ and $\bar{\delta}=\beta_{\text{\rm init}}-\beta_{0}=h$.
Denote by $\delta=\beta_{\text{\rm init}}-\beta$. Similar to the proof in [48], we have a deterministic proof on event ${\mathcal{T}}_{a}$, except that now on ${\mathcal{T}}_{a}$,
$\left\lVert X\delta\right\rVert^{2}_{n}+\left\lVert Xh\right\rVert^{2}_{n}+\lambda_{n}\left\lVert h_{T_{0}^{c}}\right\rVert_{1}\leq\left\lVert X(\beta-\beta_{0})\right\rVert^{2}_{n}+3\lambda_{n}\left\lVert h_{T_{0}}\right\rVert_{1}.$ (42)
We differentiate between three cases under event ${\mathcal{T}}_{a}$.
  1. 1.
    In the first case, suppose
    $\left\lVert X\delta\right\rVert^{2}_{n}+\left\lVert Xh\right\rVert^{2}_{n}\geq\left\lVert X(\beta-\beta_{0})\right\rVert_{n}^{2};$ (43)
    Then, $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}\leq 3\left\lVert h_{T_{0}}\right\rVert_{1},\text{ and hence }\;\;h\in{\mathcal{C}}(s_{0},3).$
    We will show that on event ${\mathcal{T}}_{a}$, for $\lambda_{n}=d_{0}\lambda\sigma$,
    $\left\lVert h_{T_{0}}\right\rVert_{2}$ $\leq$ $K(s_{0},3)\left\lVert X\beta-X\beta_{0}\right\rVert_{n}+3K^{2}(s_{0},3)\lambda_{n}\sqrt{s_{0}};$ (44)
    Moreover, the bounds on $\left\lVert X\delta\right\rVert_{n}$ and $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ (and $\left\lVert h\right\rVert_{1}$) follow from (45):
    $\left\lVert X\delta\right\rVert^{2}_{n}+\lambda_{n}\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ $\leq$ $\left\lVert X(\beta-\beta_{0})\right\rVert_{n}^{2}+(3K(s_{0},3)\lambda_{n}\sqrt{s_{0}}/2)^{2},$ (45)
    where $\left\lVert X(\beta-\beta_{0})\right\rVert_{n}\leq\lambda\sigma\sqrt{s_{0}\Lambda_{\max}(s-s_{0})}$. Then we have
    $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ $\leq$ $\left\lVert X\beta-X\beta_{0}\right\rVert^{2}_{n}/(\lambda_{n})+(3K/2)^{2}\lambda_{n}s_{0}\leq D_{1,a}\lambda\sigma s_{0},$
    $\text{ where}\;\;D_{1,a}$ $:=$ $\Lambda_{\max}(s-s_{0})/d_{0}+9d_{0}K(s_{0},3)^{2}/{4}.$ (46)
  2. 2.
    In the second case, suppose
    $\left\lVert X\delta\right\rVert^{2}_{n}+\left\lVert Xh\right\rVert^{2}_{n}\leq\lambda_{n}\left\lVert h_{T_{0}}\right\rVert_{1}.$ (47)
    We will show that on event ${\mathcal{T}}_{a}$, $h\in{\mathcal{C}}(s_{0},4)$ and moreover,
    $(2\left\lVert X\delta\right\rVert_{n})\vee\left\lVert Xh\right\rVert_{n}\leq\lambda_{n}\sqrt{s_{0}}K(s_{0},4),$ (48)
[…]
    $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}\leq 4\left\lVert h_{T_{0}}\right\rVert_{1}\leq 4\lambda_{n}K(s_{0},4)^{2}s_{0}.$ (49)
  3. 3.
    Finally, we consider
    $\lambda_{n}\left\lVert h_{T_{0}}\right\rVert_{1}\leq\left\lVert X\delta\right\rVert^{2}_{n}+\left\lVert Xh\right\rVert^{2}_{n}\leq\left\lVert X\beta-X\beta_{0}\right\rVert_{n}^{2}.$ (50)
    Then clearly (22) also holds with no need to invoke $\textnormal{{RE}}(s_{0},c_{0},X)$ condition (for any $c_{0}$) at all:
    $\left\lVert X\delta\right\rVert_{n}^{2}\vee\left\lVert Xh\right\rVert_{n}^{2}$ $\leq$ $\left\lVert X\beta-X\beta_{0}\right\rVert_{n}^{2}\leq\Lambda_{\max}(s-s_{0})\lambda^{2}\sigma^{2}s_{0}\text{ and }$
    $\left\lVert h_{T_{0}}\right\rVert_{1}$ $\leq$ $\left\lVert X\beta-X\beta_{0}\right\rVert_{n}^{2}/\lambda_{n}\asymp\lambda_{n}s_{0}.$
    Moreover, we have $\left\lVert h\right\rVert_{1}\leq 4\Lambda_{\max}(s-s_{0})\lambda\sigma s_{0}/(d_{0})$ and $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}\leq D_{1,c}\lambda\sigma s_{0}$, where $D_{1,c}=3\Lambda_{\max}(s-s_{0})/d_{0}$, since
    $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ $\leq$ $3\left\lVert X\beta-X\beta_{0}\right\rVert_{n}^{2}/\lambda_{n}\leq 3\Lambda_{\max}(s-s_{0})\lambda\sigma s_{0}/(d_{0}).$ (51)
    Hence although the vector $h$ may not satisfy the cone constraint, both components of $h$ have bounded $\ell_{1}$ norm of order $\asymp\lambda\sigma s_{0}$.
Using (45), (136), (48), and (50), we conclude that (22) holds. Although the new proof is introduced to get rid of the constant factor $2$ in front of $\left\lVert X\beta-X_{I_{0}}\beta_{I_{0}}\right\rVert_{n}=\left\lVert X(\beta-\beta_{0})\right\rVert_{n}$ in (22) in [48], we are not optimizing the constants in this paper; cf. the proof of the original Theorem 5.1 in [48]. For example, the constant $3$ in (42) can be further reduced following [11] using (125). Moreover, combining (44), (46), (49), and (51), we have the expression of (27) for $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$, namely, $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}\leq D_{1}\lambda\sigma s_{0}$. We demonstrate the tightness of these bounds in Fig. 2.
Next we bound $\left\lVert h\right\rVert_{2}$ in view of Lemma 3.1. Lemma 3.3 also provides the upper bounds on $\left\lVert h_{T_{01}}\right\rVert_{2}$, leading to the expression of $D_{0}$ as in (23). We combine the upper bounds in Lemma 3.1 and the supplementary Lemma D.2 in the proof of Lemma 3.3. We prove Lemmas 3.1 and 3.3, and Theorem 2.4 in the supplementary Sections CE, and D respectively.
Lemma 3.3.
Under the settings of Theorem 2.4, we have under event ${\mathcal{T}}_{a}$,
$\text{\bf Case 1}:$ $\left\lVert h\right\rVert_{2}\leq 2\left\lVert h_{T_{01}}\right\rVert_{2}\leq 2D_{0,a}\lambda\sigma\sqrt{s_{0}}\;\text{where}$
$D_{0,a}$ $\leq$ $\sqrt{2}\big(\sqrt{\Lambda_{\max}(s-s_{0})}K(s_{0},3)+3d_{0}K^{2}(s_{0},3)\big);$ (52)
$\text{\bf Case 2}:$ $\left\lVert h\right\rVert_{2}\leq\sqrt{5}\left\lVert h_{T_{01}}\right\rVert_{2}\leq\sqrt{10}K^{2}(s_{0},4)d_{0}\lambda\sigma\sqrt{s_{0}};$
$\text{\bf Case 3}:$ $\left\lVert h_{T_{01}}\right\rVert_{2}\leq\frac{1}{\sqrt{\Lambda_{\min}(2s_{0})}}\big(\left\lVert Xh\right\rVert_{n}+\frac{\ell(s_{0})}{\sqrt{s_{0}}}\left\lVert h_{T_{0}^{c}}\right\rVert_{1}\big)\leq D\lambda\sigma\sqrt{s_{0}},$ (53)
$\text{where}\;D:=\sqrt{\frac{\Lambda_{\max}(s-s_{0})}{\Lambda_{\min}(2s_{0})}}\big(1+\frac{3\ell(s_{0})}{d_{0}}\sqrt{\Lambda_{\max}(s-s_{0})}\big)$
and $\ell(s_{0})$ is as defined in (D.2). Moreover, under event ${\mathcal{T}}_{a}$, we have for $D_{0}$ in (23) and $D_{1}$ in (27): $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}\leq D_{1}\lambda\sigma s_{0}$,
$\left\lVert h_{T_{01}}\right\rVert_{2}$ $\leq$ $D_{0}\lambda\sigma\sqrt{s_{0}},\quad\left\lVert h_{T_{01}^{c}}\right\rVert_{2}\leq\left\lVert h_{T_{0}^{c}}\right\rVert_{1}/\sqrt{s_{0}}\leq D_{1}\lambda\sigma\sqrt{s_{0}},$
[…]
$\left\lVert\beta_{\text{\rm init}}-\beta\right\rVert_{2}$ $\leq$ $\left\lVert h\right\rVert_{2}+\lambda\sigma\sqrt{s_{0}}\leq[\sqrt{D_{0}^{2}+D_{1}^{2}}+1]\lambda\sigma\sqrt{s_{0}}.$
Remark 3.4.
One can also use $\textnormal{{RE}}(s_{0},6,X)$ in Case 2 to further tighten the constants in (51) and (53) for Case 3, for example, by considering
$\text{ in Case 2},\quad\left\lVert X\delta\right\rVert^{2}_{n}+\left\lVert Xh\right\rVert^{2}_{n}\leq 3\lambda_{n}\left\lVert h_{T_{0}}\right\rVert_{1},$
instead of (136), and in Case 3, instead of (50),
$3\lambda_{n}\left\lVert h_{T_{0}}\right\rVert_{1}$ $\leq$ $\left\lVert X\delta\right\rVert^{2}_{n}+\left\lVert Xh\right\rVert^{2}_{n}\leq\left\lVert X\beta-X\beta_{0}\right\rVert_{n}^{2}.$
Now clearly $\frac{\Lambda_{\max}(s-s_{0})}{\Lambda_{\min}(2s_{0})}\geq(\frac{\Lambda_{\max}(s-s_{0})}{\Lambda_{\min}(2s_{0})})^{1/2}$ for $s\geq 2s_{0}$ since
$\Lambda_{\min}(2s_{0})\leq\Lambda_{\min}(s_{0})\leq\Lambda_{\max}(s-s_{0}),\;\text{ and }1/\sqrt{\Lambda_{\min}(2s_{0})}\leq\sqrt{2}K(s_{0},1).$
Compared with Case 1 and Case 2, where $\textnormal{{RE}}(s_{0},k_{0},X)$ holds for $k_{0}=3,4$ respectively, we have an extra term on the bound for $\left\lVert h_{T_{01}}\right\rVert_{2}$ in (53), namely, $\ell(s_{0})\left\lVert h_{T_{0}^{c}}\right\rVert_{1}/\sqrt{s_{0}}\asymp\lambda\sigma\sqrt{s_{0}}$; cf. (51) and (D.2).
Remark 3.5.
Under the UUP condition as in Definition G.1:
$\forall s_{0}\leq s,\quad\theta_{s_{0},2s_{0}}+\delta_{2s_{0}}\leq\theta_{s,2s}+\delta_{2s}<1$ (54)
since $\theta_{s,2s}$ and $\delta_{s}$ are nondecreasing in $s$. Thus we have $\theta_{s_{0},2s_{0}}<1-\delta_{2s_{0}}\leq\Lambda_{\min}(2s_{0})$. When such tight bounds are not available, we can still use the bounds in (30) to control $\theta_{s_{0},2s_{0}}$. Moreover, suppose $\theta_{s_{0},2s_{0}}<\Lambda_{\min}(2s_{0})$, which holds under UUP (54), one can get rid of the factor $\sqrt{\Lambda_{\max}(s_{0})}$ in (D.2), where recall $\ell(s_{0})=({\theta_{s_{0},2s_{0}}}/{\sqrt{\Lambda_{\min}(2s_{0})}})\wedge\sqrt{\Lambda_{\max}(s_{0})}$, by bounding $\left\lVert h_{T_{01}}\right\rVert_{2}$ for Case 3 following (38) instead of the supplementary Lemma D.2.
4 On Type II errors and $\ell_{2}$-loss optimality
So far, we have focused on controlling Type I errors, which would be meaningless if significant variables are all missing. Our goal in this section is to show when $\beta_{\min,A_{0}}$ (15) is sufficiently large, we have $A_{0}\subset I$ while achieving the sparse oracle inequalities. Under the RE and sparse eigenvalue conditions, this result is shown in Theorem 6.3 [48], which is a corollary of Lemma 4.3. We include it here for self-containment. First, we state Proposition 4.1.
Proposition 4.1.
Let $A_{0}:=\{j:|\beta_{j}|>\sqrt{2\log p/n}\sigma\}$. Let $T_{0}$ denote positions of the $s_{0}$ largest coefficients of $\beta$ in absolute values, where $s_{0}$ is defined in (13). Let $a_{0}=\left|A_{0}\right|$ denote the cardinality of $A_{0}$ (see also (15)). Then $\forall c^{\prime}>1/2$, we have $\left|\{j\in T_{0}^{c}:|\beta_{j}|\geq\sqrt{\log p/(c^{\prime}n)}\sigma\}\right|\leq(2c^{\prime}-1)(s_{0}-a_{0})$.
Again order $\beta_{j}$’s in decreasing order of magnitude: $|\beta_{1}|\geq|\beta_{2}|\geq...\geq|\beta_{p}|$. Let $T_{0}=\{1,\ldots,s_{0}\}$. One could choose another target set: for example $\{j:|\beta_{j}|\geq\sqrt{\log p/(c^{\prime}n)}\sigma\}$, for some $\log p/2>c^{\prime}>1/2$. Moreover, we consider the consequence of setting $t_{0}\in[\sigma\sqrt{2/(n)},\sigma\sqrt{2\log p/(n)}]$. In particular, when we set $c^{\prime}=\log p/2$, we have
$\left|\{j\in T_{0}^{c}:|\beta_{j}|\geq\sigma\sqrt{2/(n)}\}\right|\leq(\log p-1)(s_{0}-a_{0}).$ (55)
We prove Proposition 4.1 in Section 6.3. We first show in Lemma 4.2 that under no restriction on $\beta_{\min}$, we achieve an oracle bound on the $\ell_{2}$ loss, which depends only on the $\ell_{2}$ loss of the initial estimator on the set $T_{0}$. Bounds in Lemma 2.3 are special cases of (56). In Lemma 4.3, we impose a lower bound on $\beta_{\min,A_{0}}$ as in (58) in order to recover the subset of variables in $A_{0}$, while achieving the nearly ideal $\ell_{2}$ loss with a sparse model $I$ in view of Lemma 2.7. We prove Lemma 4.2 and 4.3 in Sections 6.6 and 6.7 respectively.
Lemma 4.2.
(A deterministic result on the bias.) Let $\beta_{\text{\rm init}}$ be an initial estimator. Let $\lambda:=\sqrt{2\log p/n}$ and $h\in\mathbb{R}^{p}$ be as defined in Lemma 2.3. Suppose we choose a thresholding parameter $t_{0}$ and set
$I=\{j:\left|\beta_{j,\text{\rm init}}\right|\geq t_{0}\}.$
Then for ${\mathcal{D}}:=[p]\setminus I$, we have for ${\mathcal{D}}_{11}:={\mathcal{D}}\cap A_{0}$ and $a_{0}=\left|A_{0}\right|$,
$\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}^{2}$ $\leq$ $(s_{0}-a_{0})\lambda^{2}\sigma^{2}+(t_{0}\sqrt{a_{0}}+\left\lVert h_{{\mathcal{D}}_{11}}\right\rVert_{2})^{2}.$ (56)
Suppose $t_{0}<\beta_{\min,A_{0}}$ as defined in (15). Then (56) can be replaced by
$\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}^{2}$ $\leq$ $(s_{0}-a_{0})\lambda^{2}\sigma^{2}+\left\lVert h_{{\mathcal{D}}_{11}}\right\rVert_{2}^{2}\left({\beta_{\min,A_{0}}}/{(\beta_{\min,A_{0}}-t_{0})}\right)^{2}.$ (57)
Lemma 4.3.
(A deterministic result on the bias and model selection.) Let $\beta_{\text{\rm init}}$ be an initial estimator. Let $\lambda:=\sqrt{2\log p/n}$ and $h\in\mathbb{R}^{p}$ be as defined in Lemma 2.3. Suppose for $\beta_{\min,A_{0}}$ as defined in (15), it holds that
$\beta_{\min,A_{0}}\geq\left\lVert h_{A_{0}}\right\rVert_{\infty}+\min\left\{(s_{0})^{-1/2}\left\lVert h_{T_{0}^{c}}\right\rVert_{2},\;(s_{0})^{-1}\left\lVert h_{T_{0}^{c}}\right\rVert_{1}\right\}.$ (58)
Now we choose a thresholding parameter $t_{0}$ such that
$\text{for some }\;\breve{s}_{0}\in[s_{0},s),\quad\beta_{\min,A_{0}}-\left\lVert h_{A_{0}}\right\rVert_{\infty}\geq t_{0}$
$\geq$ $\min\left\{(\breve{s}_{0})^{-1/2}\left\lVert\beta_{\text{\rm init},T_{0}^{c}}\right\rVert_{2},(\breve{s}_{0})^{-1}\left\lVert\beta_{\text{\rm init},T_{0}^{c}}\right\rVert_{1}\right\}$
holds and set $I=\{j:\left|\beta_{j,\text{\rm init}}\right|\geq t_{0}\}$. Then we have
$A_{0}\subset I$ $\text{ and }\quad|I\cap T_{0}^{c}|\leq\breve{s}_{0},\quad\text{ and hence}$ (60)
$|I|\leq s_{0}+\breve{s}_{0}$ $\text{ and }\quad\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}^{2}\leq(s_{0}-a_{0})\lambda^{2}\sigma^{2}.$ (61)
Choosing the set $A_{0}$ as in Proposition 4.1 is rather arbitrary; one could for example, consider the set of variables that are strictly above $\lambda\sigma/2$. In Theorem 4.4, we show that one can indeed recover a subset $A_{0}$ of variables accurately, for $A_{0}$ as defined in Proposition 4.1, when $\beta_{\min,A_{0}}:=\min_{j\in A_{0}}|\beta_{j}|$ is large enough (relative to the $\ell_{2}$ loss of an initial estimator $\beta_{\text{\rm init}}$ under the RE condition on the set $A_{0}$); in addition, a small number of extra variables from $T_{1}\subset T_{0}^{c}:=[p]\setminus T_{0}$ are also possibly included in the model $I$. We mention in passing that changing the coefficients of $\beta_{A_{0}}$ will not change the values of $s_{0}$ or $a_{0}$, so long as their absolute values stay strictly above $\lambda\sigma$; See Section 4.2. Compared with the almost exact sparse recovery result in Theorem 1.1 [47], we have relaxed the restriction on $\beta_{\min}$: rather than requiring all non-zero entries to be large in absolute values, we only require those in a subset $A_{0}$ to be recovered to be large.
Theorem 4.4.
Under the settings of Theorem 2.1, let $\beta_{\text{\rm init}}$ be the Lasso estimator as in (2). Let $T_{0}$ denote the largest $s_{0}$ coordinates of $\beta$ in absolute values. Let $T_{1}$ be the largest $s_{0}$ positions of $\beta_{\text{\rm init}}$ outside of $T_{0}$. Let $T_{01}=T_{0}\cup T_{1}$. Suppose (5), (6), and $\textnormal{{RE}}(s_{0},4,X)$ condition hold. Let $D_{0},D_{1}$ be as in (23) and (27). Suppose for some constants $\breve{D}_{1}>D_{1}$, $\breve{D}_{0}>D_{0}$,
$\beta_{\min,A_{0}}:=\min_{j\in A_{0}}|\beta_{j}|\geq D_{0}\lambda\sigma\sqrt{s_{0}}+(\breve{D}_{1}\wedge\breve{D}_{0})\lambda\sigma$ (62)
where $\lambda:=\sqrt{2\log p/n}$. Choose a thresholding parameter $t_{0}$ and set
$I=\{j:\left|\beta_{j,\text{\rm init}}\right|\geq t_{0}\},\;\text{ where }t_{0}=(\breve{D}_{1}\wedge\breve{D}_{0})\lambda\sigma.$
Let $\hat{\beta}_{I}=(X_{I}^{T}X_{I})^{-1}X_{I}^{T}Y$ and $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)$ be the 0-extended version of $\hat{\beta}_{I}$ such that $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}_{I^{c}}=0$ and $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}_{I}=\hat{\beta}_{I}$. Then on event ${\mathcal{T}}_{a}$, we have $\left\lvert I\right\rvert\leq 2s_{0}$,
$A_{0}\subset I,\quad\;I\cap T_{01}^{c}=\emptyset,\;\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}\leq\lambda\sigma\sqrt{s_{0}-a_{0}},\text{ and}$
$\left\lVert\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)-\beta\right\rVert_{2}^{2}\leq\big(1+\frac{(\Lambda_{\max}(2s)-\Lambda_{\min}(2s))^{2}+8(1+a)}{2\Lambda_{\min}^{2}(|I|)}\big)s_{0}\lambda^{2}\sigma^{2}.$ (63)
Proof.
By definition of $T_{0}$, we have $T_{0}\subset S$ and recall ${S_{{\mathcal{D}}}}={\mathcal{D}}\cap S$. Moreover, $I\cap{S_{{\mathcal{D}}}}=\emptyset$ by definition. By Lemma 3.3, we have for on ${\mathcal{T}}_{a}$,
$\left\lVert\beta_{\text{\rm init},T_{0}^{c}}\right\rVert_{1}=\left\lVert h_{T_{0}^{c}}\right\rVert_{1}\leq D_{1}\lambda\sigma s_{0}\text{ and }\;\left\lVert h_{T_{01}}\right\rVert_{2}\leq D_{0}\lambda\sigma\sqrt{s_{0}}.$ (64)
First, we find the lower bound on $\beta_{\min,A_{0}}$ so that $A_{0}\subset I$. Suppose there exists a threshold $t_{0}>0$ such that one of the following conditions holds,
[…]
ensuring no FNs from $A_{0}$. Next, we derive a lower bound on $t_{0}$. Suppose
$t_{0}>(D_{0}\wedge D_{1})\lambda\sigma\geq\left\lVert\beta_{\text{\rm init},T_{01}^{c}}\right\rVert_{\infty}=\left\lVert h_{T_{01}^{c}}\right\rVert_{\infty},$ (66)
we can then eliminate all variables in $T_{01}^{c}$ from model $I$, where $\beta_{\text{\rm init},T_{0}^{c}}=h_{T_{0}^{c}}$ and $T_{1}\subset T_{0}^{c}$. To see the second inequality in (66), recall the $k$th largest value of $\left\lvert h_{T_{0}^{c}}\right\rvert$ obeys $\left|h_{T_{0}^{c}}\right|_{(k)}\leq\left\lVert h_{T_{0}^{c}}\right\rVert_{1}/k$ for $k\geq 1$. Hence by definition of $T_{1}$,
[…]
Moreover, by definition of $T_{1}$, the largest entry in $\left\lvert h_{T_{01}^{c}}\right\rvert$ (entrywise absolute value of vector $h_{T_{01}^{c}}$) is bounded by the average over the top $s_{0}$ largest entries of $\left\lvert h_{T_{0}^{c}}\right\rvert$: $\left|h_{T_{0}^{c}}\right|_{(s_{0}+1)}\leq\frac{1}{s_{0}}\sum_{k=1}^{s_{0}}\left|h_{T_{0}^{c}}\right|_{(k)}=\left\lVert h_{T_{1}}\right\rVert_{1}/s_{0}$, and hence
$\left\lVert\beta_{\text{\rm init},T_{01}^{c}}\right\rVert_{\infty}$ $=$ $\left\lVert h_{T_{01}^{c}}\right\rVert_{\infty}\leq\left\lVert h_{T_{1}}\right\rVert_{1}/s_{0}\leq\sqrt{s_{0}}\left\lVert h_{T_{1}}\right\rVert_{2}/s_{0}$
[…]
In summary, (62) and (66) imply that it is feasible to set
$t_{0}$ $=$ $(\breve{D}_{1}\wedge\breve{D}_{0})\lambda\sigma>(D_{1}\wedge D_{0})\lambda\sigma\;\text{ since}$
$\beta_{\min,A_{0}}$ $\geq$ $\left\lVert h_{T_{0}}\right\rVert_{2}+t_{0}>\left\lVert h_{T_{0}}\right\rVert_{\infty}+(D_{0}\wedge D_{1})\lambda\sigma.$ (67)
Then none of the variables in $T_{01}^{c}$ can be chosen in the thresholding step and hence $I\cap T_{01}^{c}=\emptyset$. Hence, $\left|I\right|\leq 2s_{0}$. On event ${\mathcal{T}}_{a}$, (33) holds by Lemma 2.7. Moreover, $\theta^{2}_{|I|,|{S_{{\mathcal{D}}}}|}$ is bounded as in Lemma 2.8 given $|I|+|{S_{{\mathcal{D}}}}|=|I\cup{S_{{\mathcal{D}}}}|\leq\left\lvert S\right\rvert+\left\lvert I\cap S^{c}\right\rvert\leq s+|I\cap T_{0}^{c}|\leq s+s_{0}\leq 2s$. Thus (63) follows immediately. ∎∎
4.1 Discussions
Theorem 4.4 is an immediate corollary of Theorem 2.1, Lemma 2.7, and Lemma 4.3, except that we now set $\breve{s}_{0}=s_{0}$ everywhere in Lemma 4.3, and assume having an upper estimate $\breve{D}_{1}$ (resp. $\breve{D}_{0}$) of $D_{1}$ (resp. $D_{0}$), so as not to depend on an “oracle” telling us an exact value. Given all other parameters held invariant, a larger $\beta_{\min,A_{0}}$ (e.g., a larger $C_{a}$ for $\beta$ in Figure 1) will result in a tighter bound on the bias $\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}^{2}$ in view of (57). When $\beta_{\min,A_{0}}^{2}$ dominates the upper bound on the RHS of (57), then $A_{0}\cap I^{c}=\emptyset$, since the loss of a single variable from $A_{0}$ will already saturate this upper bound in the order of $O(\lambda^{2}\sigma^{2}s_{0})$. Hence when $\beta_{\min,A_{0}}$ is strong enough in the sense of (4.3), there will be no false negatives from the set $A_{0}$, leading to the stronger and tighter bounds in (60) and (61) that also control false positives and the bias. This is validated in our experiments in Section 5.
In the statement of Lemma 4.3, we assume the knowledge of the error bounds on various norms of $\beta_{\text{\rm init}}-\beta$ and $h=\beta_{\text{\rm init}}-\beta_{T_{0}}$ implicitly (hence the name of “oracle”). In obtaining (62), we may substitute the bounds as derived in Lemma 3.3 in (58), as trivially, for $A_{0}\subseteq T_{0}$ and $h_{A_{0}}=\beta_{\text{\rm init},A_{0}}-\beta_{A_{0}}$,
$\left\lVert h_{A_{0}}\right\rVert_{\infty}$ $\leq$ $\left\lVert h_{A_{0}}\right\rVert_{2}\leq\left\lVert h_{T_{0}}\right\rVert_{2}\leq D_{0}\lambda\sigma\sqrt{s_{0}}\;\text{ and }$ (68)
[…]
Lemma 4.3 explains our results in Theorem 1.1 [46] as well as the numerical results. We also introduce $\breve{s}_{0}$ so that the dependency of $t_{0}$ on the knowledge of $s_{0}$ is relaxed; in particular, it can be used to express a desirable level of sparsity for the model $I$ that one wishes to select. In general, we may assume that $\beta_{\min,A_{0}}\geq 2D_{0}\lambda\sigma\sqrt{s_{0}}$ so as to have a large range of effective thresholding parameters; cf. Section 3. Thus one can increase $t_{0}$ as $\beta_{\min,A_{0}}$ increases in order to reduce the number of false positives while not increasing the number of false negatives from the active set $A_{0}$.
In this case, it is also possible to remove variables from $T_{0}^{c}$ entirely by increasing the threshold $t_{0}$ while strengthening the lower bound on $\beta_{\min,A_{0}}$ by a constant factor. For example, we may set $t_{0}>\left\lVert\beta_{\text{\rm init},T_{0}^{c}}\right\rVert_{\infty}$, rather than $\left\lVert\beta_{\text{\rm init},T_{01}^{c}}\right\rVert_{\infty}$ as in (66). Clearly, $\left\lVert\beta_{\text{\rm init},T_{0}^{c}}\right\rVert_{\infty}=\left\lVert h_{T_{1}}\right\rVert_{\infty}\leq\left\lVert h_{T_{1}}\right\rVert_{2}=O_{P}(\lambda\sigma\sqrt{s_{0}})$ and hence the condition $\beta_{\min,A_{0}}\geq\sqrt{2}\left\lVert h_{T_{01}}\right\rVert_{2}\geq\left\lVert h_{T_{0}}\right\rVert_{2}+\left\lVert h_{T_{1}}\right\rVert_{2}$ ensures
$\beta_{\min,A_{0}}-\left\lVert h_{A_{0}}\right\rVert_{\infty}$ $\geq$ $\beta_{\min,A_{0}}-\left\lVert h_{T_{0}}\right\rVert_{2}\geq\left\lVert h_{T_{1}}\right\rVert_{2}>\left\lVert\beta_{\text{\rm init},T_{0}^{c}}\right\rVert_{\infty},$
which is sufficient for (65) to hold so long as $t_{0}\asymp\left\lVert\beta_{\text{\rm init},T_{0}^{c}}\right\rVert_{\infty}\leq\left\lVert h_{T_{1}}\right\rVert_{2}$. Then following the same analysis as in Theorem 4.4, we can recover $A_{0}$ while eliminating variables from $T_{0}^{c}$; cf. (67). In general, if the strong signals are close to each other in their strength, then a small $\beta_{\min,A_{0}}$ implies that we are in a situation with low signal to noise ratio (low SNR); one needs to carefully tradeoff false positives with false negatives as shown in our numerical results in Section 5 and the supplementary Section I. Such results and their formal statements are omitted from the present paper.
Bounds on $\left\lVert h_{A_{0}}\right\rVert_{\infty}$ are in general harder to obtain than $\left\lVert h_{A_{0}}\right\rVert_{2}$. In general, we can still hope to bound $\left\lVert h_{A_{0}}\right\rVert_{\infty}$ by $\left\lVert h_{A_{0}}\right\rVert_{2}$ as done in Theorem 4.4. Having a tight bound on $\left\lVert h_{T_{0}}\right\rVert_{2}$ (or $\left\lVert h_{T_{0}}\right\rVert_{\infty}$) and $\lVert h_{T_{0}^{c}}\rVert_{2}$ naturally helps relaxing the requirement on $\beta_{\min,A_{0}}$ for Lemma 4.3, while as shown in Lemma 4.2, such tight upper bounds will help us control the model size $\left\lvert I\right\rvert$ and the bias $\left\lVert\beta_{{\mathcal{D}}}\right\rVert$ and therefore achieve a tight bound on the $\ell_{2}$ loss in the statement of Lemma 2.5. We refer to [37] for discussion and further pointers into this literature on information theoretic limits on sparse recovery.
Figure 1: In this model, the component $\beta^{(11)}$ has $a_{0}$ non-zero coordinates with the same magnitude $C_{a}\lambda\sigma=:\beta_{\min,A_{0}}$, where $C_{a}\in\{1.706,8.528\}$ and $\beta_{\min,A_{0}}\in\{0.2,1\}$; the component $\beta^{(12)}$ has $s_{0}-a_{0}$ non-zero coordinates with the same magnitude $C_{m}\lambda\sigma$, where $C_{m}=1/{\sqrt{2}}$ for $s>s_{0}$ and $C_{m}=1$ in case $s_{0}=s$; the component $\beta^{(2)}$ has $s-s_{0}$ non-zero coordinates with the same magnitude $C_{t}\lambda\sigma=:c_{t}\sigma/\sqrt{n}$. See (55). The rest are all 0s. In the exact sparse case, namely, when $s=s_{0}$, all non-zero signals are concentrated on the component $\beta^{(1)}$ without spreading across components of $\beta^{(2)}$.
4.2 Model specification
We now pause to briefly describe our experiment setup so as to further discuss variable selection in the context of Theorem 4.4. Moreover, we generate a class of models with different sparsity $s$ and $\ell_{1}$ norm on $\beta_{T_{0}^{c}}$ to shed light on this connection with the work of [43]. Let $\left\lvert\beta_{j}\right\rvert$’s be ordered as in (14). Let $S=\mathop{\text{\rm supp}\kern 0.86108pt}(\beta)$ and $A_{0},T_{0}$ be as defined in Proposition 4.1. Let $A_{0}=[a_{0}]$ and $T_{0}=[s_{0}]$. Then $T_{0}\setminus A_{0}=\{a_{0}+1,\ldots,s_{0}\}$. Let $\left\lvert S\right\rvert=s$. See Figure 1 for an illustration of the model specification. We divide $\beta$ into 4 components and write $\beta=\beta^{(11)}+\beta^{(12)}+\beta^{(2)}+\beta^{(0)}\in\mathbb{R}^{p}$, where we set for some $C_{a}>1$, $0<C_{m}\leq 1$, $c_{t}\geq 0$,
$\beta_{j}^{(11)}$ $=$ $\pm C_{a}\lambda\sigma\cdot 1_{1\leq j\leq a_{0}},\quad\beta_{j}^{(12)}=\pm C_{m}\lambda\sigma\cdot 1_{a_{0}<j\leq s_{0}},$ (69)
$\beta_{j}^{(2)}$ $=$ $\pm c_{t}(\sigma/\sqrt{n})\cdot 1_{s_{0}<j\leq s}\;\;\;\text{ and}\;\;\beta_{j}^{(0)}=0\cdot 1_{s<j\leq p},$
where $\lambda=\sqrt{2\log p/n}$ and the unspecified coordinates in each component are again set to 0. The first 3 components contain non-zero coordinates. Hence $\left\lVert\beta^{(12)}\right\rVert_{2}=C_{m}\lambda\sigma\sqrt{s_{0}-a_{0}}$ in (69). Since $C_{a}>1$, we have by (13),
[…]
$(s_{0}-a_{0})\lambda^{2}\sigma^{2}(1-C_{m}^{2})$ $\geq$ $\left\lVert\beta^{(2)}\right\rVert_{2}^{2}=\sum_{j>s_{0}}\beta_{j}^{2}=\sum_{j>s_{0}}\min(\beta_{j}^{2},\lambda^{2}\sigma^{2})$ (71)
$\geq$ $(s_{0}-a_{0})(1-C^{2}_{m})\lambda^{2}\sigma^{2}-\lambda^{2}\sigma^{2}.$ (72)
Clearly, for $\beta^{(2)}$, its $\ell_{1}$ norm $\left\lVert\beta_{T_{0}^{c}}\right\rVert_{1}=\left\lVert\beta^{(2)}\right\rVert_{1}$ is proportional to the signal strength $\left\lVert\beta^{(2)}\right\rVert_{2}^{2}$ and is inversely proportional to its $\ell_{\infty}$ norm, since by (71),
$(\text{support})\quad\left\lvert T_{0}^{c}\cap S\right\rvert$ $=$ $s-s_{0}=\left\lVert\beta^{(2)}\right\rVert_{2}^{2}/(c_{t}^{2}\sigma^{2}/n)$ (73)
$\leq$ ${(s_{0}-a_{0})(2\log p)(1-C_{m}^{2})}/{c_{t}^{2}},\;\text{where}$
$(\ell_{2})\quad\left\lVert\beta^{(2)}\right\rVert_{2}$ $=$ $\left\lVert\beta_{{T_{0}^{c}}}\right\rVert_{2}\approx\lambda\sigma\sqrt{(1-C_{m}^{2})(s_{0}-a_{0})},$ (74)
$\text{and}\;(\ell_{1})\quad\left\lVert\beta_{T_{0}^{c}}\right\rVert_{1}$ $=$ $\sqrt{\left\lvert T_{0}^{c}\cap S\right\rvert}\left\lVert\beta_{T_{0}^{c}}\right\rVert_{2}=\left\lVert\beta^{(2)}\right\rVert_{2}^{2}/[c_{t}\sigma/\sqrt{n}]$ (75)
$\leq$ $\frac{\sqrt{2\log p}(1-C_{m}^{2})}{c_{t}}\lambda\sigma(s_{0}-a_{0}).$
Here (75) is essentially tight with a matching lower bound at the same order; cf. (72). In other words, by choosing different values of $C_{t}$, we generate the model class for $\beta$ with different sparsity as in (69) and (76):
$(s-s_{0})C_{t}^{2}$ $=$ $(s_{0}-a_{0})(1-C_{m}^{2}),\text{ where}\;C_{t}=c_{t}/\sqrt{2\log p}\geq 0.$ (76)
By setting the height of $\beta^{(2)}$, or $\left\lVert\beta^{(2)}\right\rVert_{\infty}$ to be $\asymp\sigma/\sqrt{n}$, the size of its support $\left\lvert\mathop{\text{\rm supp}\kern 0.86108pt}(\beta^{(2)})\right\rvert\asymp\log p(s_{0}-a_{0})$ as desired; cf. Proposition 4.1 and (73). Finally (74) holds by (70), (71), and (72). We compare the $\ell_{1}$ condition in (75) with those in [43]; cf. (80).
We now show that the upper bound in (75) is essentially tight. First, the bound in (74) is tight, since
$\left\lVert\beta^{(2)}\right\rVert_{2}^{2}$ $\leq$ $\left\lVert\beta^{(2)}\right\rVert_{1}\left\lVert\beta^{(2)}\right\rVert_{\infty}=[c_{t}\sigma/\sqrt{n}]\left\lVert\beta^{(2)}\right\rVert_{1},\text{ where}$ (77)
$\sum_{j\in T^{c}_{0}}\min(\beta_{j}^{2},\lambda^{2}\sigma^{2})$ $=$ $\sum_{j>s_{0}}\min(\beta_{j}^{2},\lambda^{2}\sigma^{2})=\left\lVert\beta^{(2)}\right\rVert_{2}^{2}$ (78)
$\geq$ $(s_{0}-a_{0})(1-C^{2}_{m})\lambda^{2}\sigma^{2}-\lambda^{2}\sigma^{2}.$
Moreover, (75) is also essentially tight by definition since by (77) and (78),
$\left\lVert\beta^{(2)}\right\rVert_{1}$ $\geq$ $\left\lVert\beta^{(2)}\right\rVert_{2}^{2}/[c_{t}\sigma/\sqrt{n}]\geq((1-C^{2}_{m})(s_{0}-a_{0})-1)\lambda^{2}\sigma/[c_{t}/\sqrt{n}]$ (79)
$\geq$ $((1-C^{2}_{m})(s_{0}-a_{0})-1)\lambda\sigma\frac{\sqrt{2\log p}}{c_{t}}\;\text{ and moreover}$
$\sum_{j\in A_{0}^{c}}\left\lvert\beta_{j}\right\rvert$ $=$ $\left\lVert\beta^{(2)}\right\rVert_{1}+C_{m}\lambda\sigma(s_{0}-a_{0})=\Omega(\lambda\sigma(s_{0}-a_{0}){\sqrt{(\log p)}}/{c_{t}}).$
Finally, (78) follows from (71) and the fact that $\left\lVert\beta^{(12)}\right\rVert_{2}^{2}=\sum_{a_{0}<j\leq s_{0}}\min(\beta_{j}^{2},\lambda^{2}\sigma^{2})=C^{2}_{m}\lambda^{2}\sigma^{2}(s_{0}-a_{0})$.
4.3 Conditions in [43]
Without loss of generality, we order $\left\lvert\beta_{1}\right\rvert\geq\left\lvert\beta_{2}\right\rvert\geq\ldots\geq\left\lvert\beta_{p}\right\rvert$ as in (14). For the Lasso solution $\beta_{\text{\rm init}}$, it is assumed that $\lambda_{n}\geq 2\lambda\sigma\sqrt{(1+a)c^{*}}$ in (2) in Theorem 3 in [43]. Hence we assume $\lambda_{n}\asymp 2\lambda\sigma\sqrt{(1+a)c^{*}}$ throughout our discussion. Roughly speaking, we interpret $H_{q}=\{1,\ldots,q\}$ as the set of large coordinates in $\beta$ that one aims to recover in the settings of [43]. For convenience, we denote by $H_{0}:=[p]\setminus H_{q}=\{q+1,\ldots,p\}$ and $T_{1}^{*}\subset H_{0}$ the index set $\{j\in H_{0}:\left\lvert\beta_{j}\right\rvert\geq\lambda\sigma\}$. The sparsity assumption by [43] is set with the $\ell_{1}$-sparsity on $\beta_{H_{0}}$,
$\sum_{j>q}^{p}\left\lvert\beta_{j}\right\rvert\leq\eta_{1}=\tilde{O}(r_{1}^{2}q\lambda\sigma),\;\text{ where }\eta_{1}\leq 2\lambda\sigma\frac{r_{1}^{2}q}{\sqrt{c^{*}}}=O(\frac{q^{*}}{\sqrt{c^{*}}}\lambda\sigma);$ (80)
Here the $\tilde{O}$ notation may hide some constants including $c^{*}>1$ and $q^{*}>4r_{1}^{2}q$, where $q^{*},c_{*},c^{*},q,r_{1}$ (and $r_{2}$) are all allowed to depend on $n$; cf. (82) to (84). This target set $H_{q}$ is slightly more restrictive than the set $A_{0}=\{1,\ldots,a_{0}\}$, since in $H_{0}$, one may still see large signals; cf. the supplementary (113) and the proof of Theorem 3 [43]. To resolve this discrepancy and to properly interpret results in [43], we first extend $H_{q}$ by $T_{1}^{*}\subset H_{0}$ and denote this extended set by $L(q)$, where
$L(q)$ $:=$ $H_{q}\cup T_{1}^{*}=\{j:\left\lvert\beta_{j}\right\rvert\geq\lambda\sigma\},\;\left\lvert T_{1}^{*}\right\rvert\leq\eta_{1}/(\lambda\sigma)=\tilde{O}(2r_{1}^{2}q),$
[…]
By construction, $A_{0}\subseteq L(q)\subset T_{0}$ by definition of $A_{0}$ (15) and $T_{0}$. Hence
$a_{0}\leq\left\lvert L(q)\right\rvert\leq s_{0},\quad q^{*}=\Omega(s_{0}\sqrt{c^{*}}),\text{ so long as }a_{0}\asymp s_{0};$ (81)
Moreover, [43] require the following $\eta_{2}$ condition in (82) and impose the sparse Riesz conditions with rank $q^{*}$ as in (84) on $X$:
$\eta_{2}/\sqrt{n}:=\max_{A\subset H_{0}}\lVert\sum_{j\in A}\beta_{j}X_{j}\rVert_{2}/\sqrt{n}=\tilde{O}(2r_{2}\sqrt{q}\lambda\sigma),$ (82)
$c_{*}\leq{\left\lVert X_{A}v\right\rVert_{2}^{2}}/{(n\left\lVert v\right\rVert_{2}^{2})}\leq c^{*}\quad\forall A\;\text{with}\quad\left|A\right|=q^{*}\;\text{ and }\;v\in\mathbb{R}^{q^{*}}.$ (83)
Then, by Eq. (2.15) – (2.18) and (3.1) [43],
$M_{1}^{*}q+1:=(2+4r_{1}^{2}+4\sqrt{\kappa}r_{2}+4\kappa)q+1\leq q^{*},\;\;\text{ where}\;\;\kappa:=c^{*}/c_{*},$ (84)
and $r_{1}$ and $r_{2}$ are the same as in (80) and (82) respectively. Both conditions are needed to show an upper bound on $\left\lVert\beta_{H_{0}}\right\rVert_{2}=\tilde{O}(\sqrt{q^{*}}\sigma\lambda)$; cf. (85). Let $A_{1}:=\mathop{\text{\rm supp}\kern 0.86108pt}(\beta_{\text{\rm init}})\cup H_{q}$. Moreover, since $A_{1}^{c}\subset H_{0}$, we have by (80),
$\left\lVert\beta_{\text{\rm init},A_{1}^{c}}-\beta_{A_{1}^{c}}\right\rVert_{1}$ $\leq$ $\left\lVert\beta_{H_{0}}\right\rVert_{1}\leq\eta_{1}\;\;\text{ and }$
$\left\lVert\beta_{\text{\rm init},A_{1}^{c}}-\beta_{A_{1}^{c}}\right\rVert_{2}^{2}$ $\leq$ $\left\lVert\beta_{H_{0}}\right\rVert_{2}^{2}=\tilde{O}(\frac{r_{2}}{\sqrt{c_{*}}}q^{*}\lambda^{2}\sigma^{2}).$ (85)
[43] use (84), (85), and the $\eta_{1}$ and $\eta_{2}$ conditions to bound
$\left\lVert\beta_{\text{\rm init}}-\beta\right\rVert_{2}=\tilde{O}_{P}(\sqrt{q^{*}}\lambda\sigma),\;\text{ and }\;\left\lVert\beta_{\text{\rm init}}-\beta\right\rVert_{1}=\tilde{O}_{P}(q^{*}\lambda\sigma),$ (86)
which correctly depend on $q^{*}$. Recall in our example in Section 4.2, $\left\lVert\beta_{A_{0}^{c}}\right\rVert_{2}\leq(s_{0}-a_{0})^{1/2}\lambda\sigma$, where $H_{q}\subset A_{0}$ and $T_{0}^{c}\subseteq([p]\setminus L(q))\subseteq A_{0}^{c}\subset H_{0}$, and
$\left\lVert\beta_{A_{0}^{c}}\right\rVert_{1}$ $\approx$ $\left\lvert s_{0}-a_{0}\right\rvert\lambda\sigma(c_{t}+\sqrt{\log p})/(\sqrt{2}c_{t})\text{ since }$ (87)
[…]
$\left\lVert\beta_{T_{0}\setminus A_{0}}\right\rVert_{1}$ $=$ $\left\lVert\beta^{(12)}\right\rVert_{1}=\left\lvert s_{0}-a_{0}\right\rvert C_{m}\lambda\sigma=\left\lvert s_{0}-a_{0}\right\rvert\lambda\sigma/\sqrt{2}.$
Notice that the tight upper bound in (87) (with matching lower bound) has an extra $\sqrt{\log p}$ factor compared to the $\ell_{1}$-sparsity condition in (80), where it is understood that $q^{*}\asymp s_{0}$ in order for the error $\left\lVert\beta_{\text{\rm init}}-\beta\right\rVert_{2}$ and $\left\lVert X(\beta_{\text{\rm init}}-\beta)\right\rVert_{2}$ in Theorem 3 [43] to match those in Theorem 2.4; cf. the supplementary (108). In summary, although we impose sparsity in the $\ell_{0}$ sense, the actual lower bound on $\left\lVert\beta_{A_{0}^{c}}\right\rVert_{1}$,
$\tilde{\eta}_{1}:=\left\lVert\beta_{A_{0}^{c}}\right\rVert_{1}>\left\lVert\beta_{T_{0}^{c}}\right\rVert_{1}=\Omega(\sqrt{\log p}\lambda\sigma(s_{0}-a_{0}))$
covers convergence results not available in  [43]. More explicitly, in (80), it is necessary for the $\ell_{1}$ norm on $\beta_{H_{0}}$ to satisfy
$\left\lVert\beta_{H_{0}}\right\rVert_{1}<\eta_{1}=\tilde{O}(r_{1}^{2}q\lambda\sigma)=O(1)(q^{*}\lambda\sigma)=O(1)(s_{0}\lambda\sigma),$
while in order for $\eta_{1}\geq\left\lVert\beta_{H_{0}}\right\rVert_{1}\geq\left\lVert\beta_{A_{0}^{c}}\right\rVert_{1}$ to hold, we need to set
$\;q^{*}\geq\left\lvert s_{0}-a_{0}\right\rvert\frac{c_{t}+\sqrt{\log p}}{\sqrt{2}c_{t}}\text{ so that }\eta_{1}\asymp q^{*}\lambda\sigma>\left\lVert\beta_{A_{0}^{c}}\right\rVert_{1};$ (88)
cf. (79) and (4.2). On the other hand, suppose $q^{*}\asymp\sqrt{\log p}(s_{0}-a_{0})$ is allowed, then this extra $\sqrt{\log p}$ factor in (88) will inevitably appear in the upper bounds derived in (86) and (85), resulting in worse $\ell_{2}$ loss while simultaneously, the bounded sparse eigenvalue conditions need to hold for design matrix $X$ with rank $q^{*}=\Omega((s_{0}-a_{0})\sqrt{\log p})$. See the supplementary Section 7.
$\ell_{1}$ error on $h$ versus on $\delta$. Recall $h=\beta_{\text{\rm init}}-\beta^{\mathop{\natural}\kern 0.60275pt}({T_{0}})$ and $\delta=\beta_{\text{\rm init}}-\beta$. Since $\mathop{\text{\rm supp}\kern 0.86108pt}(\beta_{\text{\rm init}})\asymp q^{*}$ by the supplementary (106), these results on the $\ell_{1}$ error $\left\lVert\delta\right\rVert_{1}$, where $\delta=\beta_{\text{\rm init}}-\beta$, and the Lasso support in [43], cf. (86) and the supplementary (106), are different from Theorem 2.4 in the present paper, as we do not aim to bound the Lasso support directly. A more subtle point is that since $q^{*}\asymp\left|A_{1}\right|$, we can crudely interpret $q^{*}$ as the size of $\mathop{\text{\rm supp}\kern 0.86108pt}(\beta_{\text{\rm init}})$, given $q\ll q^{*}$;
In our theory, indeed, the Lasso support can be much larger than $s_{0}$ while the $\ell_{1}$-, $\ell_{2}$-norm bounds on $h$ are quite tight in the sense of Theorem 2.4. Hence the SRC assumptions (83) are somewhat similar and closely related to the sparse eigenvalue conditions in Theorem 2.4, where $\textnormal{{RE}}(s_{0},4,X)$ also holds with $K(s_{0},4)<\infty$. On the one hand, the SRC condition is set at a rank $q^{*}\asymp\left|\mathop{\text{\rm supp}\kern 0.86108pt}(\beta_{\text{\rm init}})\right|$, which may be significantly larger than $s_{0}$ in our model. On the other hand, the $\ell_{2}$ and $\ell_{1}$ error on $h$ in Theorem 2.4 depend on $s_{0}$, as well as $K^{2}(s_{0},4)$, $\Lambda_{\max}(2s_{0})$, and $\Lambda_{\min}(2s_{0})$ (cf. (30)), and $\Lambda_{\max}(s-s_{0})\asymp\Lambda_{\max}(\log p(s_{0}-a_{0})/c_{t}^{2})$, without an explicit condition on (82).
We demonstrate the tightness of these theoretical bounds in Figures  2 and 3. Consider (69) in particular. In our experiments, $c_{t}$ is an absolute constant $\in[0.527,1.658]$ (See Table 2), and the signs of the non-zero values in $\beta^{(2)}$ are chosen from $\{\pm 1\}$ at random. Hence, we have a longer tail (in the sense of a larger $\left\lVert\beta^{(2)}\right\rVert_{1}$ with many small coordinates) and we expect the Lasso estimate to have a larger $\left\lVert\delta\right\rVert_{1}$. See Table 2, where we fix $\sigma=1$, $a_{0}=30$, $s_{0}=50$, and $C_{m}=1/\sqrt{2}$ for $s>s_{0}$.
As shown in the bottom two plots in Figure 2, where we set $(p,n,s_{0},a_{0},\gamma,C_{a})=(2048,1600,50,30,0.7,1.706)$, we observe that $\left\lVert\delta\right\rVert_{1}$ dominates $\left\lVert h\right\rVert_{1}$ consistently across all $s\in\{130,370,511\}$ as the Lasso penalty $\lambda_{n}=f_{p}\lambda\sigma$ varies. In particular while upon rescaling, all curves corresponding to different values of $s$ align well for the $\ell_{2}$-error, namely, $\left\lVert\delta\right\rVert_{2}$, $\left\lVert h_{T_{0}}\right\rVert_{2}$ (right plot) and the $\ell_{1}$-norm $\left\lVert h\right\rVert_{1}$, $\left\lVert h_{T_{0}}\right\rVert_{1}$, and $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ (left plot), the same is not true for $\left\lVert\delta_{T_{0}^{c}}\right\rVert_{1}$ and $\left\lVert\delta\right\rVert_{1}$. However, we have by the triangle inequality,
$\left\lvert\left\lVert\delta\right\rVert_{1}-\left\lVert h\right\rVert_{1}\right\rvert$ $=$ $\left\lvert\left\lVert\delta_{T_{0}^{c}}\right\rVert_{1}-\left\lVert h_{T_{0}^{c}}\right\rVert_{1}\right\rvert\leq\left\lVert\beta_{T_{0}^{c}}\right\rVert_{1}.$ (89)
We show the lower bound on $\left\lVert\beta_{T_{0}^{c}}\right\rVert_{1}$ in (79). Hence the error $\left\lVert\delta_{T_{0}^{c}}\right\rVert_{1}$ increases as $s$ increases (as $c_{t}$ decreases) but also remains bounded, as predicted by (74), (75) and (89). See Sections 5.1.
4.4 Variable selection in $A_{0}$
Note that $\beta_{\min,A_{0}}=\left\lVert\beta^{(11)}\right\rVert_{\infty}=C_{a}\lambda\sigma$, $\left\lVert\beta^{(12)}\right\rVert_{\infty}=C_{m}\lambda\sigma$, and $\left\lVert\beta^{(2)}\right\rVert_{\infty}=C_{t}\lambda\sigma=c_{t}\sigma/\sqrt{n}$. We mention up front that no matter where we put the threshold, some of the signals in $\beta^{(12)}$ will be lost so long as $t_{0}\asymp\lambda\sigma$ since $C_{m}\leq 1$; cf. Fig. 3. In our experiments, we can consistently recover those signals in $A_{0}$ for $\beta_{\min,A_{0}}:=\min_{j\in A_{0}}|\beta_{j}|\asymp\lambda\sigma\sqrt{s_{0}}$ (in case $C_{a}=8.528$ for the model class in Figure 1), but this is not the case when $\beta_{\min,A_{0}}\asymp\lambda\sigma$ (in case $C_{a}=1.706$). This difference occurs despite the nearly identical $\ell_{p},p=1,2$ norm bounds on the estimation error for the initial Lasso estimator $\beta_{\text{\rm init}}$ for the two models of $\beta$; See the top two panels in Figure 2, where curves corresponding to different values of $\beta_{\min,A_{0}}$ with $C_{a}\lambda\sigma=0.2,1$ align well across different values of $s\in\{130,370,511\}$; This is true for $\left\lVert h_{T_{0}}\right\rVert_{2}$, $\left\lVert h_{T_{0}}\right\rVert_{1}$, $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$, and $\left\lVert\delta\right\rVert_{2}$ under the same design matrix $X$ of dimension $1600\times 2048$. The relative effect on variable selection in component $\beta^{(12)},\beta^{(2)}$, and $\beta^{(0)}$ follows the same trend for both $C_{a}$ settings. The relative effect on variable selection in component $\beta^{(11)}$ is much more significant in case $\beta_{\min,A_{0}}=0.2$ ($C_{a}=1.706$), since the largest error magnitude may reach the full signal strength, while for $\beta_{\min,A_{0}}=1$ ($C_{a}=8.528$), the magnitude of the error is only a fraction $\propto 1/{\sqrt{s_{0}}}$ of the signal strength in $A_{0}$ and hence $A_{0}\subset I$ so long as $t_{0}=o(\lambda\sigma\sqrt{s_{0}})$.
Table 1: Evaluation metrics for variable selection. II is the estimated model, and ${\mathcal{D}}=I^{c}$.
Metric Definition Meaning
TP $I\cap T_{0}$ Selected variables in $T_{0}$
FP $I\cap T_{0}^{c}$ Selected variables in $T_{0}^{c}$
TN ${\mathcal{D}}\cap T_{0}^{c}$ Not selected variables in $T_{0}^{c}$
FN ${\mathcal{D}}\cap T_{0}$ Not selected variables in $T_{0}$
5 Numerical results
In this section, we present results from numerical simulations to validate the theoretical analysis presented in previous sections. We consider Gaussian random matrices for the design $X$ with both $p\times p$ identity and Toeplitz covariance. We refer to the former as i.i.d. Gaussian ensemble, where $X_{ij}\sim N(0,1)$ for all $i,j$, and the latter as Toeplitz ensemble, where the covariance matrix $T(\gamma)$ for each row vector in $\mathbb{R}^{p}$ is given by $T(\gamma)]_{i,j}=\gamma^{|i-j|}$. The design matrix dimensions are either $(p=1024,n=800)$ or $(p=2048,n=1600)$. To evaluate the impact of nominal sparsity $s$ on the recovery of $s_{0}$ components, we use $\beta$ as constructed as in Section 4.2. We have the following steps.
  1. 1.
    Generate input $\beta\in\mathbb{R}^{p}$ as shown in Fig. 1. $\beta$ is determined by the parameters $(C_{a},C_{m},C_{t})$, $\lambda=\sqrt{2\log p/n}$, and $\sigma$ in the noise $\epsilon$. Here, we fix $a_{0}=30,s_{0}=50$, and $\sigma=1$. The signs and positions of the non-zero coordinates are chosen at random. See Table 2.
  2. 2.
    Generate a Gaussian ensemble $X_{n\times p}$ with independent rows, which is then normalized to have column $\ell_{2}$-norm $\sqrt{n}$. We consider two types of design: i.i.d. Gaussian ensemble, and Toeplitz ensemble as mentioned above.
  3. 3.
    Compute $Y=X\beta+\epsilon$, where the noise $\epsilon\sim N(0,\sigma^{2}I_{n})$ is generated with $I_{n}$ being the $n\times n$ identity matrix.
  4. 4.
    Feed $Y$ and $X$ to the Thresholded Lasso algorithm to estimate $\beta$ as described in Section 1. We call the Lasso procedure $\textsf{LARS}(Y,X)$  [15] to compute the full regularization path. We select the $\beta_{\text{\rm init}}$ from this output path with penalty parameter $\lambda_{n}=f_{p}\lambda\sigma$. We then threshold $\beta_{\text{\rm init}}$ with threshold $t_{0}=f_{t}\lambda\sigma$, and run OLS to obtain $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)$.
We set $C_{m}=1$ for $s=s_{0}$ so that $\left\lVert\beta^{(12)}\right\rVert_{2}=(s_{0}-a_{0})\lambda\sigma$ and $\left\lVert\beta^{(2)}\right\rVert_{2}=0$. For $s>s_{0}$, we fix $C_{m}=1/\sqrt{2}$, and set $\left\lVert\beta^{(12)}\right\rVert_{2}=\left\lVert\beta^{(2)}\right\rVert_{2}=\frac{1}{\sqrt{2}}\lambda\sigma\sqrt{s_{0}-a_{0}}$ in (69) and (73). Hence the upper bound in (71) becomes an equality for both scenarios. When we lower $C_{t}$ ($c_{t}$) in (76), we will have a $\beta$ with a larger $\mathop{\text{\rm supp}\kern 0.86108pt}(\beta)$ with many small coefficients. In particular, $s-s_{0}$ (the length of $\beta^{(2)}$) and $\left\lVert\beta^{(2)}\right\rVert_{1}$ increase as $c_{t}$ decreases, but $\left\lVert\beta^{(2)}\right\rVert_{2}$ remains the same; cf. (73) and (74). There are two main tuning parameters: $\lambda_{n}=f_{p}\lambda\sigma$ and $t_{0}=f_{t}\lambda\sigma$, where $f_{p},f_{t}>0$. For each experiment, after we generate $\beta\in\mathbb{R}^{p}$ in Step 1, we repeat Steps 2-4 100 times, and compute averages after 100 runs. Due to limited space, we only present results from experiments using Toeplitz ensemble with $\gamma\in\{0.3,0.7\}$, but we observe similar trends for other design matrices such as the i.i.d. Gaussian ensemble. In the present context, we adopt a more stringent definition of metrics for variable selection evaluation. More specifically, we define True positives (TPs) as those variables from $I\cap T_{0}$, and False positives (FPs) refer to variables in $I\cap T_{0}^{c}$. Note that this interpretation naturally flags many more variables as FPs, which in the conventional notion would have been counted as TPs. False negatives (FNs) refer to variables from ${\mathcal{D}}\cap T_{0}$ where ${\mathcal{D}}=I^{c}$. True negatives (TNs) refer to variables from ${\mathcal{D}}\cap T_{0}^{c}$. See Table 1.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 2: $p=2048,n=1600$. Left column: $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$, $\left\lVert h_{T_{0}}\right\rVert_{1}$, and $\left\lVert h\right\rVert_{1}$ as Lasso penalty ($f_{p}$) increases across different sparsity $s\in\{130,370,511\}$. Right column: plots of $\left\lVert h_{T_{0}}\right\rVert_{2}$ and $\left\lVert\delta\right\rVert_{2}$. In the top panel, we fix $\gamma=0.3$, and compare two cases of $C_{a}\lambda\sigma\in\{0.2,1\}$. In the middle panel, we fix $C_{a}\lambda\sigma=1$ and compare two cases of $\gamma\in\{0.3,0.7\}$. In the bottom panel, we zoom in on one case with $\gamma=0.7,C_{a}\lambda\sigma=0.2$, and we plot $\left\lVert\delta\right\rVert_{1}$ together with $\left\lVert\delta\right\rVert_{2}$ in the bottom right panel.
Refer to caption Refer to caption
(a) (b)
Refer to caption Refer to caption
(c) (d)
Figure 3: $p=2048,n=1600,\gamma=0.7$. Plots of model size ($|I|$), number of TPs and FPs, as threshold increases. Note $|I|=$ TPs + FPs. In (a) and (b), Lasso penalty factor $f_{p}=0.3$ is fixed, and in panel (a) $s\in\{130,511,710\}$, and in panel (b) $s\in\{50,130\}$. In panels (c) and (d), we plot the same metrics across different $f_{p}\in\{0.1,0.3,0.7\}$ with fixed $s=130$. In all panels, the 3 dotted vertical lines from left to right represent $C_{m}\lambda\sigma/2,C_{m}\lambda\sigma$ and $\lambda\sigma$. The model size remains invariant and hence the diagonal dashed lines all stay flat for $\lambda\sigma<t_{0}\leq 2\lambda\sigma$ for $\beta_{\min,A_{0}}=1$.
5.1 $\ell_{1}$ and $\ell_{2}$ error bounds for $\beta_{\text{\rm init}}$
Denote by $\beta_{A}$ the restriction of $\beta$ to the set $A\subset[p]$, with all other coordinates set to zero. We also use $\beta^{(1)}:=\beta^{(11)}+\beta^{(12)}=\beta_{T_{0}}$ and $\beta^{(2)}+\beta^{(0)}=\beta_{T_{0}^{c}}$ throughout our discussion. Recall $h=\beta_{\text{\rm init}}-\beta_{T_{0}}$ and $\delta=\beta_{\text{\rm init}}-\beta$. First, we investigate $\ell_{1}$ and $\ell_{2}$ error bounds for $\delta$ and $h$, with the Lasso penalty being $\lambda_{n}=f_{p}\lambda\sigma$. In our experiments, we vary the correlation parameter $\gamma$, $\beta_{\min,A_{0}}$ and nominal sparsity $s$ under model class (76).
In the top and middle rows of Fig. 2, we plot $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$, $\left\lVert h_{T_{0}}\right\rVert_{1}$, and $\left\lVert h\right\rVert_{1}$ in the left column, and $\left\lVert h_{T_{0}}\right\rVert_{2}$ and $\left\lVert\delta\right\rVert_{2}$ in the right column. Across all plots, we observe that curves for both $\left\lVert h_{T_{0}}\right\rVert_{1}$ and $\left\lVert h_{T_{0}}\right\rVert_{2}$ decrease slightly first and then increase as $f_{p}$ increases, due to some variables with significant coefficients being eliminated. We observe that as $f_{p}$ increases, $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ decreases quickly for all nominal sparsity $s$ ($y$-axis is in log scale). Even though values in $\beta^{(2)}$ are non-zero, due to their small magnitude, they are essentially treated the same way as the zeros in $\beta^{(0)}$ (as they should be). Moreover, all curves for $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ and $\left\lVert h\right\rVert_{1}$ align well for the same $\gamma\in\{0.3,0.7\}$ as predicted by Lemma 3.3 and Theorem 2.4 since we fix $s_{0}$ while varying sparsity $s\in\{130,370,511\}$.
Dependence on $\gamma$. In the top row, all curves are closely aligned, validating that $\beta_{\min,A_{0}}$ (for different values of $C_{a}$) does not impact these error metrics significantly. In contrast, in the middle row, we observe that curves for $\left\lVert h_{T_{0}}\right\rVert_{1}$ and $\left\lVert h_{T_{0}}\right\rVert_{2}$, shift downwards slightly for $\gamma=0.3$. This is expected since for the design matrix $X$ with smaller $\gamma$, the incoherence parameters appearing in Theorem 2.4 will be smaller. The two sets of curves for $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ and $\left\lVert h\right\rVert_{1}$ cross each other for $\gamma=0.3$ and $0.7$ with a small gap (middle left panel), but eventually the set with a larger $\gamma$ stays on the top. The gap is potentially caused by the non-linear interactions between $\gamma$ and the penalty parameters throughout the entire Lasso path.
$\ell_{1}$ and $\ell_{2}$ error for $\delta$. In the right column, for the $\ell_{2}$ error for estimating $\beta$ with $\beta_{\text{\rm init}}$, we observe the typical V-shaped curves as $f_{p}$ increases from 0 to $1$, since $\left\lVert\delta\right\rVert_{2}$ reaches a minimum and then increases again as the penalty $\lambda_{n}$ increases. In the bottom right panel of Fig. 2, we plot $\left\lVert\delta\right\rVert_{1}$ and $\left\lVert\delta\right\rVert_{2}$, and also $\left\lVert h\right\rVert_{1}$ (middle solid curves) for reference. All solid curves in the bottom corresponding to the same pair of $(s_{0},\gamma)$ again align well as predicted by Theorem 2.4 (bottom right panel). In the same panel, we observe the three dashed curves corresponding to $\left\lVert\delta\right\rVert_{1}$ are clearly separated under different sparsity $s\in\{130,370,511\}$, and stacked in descending order as $s$ decreases, that is, when $\beta$ becomes more sparse in (89). We know $\left\lVert\delta\right\rVert_{1}=\left\lVert\delta_{T_{0}}\right\rVert_{1}+\left\lVert\delta_{T_{0}^{c}}\right\rVert_{1}$, where $\left\lVert\delta_{T_{0}}\right\rVert_{1}=\left\lVert h_{T_{0}}\right\rVert_{1}$. In contrast, as shown in the left panels, $\left\lVert h_{T_{0}}\right\rVert_{1}$, $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ and hence $\left\lVert h\right\rVert_{1}$ all align well across different $s$ for the same $\gamma$. Hence, the difference in $\left\lVert\delta\right\rVert_{1}$ is due to the component $\left\lVert\delta_{T_{0}^{c}}\right\rVert_{1}$, which depends on the sparsity. This phenomenon is expected and explained in Section 4.3. The influence of $\gamma$ on $\left\lVert\delta\right\rVert_{2}$ follows a similar trend as that for $\left\lVert h_{T_{0}}\right\rVert_{2}$. Hence, all error curves corresponding to a larger $\gamma$ ($0.7$) consistently dominate those with a smaller $\gamma=0.3$ for $\left\lVert h_{T_{0}}\right\rVert_{1}$, $\left\lVert h_{T_{0}}\right\rVert_{2}$, and $\left\lVert\delta\right\rVert_{2}$ in the middle panel.
5.2 Variable selection with thresholding
In Figure 3, we plot in all panels the final model size $\left\lvert I\right\rvert$ (top right solid black curves), the number of TPs in model II ($\left|I\cap T_{0}\right|$, middle diagonal red dashed lines/curves), and the number of FPs from $T_{0}^{c}$, ($\left|I\cap T_{0}^{c}\right|$, left bottom dotted green curves), as functions of threshold $t_{0}$ for $\beta_{\min,A_{0}}=C_{a}\lambda\sigma\in\{0.2,1\}$ and $C_{a}\in\{1.706,8.528\}$. In panel (a), all curves align well across different $s\in\{130,370,511\}$. At $t_{0}=C_{m}\lambda\sigma$ (middle dotted vertical line), the model size is only slightly above $\left\lvert A_{0}\right\rvert=30$ for both cases of $C_{a}$. When the top curves (plotting $\left|I\right|$) touch upon or cross over the horizontal line of $y=30$ at $t_{0}=t^{\prime}$, where $t^{\prime}\in(\lambda\sigma/2,\lambda\sigma)$, the model contains $A_{0}$ exactly. For $\beta_{\min,A_{0}}=1$, all coordinates in $\beta^{(11)}$ remain in model $I$ so long as $t_{0}\leq 2\lambda\sigma$. For $\beta_{\min,A_{0}}=0.2$ ($C_{a}=1.706$), the model size $\left\lvert I\right\rvert$ continues to shrink till 0, as $t_{0}$ increases.
In Fig.  3 panel (b), we compare the exact $s_{0}$-sparse case ($s=50$) with the almost $s_{0}$-sparse case ($s=130$). Under the same $t_{0}$, we observe that the exact sparse case recovers more non-zero components from $T_{0}$ (higher red dashed lines with more TPs and fewer FNs for $t_{0}<1.5\lambda\sigma$) and less from $T_{0}^{c}$ (lower green lines with fewer FPs for $t_{0}<\lambda\sigma$), since for $s=s_{0}$, $\beta_{j}^{(12)}$ is larger in magnitude with $C_{m}=1$ and $\beta_{T_{0}^{c}}=0$. Panels (c) and (d) show that as the Lasso penalty ($f_{p}$) increases, model sizes further decrease, and hence $\left\lvert I\cap T_{0}\right\rvert$ and $\left\lvert I\cap T_{0}^{c}\right\rvert$ both decrease. Hence all three sets of curves ($|I|$, TPs, FPs) shift downward as $f_{p}$ increases since Lasso is able to remove some less significant variables as an initial estimator. However, FPs remain at a high level without thresholding or when the threshold is small. This is true for both $C_{a}$ settings.
False negatives. Recall FNs = $s_{0}$ - TPs. The primary distinction between the two settings of $C_{a}$ lies in the FNs from $\beta^{(11)}$, since we will lose some variables from $\beta^{(12)}$ inevitably for both choices of $C_{a}$, no matter where we put $t_{0}$, as $\beta^{(12)}_{\text{\rm init},j}$ may fall within the range of $\pm c\lambda\sigma$ for any $c\in(0,1]$. The entries of the Lasso estimate $\beta_{\text{\rm init}}^{(12)}$ of $\beta^{(12)}$ are indeed spread across the interval of $\lambda\sigma,\lambda\sigma]$ since $\left\lvert\beta_{\text{\rm init},j}^{(12)}\right\rvert\leq\lambda\sigma$. This is indicated in Fig. 3 by the negative slope in the red diagonal lines, where coordinates in $\beta^{(12)}$ are regularly cut as $t_{0}$ increases. Larger $\beta_{\min,A_{0}}$ means variables in $A_{0}$ will be kept over a longer range of $t_{0}$, as the remaining TPs are all from $A_{0}$ after a certain threshold. Indeed, for $\beta_{\min,A_{0}}=1$ ($C_{a}=8.528$), we observe in Fig. 3 panel (c), the curves for TPs (red dashed slanted lines) have a changing slope at around $t_{0}=t^{\prime}$, where $t^{\prime}\in(\lambda\sigma/2,\lambda\sigma]$ for $f_{p}=0.3$, and then flattens out along the horizontal line of $y=30$ until $t_{0}=4\lambda\sigma$. For $\beta_{\min,A_{0}}=0.2$ ($C_{a}=1.706$, panel (d)), the dashed diagonal line intersects the horizontal line of $y=30$, and continues with the downward trend until it reaches $y=0$ while losing all true variables.
False positives. We observe in Fig. 3 panel (a) that FPs drop sharply in both $C_{a}$ cases, and the rate is the same for all $s\in\{130,511,770\}$ as $t_{0}$ increases, whereas TPs drop with a slower rate due to their larger estimated values. At $t_{0}=\lambda\sigma/2$, the model size is about but slightly below $50$ with both FNs and FPs. By our theory, the coordinates in $\beta^{(2)}$ with small coefficients are at the noise level and hence are neither guaranteed nor necessary to be included in the model $I$. Roughly speaking, the largest magnitude of the Lasso estimate $\beta_{\text{\rm init}}^{(2)}$ and $\beta_{\text{\rm init}}^{(0)}$ as well as their $\ell_{\infty}$ norm error are nearly all bounded by $\lambda\sigma$ in absolute values, and hence $t_{0}\asymp\lambda\sigma$ is effective in controlling the number of variables selected from $T_{0}^{c}$ (False Positives, cf. Table 1), as shown on the left bottom dotted green curves on all panels in Figure 3. On each curve, the number of FPs drops quickly as $t_{0}$ increases and transitions to a horizontal line at $t_{0}\approx 0.8\lambda\sigma$, as predicted by the lower bound on $t_{0}$ in (66).
$\ell_{2}$-norm error of $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)$. In this section, we show that the previously stated tradeoffs between FPs and TPs in Thresholded Lasso does not come at the cost of an increased $\ell_{2}$ error of the final estimator. We plot in Fig. 4, the $\ell_{2}$-norm error of the final estimate, defined as $\left\lVert\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)-\beta\right\rVert_{2}$, we observe that for all the cases of Lasso penalty ($f_{p}$ values), for a wide range of $t_{0}$, $\left\lVert\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)-\beta\right\rVert_{2}$ stays at the same level or below $\left\lVert\delta\right\rVert_{2}$. This is as predicted by Theorems 2.1 and 2.4. For lower $f_{p}$ values, such as $0.1,0.3$, $\left\lVert\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)-\beta\right\rVert_{2}$ further decreases as threshold $t_{0}$ increases. This is because small coordinates in $\beta_{\text{\rm init}}$ with values below $t_{0}$ will be gone with thresholding, while variables in $A_{0}$ remain intact due to their larger magnitudes. When $f_{p}$ increases from $0.3$ to $1$, $\left\lVert\delta\right\rVert_{2}$ has a V-shaped curve, due to the loss of significant coordinates in $T_{0}$. This is shown in Fig. 4, where the horizontal lines corresponding to $f_{p}=1$ is higher than the ones correspond to $f_{p}=0.5$ or $0.7$; See also Fig. 2.
s $C_{a}$ $C_{a}\lambda\sigma$ $C_{m}$ $C_{m}\lambda\sigma$ $C_{t}$ $C_{t}\lambda\sigma$ $c_{t}$ $\left\lVert\beta^{(11)}\right\rVert_{1}$ $\left\lVert\beta^{(12)}\right\rVert_{1}$ $\left\lVert\beta^{(2)}\right\rVert_{1}$ $\left\lVert\beta^{(11)}\right\rVert_{2}$ $\left\lVert\beta^{(12)}\right\rVert_{2}$ $\left\lVert\beta^{(2)}\right\rVert_{2}$
$p=1024,n=800,\lambda\sigma=0.158$
50 6.325 1.0 1.000 0.158 0.000 0.000 0.000 30 3.162 0.000 5.477 0.707 0.000
130 6.325 1.0 0.707 0.112 0.354 0.056 1.581 30 2.236 4.472 5.477 0.500 0.500
370 6.325 1.0 0.707 0.112 0.177 0.028 0.791 30 2.236 8.944 5.477 0.500 0.500
511 6.325 1.0 0.707 0.112 0.147 0.023 0.659 30 2.236 10.738 5.477 0.500 0.500
770 6.325 1.0 0.707 0.112 0.118 0.019 0.527 30 2.236 13.416 5.477 0.500 0.500
50 1.265 0.2 1.000 0.158 0.000 0.000 0.000 6 3.162 0.000 1.095 0.707 0.000
130 1.265 0.2 0.707 0.112 0.354 0.056 1.581 6 2.236 4.472 1.095 0.500 0.500
370 1.265 0.2 0.707 0.112 0.177 0.028 0.791 6 2.236 8.944 1.095 0.500 0.500
511 1.265 0.2 0.707 0.112 0.147 0.023 0.659 6 2.236 10.738 1.095 0.500 0.500
770 1.265 0.2 0.707 0.112 0.118 0.019 0.527 6 2.236 13.416 1.095 0.500 0.500
$p=2048,n=1600,\lambda\sigma=0.117$
50 8.528 1.0 1.000 0.117 0.000 0.000 0.000 30 2.345 0.000 5.477 0.524 0.000
130 8.528 1.0 0.707 0.083 0.354 0.041 1.658 30 1.658 3.317 5.477 0.371 0.371
370 8.528 1.0 0.707 0.083 0.177 0.021 0.829 30 1.658 6.633 5.477 0.371 0.371
511 8.528 1.0 0.707 0.083 0.147 0.017 0.691 30 1.658 7.963 5.477 0.371 0.371
770 8.528 1.0 0.707 0.083 0.118 0.014 0.553 30 1.658 9.950 5.477 0.371 0.371
50 1.706 0.2 1.000 0.117 0.000 0.000 0.000 6 2.345 0.000 1.095 0.524 0.000
130 1.706 0.2 0.707 0.083 0.354 0.041 1.658 6 1.658 3.317 1.095 0.371 0.371
370 1.706 0.2 0.707 0.083 0.177 0.021 0.829 6 1.658 6.633 1.095 0.371 0.371
511 1.706 0.2 0.707 0.083 0.147 0.017 0.691 6 1.658 7.963 1.095 0.371 0.371
770 1.706 0.2 0.707 0.083 0.118 0.014 0.553 6 1.658 9.950 1.095 0.371 0.371
Table 2: In this table, we list the actual values of $s$ generated through (76) for $\beta$ configurations with $a_{0}=30,s_{0}=50.$ We also list magnitudes of each $\beta$ component and their $\ell_{1}$ and $\ell_{2}$ norms. Here $c_{t}=C_{t}\sqrt{2\log p}$. The component $\beta^{(11)}=\beta_{A_{0}}$ has $a_{0}=30$ non-zero coordinates with magnitude $C_{a}\lambda\sigma$, where $C_{a}>1$. The component $\beta^{(12)}=\beta_{T_{0}\setminus A_{0}}$ has $s_{0}-a_{0}=20$ non-zero coordinates with magnitude $C_{m}\lambda\sigma$, where $C_{m}=1/{\sqrt{2}}$ for $s>s_{0}$. The component $\beta^{(2)}=\beta_{S\setminus T_{0}}$ consists of $s-s_{0}$ non-zero coordinates of magnitude $c_{t}\sigma/\sqrt{n}$, where $c_{t}$ is an absolute constant.
Refer to caption Refer to caption
Figure 4: $p=2048,n=1600$, $s=130$. Plots of $\left\lVert\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.54248pt}}(I)-\beta\right\rVert_{2}$, for $C_{a}\lambda\sigma\in\{0.2,1\}$ and $\gamma\in\{0.3,0.7\}$. The horizontal lines correspond to the $\ell_{2}$-norm error of Lasso estimate $\beta_{\text{\rm init}}$, namely, $\left\lVert\delta\right\rVert_{2}$.
In practice, we recommend using cross-validation to set the Lasso penalty $\lambda_{n}$, which typically ranges between $0.3\lambda\sigma$ and $0.7\lambda\sigma$ for this example, and then apply thresholding.
6 Proofs
6.1 Proof of Theorem 2.1
It holds by definition of $S_{{\mathcal{D}}}$ that $I\cap S_{{\mathcal{D}}}=\emptyset$. One can check via the proof of Theorem 2.4 that (19) holds for $D_{0}^{\prime},D_{1}$ as defined in (26) and (27) respectively. Hence by Lemma 2.3, we have on event ${\mathcal{T}}_{\alpha}$ $|I|\leq 2s_{0}$ and $|I\cup S_{{\mathcal{D}}}|\leq s+s_{0}\leq 2s$ for $C_{4}\geq D_{1}$; See also (90). Moreover, $\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}\leq\sqrt{(D^{\prime}_{0}+C_{4})^{2}+1}\lambda\sigma\sqrt{s_{0}}.$ We have by Lemma 2.7, on event ${\mathcal{T}}_{\alpha}$, for $\lambda=\sqrt{2\log p/n}$, $\left\lvert I\right\rvert<2s_{0}$, for $\hat{\beta}=\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)$,
$\left\lVert\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)-\beta\right\rVert_{2}^{2}$ $\leq$ $\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}^{2}\left(1+\frac{2\theta^{2}_{|I|,|{S_{{\mathcal{D}}}}|}}{\Lambda_{\min}^{2}(|I|)}\right)+\frac{2|I|(1+a)\sigma^{2}\lambda^{2}}{\Lambda^{2}_{\min}(|I|)}$
$\leq$ $\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}^{2}\big(1+\frac{2\theta^{2}_{\left\lvert S_{{\mathcal{D}}}\right\rvert,|I|}}{\Lambda_{\min}^{2}(2s_{0})}\big)+\frac{4(1+a)}{\Lambda_{\min}^{2}(2s_{0})}\sigma^{2}\lambda^{2}s_{0}$
$\leq$ $\lambda^{2}\sigma^{2}s_{0}((D^{\prime}_{0}+C_{4})^{2}+1)\big(1+\frac{2\theta^{2}_{\left\lvert S_{{\mathcal{D}}}\right\rvert,|I|}}{\Lambda_{\min}^{2}(2s_{0})}+\frac{4}{9}\big),$
where $\Lambda_{\min}(|I|)\geq\Lambda_{\min}(2s_{0})$ by (31) since $\left\lvert I\right\rvert<2s_{0}$, and the fact that ${4(1+a)}/{\Lambda_{\min}^{2}(2s_{0})}\leq\frac{4}{9}(D^{\prime}_{0})^{2}$ since
$(D^{\prime}_{0})^{2}$ $\geq$ $9d^{2}_{0}K^{4}(s_{0},4)\geq{9(1+a)}/{\Lambda^{2}_{\min}(2s_{0})},\quad\text{ where}$
[…]
Now (17) clearly holds with $D_{4}^{2}=((D^{\prime}_{0}+C_{4})^{2}+1)\big(1+\frac{2\theta_{|I|,|{S_{{\mathcal{D}}}}|}^{2}}{\Lambda_{\min}^{2}(2s_{0})}+\frac{4}{9}\big)$, and (18) holds by Lemma 2.8. ∎
6.2 Proof of Lemma 2.3
Without loss of generality, we order $\left\lvert\beta_{1}\right\rvert\geq\left\lvert\beta_{2}\right\rvert\geq\ldots\geq\left\lvert\beta_{p}\right\rvert$. Then $T_{0}=\{1,\ldots,s_{0}\}$. Let $T_{1}$ be the largest $s_{0}$ positions of $\beta_{\text{\rm init}}$ outside of $T_{0}$. Then
$\left|I\cap T^{c}_{0}\right|$ $\leq$ $\frac{\left\lVert\beta_{\text{\rm init},T^{c}_{0}}\right\rVert_{1}}{(C_{4}\lambda\sigma)}=\frac{\left\lVert h_{T^{c}_{0}}\right\rVert_{1}}{(C_{4}\lambda\sigma)}\leq s_{0}D_{1}/C_{4}.$
Thus $|I|=|I\cap T_{0}|+|I\cap T_{0}^{c}|\leq s_{0}+s_{0}D_{1}/C_{4}$; Now  (20) holds since $T_{0}\subseteq S$ and hence
$|I\cup S|=|S|+|I\cap S^{c}|\leq s+|I\cap T_{0}^{c}|\leq s+s_{0}D_{1}/C_{4}.$ (90)
We now bound $\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}^{2}$. Denote by
$\beta_{j}^{(1)}\;=\;\beta_{j}\cdot 1_{j\leq s_{0}}\;\;\text{ and}\;\beta_{j}^{(2)}\;=\;\beta_{j}\cdot 1_{j>s_{0}}.$
Let $\beta_{{\mathcal{D}}}=\beta_{{\mathcal{D}}}^{(1)}+\beta_{{\mathcal{D}}}^{(2)}$, where $\beta_{{\mathcal{D}}}^{(1)}:=(\beta_{j})_{j\in T_{0}\cap{\mathcal{D}}}$ consists of coefficients that are significant relative to $\lambda\sigma$, but are dropped as $\left\lvert\beta_{j,\text{\rm init}}\right\rvert<t_{0}$, and $\beta_{{\mathcal{D}}}^{(2)}$ consists of those strictly below $\lambda\sigma$ in magnitude that are dropped; cf. (14). Hence
$\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}^{2}=\left\lVert\beta_{{\mathcal{D}}}^{(1)}\right\rVert_{2}^{2}+\left\lVert\beta_{{\mathcal{D}}}^{(2)}\right\rVert_{2}^{2}.$ (91)
Now it is clear $\beta_{{\mathcal{D}}}^{(2)}$ is bounded given (14), indeed, we have for $\lambda=\sqrt{2\log p/n}$,
$\left\lVert\beta_{{\mathcal{D}}}^{(2)}\right\rVert_{2}^{2}\leq\left\lVert\beta^{(2)}\right\rVert_{2}^{2}=\sum_{j>s_{0}}\beta_{j}^{2}=\sum_{j>s_{0}}\min(\beta_{j}^{2},\lambda^{2}\sigma^{2})\leq s_{0}\lambda^{2}\sigma^{2},$ (92)
where the second equality is by (14) and the last inequality is by definition of $s_{0}$ as in (13). Let ${\mathcal{D}}_{1}={\mathcal{D}}\cap T_{0}$, where $|{\mathcal{D}}_{1}|<s_{0}$. Now by the triangle inequality,
$\left\lVert\beta_{{\mathcal{D}}}^{(1)}\right\rVert_{2}=\left\lVert\beta_{{\mathcal{D}}_{1}}\right\rVert_{2}$ $\leq$ $\left\lVert\beta_{{\mathcal{D}}_{1},\text{\rm init}}\right\rVert_{2}+\left\lVert\beta_{{\mathcal{D}}_{1},\text{\rm init}}-\beta_{{\mathcal{D}}_{1}}\right\rVert_{2}$ (93)
$\leq$ $t_{0}\sqrt{\left|{\mathcal{D}}_{1}\right|}+\left\lVert h_{T_{0}}\right\rVert_{2}\leq(C_{4}+D_{0}^{\prime})\lambda\sigma\sqrt{s_{0}},$
where we used the fact that $\left\lVert\beta_{{\mathcal{D}}_{1},\text{\rm init}}-\beta_{{\mathcal{D}}_{1}}\right\rVert_{2}=\left\lVert h_{{\mathcal{D}}_{1}}\right\rVert_{2}\leq\left\lVert h_{T_{0}}\right\rVert_{2}\leq D^{\prime}_{0}\lambda\sigma\sqrt{s_{0}}$ by (19). Hence (21) holds given (91), (92), and (93). ∎
6.3 Proof of Proposition 4.1
Recall that $|\beta_{j}|\leq\lambda\sigma$ for all $j>a_{0}$ as defined in (15); hence for $\lambda=\sqrt{2\log p/n}$, we have by (104), $\sum_{i>a_{0}}^{p}\min(\beta_{i}^{2},\lambda^{2}\sigma^{2})=\sum_{i>a_{0}}^{s}\beta_{i}^{2}\leq(s_{0}-a_{0})\lambda^{2}\sigma^{2}$; hence
$\left|\{j\in A_{0}^{c}:|\beta_{j}|\geq\sqrt{\log p/(c^{\prime}n)}\sigma\}\right|$ $\leq$ $2c^{\prime}(s_{0}-a_{0})\text{ where }\left|T_{0}\setminus A_{0}\right|=s_{0}-a_{0}.$
Now given that $\left\lvert\beta_{i}\right\rvert\geq\left\lvert\beta_{j}\right\rvert$ for all $i\in T_{0},j\in T_{0}^{c}$, the proposition holds. ∎
6.4 Proof of Lemma 2.5
Recall that the random variable $\left\lVert Q\right\rVert_{2}^{2}\sim\chi_{m}^{2}$ is distributed according to the chi-square distribution where $\left\lVert Q\right\rVert_{2}^{2}=\sum_{i=1}^{m}Q_{i}^{2}$ with $Q_{i}\sim N(0,1)$ that are independent and normally distributed. By [20],
${\mathbb{P}}\left(\frac{\chi^{2}_{m}}{m}-1\leq-\varepsilon\right)$ $\leq$ $\exp(-m\varepsilon^{2}/4)\;\;\text{ for}\;\;0\leq\varepsilon\leq 1,$ (94)
${\mathbb{P}}\left(\frac{\chi^{2}_{m}}{m}-1\geq\varepsilon\right)$ $\leq$ $\exp(-3m\varepsilon^{2}/(16))\;\;\text{ for}\;\;0\leq\varepsilon\leq 1/2.$ (95)
Although we need to bound the bad event only on one side, we provide a tight bound on the norm of $\left\lVert Q\right\rVert_{2}$ with (94) and (95). Let $\left\lvert I\right\rvert=m$. Thus we have for $Q=(Q_{1},\ldots,Q_{m})$ where $Q_{j}\sim\text{i.i.d}\;N(0,1)$, for $\delta<1/2$,
${\mathbb{P}}\left(\left\lvert\frac{1}{m}\sum_{j=1}^{m}Q_{j}^{2}-1\right\rvert\geq\delta\right)=:{\mathbb{P}}\left(\mathcal{Q}\right)=$
${\mathbb{P}}\left(\frac{1}{m}\sum_{j=1}^{m}Q_{j}^{2}-1\geq\delta\right)+{\mathbb{P}}\left(\frac{1}{m}\sum_{i=1}^{m}Q_{j}^{2}-1\leq\delta\right)$
$=$ ${\mathbb{P}}\left(\frac{\chi^{2}_{m}}{m}-1\geq\delta\right)+{\mathbb{P}}\left(\frac{\chi^{2}_{m}}{m}-1\leq-\delta\right)\leq\exp(-3m\delta^{2}/(16))+\exp(-m\delta^{2}/4).$
Note that $X_{I^{c}}\beta_{I^{c}}=X_{{S_{{\mathcal{D}}}}}\beta_{{S_{{\mathcal{D}}}}}$. We have
$\hat{\beta}_{I}=(X_{I}^{T}X_{I})^{-1}X_{I}^{T}Y=(X_{I}^{T}X_{I})^{-1}X_{I}^{T}(X_{I}\beta_{I}+X_{I^{c}}\beta_{I^{c}}+\epsilon)$
$\;\;\;\;=\beta_{I}+(X_{I}^{T}X_{I})^{-1}X_{I}^{T}X_{{S_{{\mathcal{D}}}}}\beta_{{S_{{\mathcal{D}}}}}+(X_{I}^{T}X_{I})^{-1}X_{I}^{T}\epsilon;$
Hence by the triangle inequality,
$\left\lVert\hat{\beta}_{I}-\beta_{I}\right\rVert_{2}$ $\leq$ $\left\lVert(X_{I}^{T}X_{I})^{-1}X_{I}^{T}X_{{S_{{\mathcal{D}}}}}\beta_{{S_{{\mathcal{D}}}}}+(X_{I}^{T}X_{I})^{-1}X_{I}^{T}\epsilon\right\rVert_{2}$ (96)
$\leq$ $\left\lVert(X_{I}^{T}X_{I})^{-1}X_{I}^{T}X_{{S_{{\mathcal{D}}}}}\beta_{{S_{{\mathcal{D}}}}}\right\rVert_{2}+\left\lVert(X_{I}^{T}X_{I})^{-1}X_{I}^{T}\epsilon\right\rVert_{2},$
where the two terms are bounded below as follows.
First notice that $w/\sigma=(X_{I}^{T}X_{I})^{-1}X_{I}^{T}\epsilon/\sigma$ is a mean zero Gaussian random vector with covariance $(X_{I}^{T}X_{I})^{-1}$, since ${\epsilon_{i}}/{\sigma}\sim N(0,1)$ and
$\frac{1}{\sigma^{2}}{\textbf{E}}(ww^{T})$ $=$ ${\textbf{E}}\left((X_{I}^{T}X_{I})^{-1}X_{I}^{T}[(\epsilon/\sigma)\otimes\epsilon/\sigma]X_{I}(X_{I}^{T}X_{I})^{-1}\right)$
$=$ $(X_{I}^{T}X_{I})^{-1}X_{I}^{T}\frac{1}{\sigma^{2}}{\textbf{E}}(\epsilon\epsilon^{T})X_{I}(X_{I}^{T}X_{I})^{-1}=(X_{I}^{T}X_{I})^{-1}.$
Then on event $\mathcal{Q}^{c}$, which holds with probability at least $1-2\exp(-3m/64)$
$\left\lVert(X_{I}^{T}X_{I})^{-1}X_{I}^{T}\epsilon/\sigma\right\rVert_{2}^{2}$ $=$ $Q^{T}(X_{I}^{T}X_{I})^{-1}Q\leq\Lambda_{\max}\left((X_{I}^{T}X_{I})^{-1}\right)\left\lVert Q\right\rVert_{2}^{2}$ (97)
$\leq$ $\frac{3m}{2\Lambda_{\min}\left(X_{I}^{T}X_{I}\right)}\leq\frac{3m}{2n\Lambda_{\min}(\left\lvert I\right\rvert)}$, (98)
where we used an upper bound on $\left\lVert Q\right\rVert_{2}^{2}\leq 3m/2$, and the fact that
$\Lambda_{\max}\left((X_{I}^{T}X_{I})^{-1}\right)=\frac{1}{\Lambda_{\min}\left((X_{I}^{T}X_{I})\right)}=\frac{1}{n\Lambda_{\min}\left((X_{I}^{T}X_{I})/n\right)}\leq\frac{1}{n\Lambda_{\min}(\left\lvert I\right\rvert)}$.
Here $\left\lVert Q\right\rVert_{2}^{2}\sim\chi^{2}_{m}$. We now focus on bounding the first term in (96). Let $P_{I}$ denote the orthogonal projection onto $I$. Clearly, $I\cap S_{{\mathcal{D}}}=\emptyset$. Let
$c=(X_{I}^{T}X_{I})^{-1}X_{I}^{T}X_{{S_{{\mathcal{D}}}}}\beta_{{S_{{\mathcal{D}}}}}\quad\quad\text{ and }\quad X_{I}c=P_{I}X_{{S_{{\mathcal{D}}}}}\beta_{{S_{{\mathcal{D}}}}}.$
By the disjointness of $I$ and $S_{{\mathcal{D}}}$}, we have for $P_{I}X_{{S_{{\mathcal{D}}}}}\beta_{{S_{{\mathcal{D}}}}}:=X_{I}c$,
$\left\lVert P_{I}X_{{S_{{\mathcal{D}}}}}\beta_{{S_{{\mathcal{D}}}}}\right\rVert_{2}^{2}$ $=$ $\;\langle{\,P_{I}X_{{S_{{\mathcal{D}}}}}\beta_{{S_{{\mathcal{D}}}}},X_{{S_{{\mathcal{D}}}}}{\beta_{{S_{{\mathcal{D}}}}}}\,}\rangle\;=\;\langle{\,X_{I}c,X_{{S_{{\mathcal{D}}}}}{\beta_{{S_{{\mathcal{D}}}}}}\,}\rangle\;$
$\leq$ $n\theta_{|I|,|{S_{{\mathcal{D}}}}|}\left\lVert c\right\rVert_{2}\left\lVert\beta_{{S_{{\mathcal{D}}}}}\right\rVert_{2}\leq n\theta_{|I|,|{S_{{\mathcal{D}}}}|}\left\lVert\beta_{{S_{{\mathcal{D}}}}}\right\rVert_{2}\frac{\left\lVert P_{I}X_{{S_{{\mathcal{D}}}}}\beta_{{S_{{\mathcal{D}}}}}\right\rVert_{2}}{\sqrt{n\Lambda_{\min}(|I|)}},$
 $\;\text{ where }\;\;\left\lVert c\right\rVert_{2}$ $\leq$ $\frac{\left\lVert X_{I}c\right\rVert_{2}}{\sqrt{n\Lambda_{\min}(|I|)}}\leq\frac{\left\lVert P_{I}X_{{S_{{\mathcal{D}}}}}\beta_{{S_{{\mathcal{D}}}}}\right\rVert_{2}}{\sqrt{n\Lambda_{\min}(|I|)}}.$ (99)
Hence
$\left\lVert P_{I}X_{{S_{{\mathcal{D}}}}}\beta_{{S_{{\mathcal{D}}}}}\right\rVert_{2}$ $\leq$ $\frac{\sqrt{n}\theta_{|I|,|{S_{{\mathcal{D}}}}|}}{\sqrt{\Lambda_{\min}(|I|)}}\left\lVert\beta_{{S_{{\mathcal{D}}}}}\right\rVert_{2},\quad\text{ where }\quad\left\lVert\beta_{{S_{{\mathcal{D}}}}}\right\rVert_{2}=\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}$
$\text{ and }\quad\left\lVert c\right\rVert_{2}$ $\leq$ ${\theta_{|I|,|{S_{{\mathcal{D}}}}|}\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}}/{\Lambda_{\min}(|I|)}.$ (100)
Now we have on ${\mathcal{T}}_{a}$, by (97),
$\left\lVert\hat{\beta}_{I}-\beta_{I}\right\rVert_{2}$ $\leq$ $\left\lVert(X_{I}^{T}X_{I})^{-1}X_{I}^{T}X_{{S_{{\mathcal{D}}}}}\beta_{{S_{{\mathcal{D}}}}}\right\rVert_{2}+\left\lVert(X_{I}^{T}X_{I})^{-1}X_{I}^{T}\epsilon\right\rVert_{2}$
$\leq$ $\frac{\theta_{|I|,|{S_{{\mathcal{D}}}}|}}{\Lambda_{\min}(|I|)}\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}+\frac{\sqrt{3|I|}\sigma}{\sqrt{2n\Lambda_{\min}(|I|)}}.$
Now the lemma holds given that $\hat{\beta}_{I}^{\mathop{\text{\rm ols}\kern 0.60275pt}}=\hat{\beta}_{I}$, $\hat{\beta}_{I^{c}}^{\mathop{\text{\rm ols}\kern 0.60275pt}}=0$, and hence
$\left\lVert\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}-\beta\right\rVert_{2}^{2}$ $=$ $\left\lVert\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}_{I}-\beta_{I}\right\rVert_{2}^{2}+\left\lVert\beta_{I^{c}}\right\rVert_{2}^{2}=\left\lVert\hat{\beta}_{I}-\beta_{I}\right\rVert_{2}^{2}+\left\lVert\beta_{I^{c}}\right\rVert_{2}^{2}$
$\leq$ $\frac{2\theta^{2}_{|I|,|{S_{{\mathcal{D}}}}|}}{\Lambda^{2}_{\min}(|I|)}\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}^{2}+\frac{3|I|\sigma^{2}}{n\Lambda_{\min}(|I|)}+\left\lVert\beta_{{S_{{\mathcal{D}}}}}\right\rVert_{2}^{2}.$
Remark 6.1.
We can also derive the lower bound on event $\mathcal{Q}^{c}$ for $\left|I\right|=m$
$\left\lVert(X_{I}^{T}X_{I})^{-1}X_{I}^{T}\epsilon/\sigma\right\rVert_{2}^{2}$ $=$ $Q^{T}(X_{I}^{T}X_{I})^{-1}Q\geq\Lambda_{\min}\left((X_{I}^{T}X_{I})^{-1}\right)\left\lVert Q\right\rVert_{2}^{2}$
$\geq$ $\frac{m}{2\Lambda_{\max}\left(X_{I}^{T}X_{I}\right)}=\frac{m}{2n\Lambda_{\max}(\left\lvert I\right\rvert)}$,
where we used an upper bound on $\left\lVert Q\right\rVert_{2}^{2}\geq m/2$, $\left\lVert Q\right\rVert_{2}^{2}\sim\chi^{2}_{m}$, and the fact that
$\Lambda_{\min}\left((X_{I}^{T}X_{I})^{-1}\right)=\frac{1}{\Lambda_{\max}\left((X_{I}^{T}X_{I})\right)}=\frac{1}{n\Lambda_{\max}\left((X_{I}^{T}X_{I})/n\right)}=:\frac{1}{n\Lambda_{\max}(\left\lvert I\right\rvert)}$.
Now suppose $\left\lvert I\right\rvert\asymp 2s_{0}$, then
$\left\lVert(X_{I}^{T}X_{I})^{-1}X_{I}^{T}\epsilon/\sigma\right\rVert_{2}^{2}$ $\geq$ $\frac{m}{2n\Lambda_{\max}(2s_{0})}.$
6.5 Proof of Lemma 2.7
The only part we make modification is the following: On event ${\mathcal{T}}_{\alpha}$, we have by (31), for an arbitrary set $I$ of size $\left|I\right|\leq 2s$,
$\left\lVert(X_{I}^{T}X_{I})^{-1}X_{I}^{T}\epsilon\right\rVert_{2}\leq\left\lVert(X_{I}^{T}X_{I}/n)^{-1}\right\rVert_{2}\left\lVert X_{I}^{T}\epsilon/n\right\rVert_{2}\leq\frac{\sqrt{|I|}\lambda_{\sigma,a,p}}{\Lambda_{\min}(|I|)}.$ (103)
Now we have on ${\mathcal{T}}_{a}$, by (103), (99), (100), and (6.4),
$\left\lVert\hat{\beta}_{I}-\beta_{I}\right\rVert_{2}$ $\leq$ $\left\lVert(X_{I}^{T}X_{I})^{-1}X_{I}^{T}X_{{S_{{\mathcal{D}}}}}\beta_{{S_{{\mathcal{D}}}}}\right\rVert_{2}+\left\lVert(X_{I}^{T}X_{I})^{-1}X_{I}^{T}\epsilon\right\rVert_{2}$
$\leq$ $\frac{\theta_{|I|,|{S_{{\mathcal{D}}}}|}}{\Lambda_{\min}(|I|)}\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}+\frac{\sqrt{|I|}\sigma\sqrt{1+a}\lambda}{\Lambda_{\min}(|I|)}.$
and thus (33) holds since $\left\lVert\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)-\beta\right\rVert_{2}^{2}=\left\lVert\hat{\beta}_{I}-\beta_{I}\right\rVert_{2}^{2}+\left\lVert\beta_{I^{c}}\right\rVert_{2}^{2}$ by (6.4). ∎
6.6 Proof of Lemma 4.2
We write $\beta=\beta^{(11)}+\beta^{(12)}+\beta^{(2)}$ where
$\beta_{j}^{(11)}=\beta_{j}\cdot 1_{1\leq j\leq a_{0}},\;\beta_{j}^{(12)}=\beta_{j}\cdot 1_{a_{0}<j\leq s_{0}},\;\text{ and }\;\beta^{(2)}=\beta_{j}\cdot 1_{j>s_{0}}.$
By definition of $s_{0}$ as in (13), we have $\sum_{i=1}^{p}\min(\beta_{i}^{2},\lambda^{2}\sigma^{2})\leq s_{0}\lambda^{2}\sigma^{2}$. Now
$\sum_{j\leq a_{0}}\min(\beta_{j}^{2},\lambda^{2}\sigma^{2})$ $=$ $a_{0}\lambda^{2}\sigma^{2},\text{ since }\left\lvert\beta_{j}\right\rvert\geq\lambda\sigma,\;\text{ and hence}$
$\sum_{j>a_{0}}\min(\beta_{j}^{2},\lambda^{2}\sigma^{2})$ $=$ $\left\lVert\beta^{(12)}+\beta^{(2)}\right\rVert_{2}^{2}\leq(s_{0}-a_{0})\lambda^{2}\sigma^{2}.$ (104)
For ${\mathcal{D}}_{11}={\mathcal{D}}\cap A_{0}$, we have ${\mathcal{D}}_{11}\subset A_{0}\subset T_{0}\subset S$, and $|{\mathcal{D}}_{11}|\leq a_{0}$. Then $\left\lVert\beta_{{\mathcal{D}}_{11},\text{\rm init}}\right\rVert_{\infty}<t_{0}$. Let $\beta_{{\mathcal{D}}}^{(11)}:=(\beta_{j})_{j\in A_{0}\cap{\mathcal{D}}}$. Now by the triangle inequality,
$\left\lVert\beta_{{\mathcal{D}}}^{(11)}\right\rVert_{2}$ $=$ $\left\lVert\beta_{{\mathcal{D}}_{11}}\right\rVert_{2}\leq\left\lVert\beta_{{\mathcal{D}}_{11},\text{\rm init}}\right\rVert_{2}+\left\lVert\beta_{{\mathcal{D}}_{11},\text{\rm init}}-\beta_{{\mathcal{D}}_{11}}\right\rVert_{2}$ (105)
$\leq$ $t_{0}\sqrt{\left|{\mathcal{D}}_{11}\right|}+\left\lVert h_{{\mathcal{D}}_{11}}\right\rVert_{2}\leq t_{0}\sqrt{a_{0}}+\left\lVert h_{{\mathcal{D}}{11}}\right\rVert_{2};$
[…]
Now (57) holds when we replace the crude bound of $|{\mathcal{D}}_{11}|\leq a_{0}$ with
[…]
$\left\lVert\beta_{{\mathcal{D}}}^{(11)}\right\rVert_{2}$ $\leq$ $t_{0}\frac{\left\lVert h_{{\mathcal{D}}_{11}}\right\rVert_{2}}{\beta_{\min,A_{0}}-t_{0}}+\left\lVert h_{{\mathcal{D}}_{11}}\right\rVert_{2}=\left\lVert h_{{\mathcal{D}}_{11}}\right\rVert_{2}\frac{\beta_{\min,A_{0}}}{\beta_{\min,A_{0}}-t_{0}}.$
6.7 Proof of Lemma 4.3
By the choice of $t_{0}$ in (4.3) and by (58),
$\min_{i\in A_{0}}\left\lvert\beta_{\text{\rm init},i}\right\rvert\geq\left\lvert\beta_{\min,A_{0}}-\left\lVert h_{A_{0}}\right\rVert_{\infty}\right\rvert\geq t_{0},\quad\text{and}\quad{\mathcal{D}}_{11}=\emptyset.$
Thus by (4.3), we can bound $|I\cap T_{0}^{c}|$, by either
$|I\cap T_{0}^{c}|\leq{\left\lVert\beta_{T_{0}^{c},\text{\rm init}}\right\rVert_{1}}/{t_{0}}\leq\breve{s}_{0}\quad\text{ or }\quad|I\cap T_{0}^{c}|\leq{\left\lVert\beta_{T_{0}^{c},\text{\rm init}}\right\rVert_{2}^{2}}/{t_{0}^{2}}\leq\breve{s}_{0},$
depending on which one is applicable. The rest follows from the proof of Lemma 4.2. ∎
7 Proof sketch of Theorem 3 in [43]
Let $\hat{S}:=\mathop{\text{\rm supp}\kern 0.86108pt}(\beta_{\text{\rm init}})$. Following [43], we use $A_{1}$ to denote the union of support set for $\beta_{\text{\rm init}}$ and $H_{q}$:
$A_{1}:=\{j:\beta_{\text{\rm init},j}\not=0\;\text{or}\;j\in H_{q}\}=\hat{S}(\lambda)\cup H_{q}.$
Then clearly, $A_{1}^{c}\subset H_{0}=\{q+1,\ldots,p\}$. In fact, by assumption (84), $\left|H_{q}\cup T_{1}^{*}\right|<q^{*}/(2\sqrt{c^{*}})$ as shown earlier. From this, we know that $q^{*}=\Omega(s_{0})$; cf. (81).
Under the SRC (83) and sparsity conditions (84), (80), and (82), and with suitable choices of the penalty $\lambda_{n}=2\lambda\sigma\sqrt{(1+a)c^{*}}$ for some $a>0$, the following statements (along with other results) hold with high probability (w.h.p.),
$(A_{1})\;\;\left|\mathop{\text{\rm supp}\kern 0.86108pt}(\beta_{\text{\rm init}})\cup H_{q}\right|$ $=$ $\left|A_{1}\right|\leq M_{1}^{*}q<q^{*};\text{cf. (2.21) and (3.1)},$ (106)
$\left\lVert\beta_{\text{\rm init},A_{1}}-\beta_{A_{1}}\right\rVert_{2}$ $\leq$ $O(1)\lambda\sigma\sqrt{\left\lvert A_{1}\right\rvert}\leq O(1)\lambda\sigma\sqrt{q^{*}},$ (107)
$\left\lVert X(\beta_{\text{\rm init}}-\beta)\right\rVert_{2}/\sqrt{n}$ $\leq$ $O(1)\sigma\lambda\sqrt{\left\lvert A_{1}\right\rvert}\quad\text{cf. (3.5)}.$ (108)
By definition of $A_{1}:=\mathop{\text{\rm supp}\kern 0.86108pt}(\beta_{\text{\rm init}})\cup H_{q}$, it holds that $\beta_{\text{\rm init},j}=0,\;\forall j\in A_{1}^{c}$. Moreover, since $A_{1}^{c}\subset H_{0}$, we have by (80),
$\left\lVert\beta_{\text{\rm init},A_{1}^{c}}-\beta_{A_{1}^{c}}\right\rVert_{1}$ $\leq$ $\left\lVert\beta_{H_{0}}\right\rVert_{1}\leq\eta_{1}=O(\frac{r_{1}^{2}q}{\sqrt{c^{*}}}2\lambda\sigma)=\tilde{O}(q^{*}\lambda\sigma),$
$\left\lVert\beta_{\text{\rm init},A_{1}^{c}}-\beta_{A_{1}^{c}}\right\rVert_{2}^{2}$ $\leq$ $\left\lVert\beta_{H_{0}}\right\rVert_{2}^{2}=\tilde{O}(\frac{r_{2}}{\sqrt{c_{*}}}q^{*}\lambda^{2}\sigma^{2}).$ (109)
To show (109), we first bound the $\ell_{2}$ norm on the following set $T_{1}^{*}=\{j:j\in H_{0},\left\lvert\beta_{j}\right\rvert\geq\lambda\sigma\}$ by (82),
$\left\lVert\beta_{T_{1}^{*}}\right\rVert_{2}^{2}$ $=$ $\sum_{j>q}^{p}\beta_{j}^{2}I(\left\lvert\beta_{j}\right\rvert\geq\lambda\sigma)\leq\left\lVert\sum_{j\in T_{1}^{*}}\beta_{j}x_{j}\right\rVert_{2}^{2}/(nc_{*})$
$\leq$ $\eta_{2}^{2}/(nc_{*})=\tilde{O}(q(2r_{2}\sigma\lambda)^{2}/{c_{*}});$
Now clearly, (109) holds since for all $j\in H_{0}\setminus T_{1}^{*}$, $\left\lvert\beta_{j}\right\rvert<\lambda\sigma$, and hence
$\left\lVert\beta_{H_{0}\setminus T_{1}^{*}}\right\rVert_{2}^{2}$ $=$ $\left\lVert\beta_{[p]\setminus L(q)}\right\rVert_{2}^{2}\leq\left\lVert\beta_{H_{0}\setminus T_{1}^{*}}\right\rVert_{1}(\lambda\sigma)$ (110)
$\leq$ $\eta_{1}\lambda\sigma=\frac{r_{1}^{2}}{\sqrt{c^{*}}}q2\sigma^{2}\lambda^{2}=\tilde{O}(2r_{1}^{2}q(\sigma\lambda)^{2}).$ (111)
Putting these two sets together, one can obtain (109), since
$\left\lVert\beta_{H_{0}}\right\rVert_{2}^{2}$ $=$ $\left\lVert\beta_{T_{1}^{*}}\right\rVert_{2}^{2}+\left\lVert\beta_{H_{0}\setminus T_{1}^{*}}\right\rVert_{2}^{2}\leq\tilde{O}((\frac{r_{2}^{2}}{c_{*}}+r_{1}^{2})q(2\sigma\lambda)^{2})$ (112)
$=$ $\tilde{O}(\frac{r_{2}}{\sqrt{c_{*}}}q^{*}\sigma^{2}\lambda^{2})\text{ since }\;\;4(\frac{r_{2}}{\sqrt{c_{*}}}+r_{1}^{2})q\leq q^{*}.$
Then one obtains (86) by (80), (107) and (109). However, the largest signal in $H_{0}$ can be $\asymp\sqrt{q}\sigma\lambda$ via the $\eta_{2}$ condition (82):
$\left\lVert\beta_{H_{0}}\right\rVert_{\infty}\leq\sqrt{\eta_{2}^{2}/(nc_{*})}=\tilde{O}(\frac{2r_{2}}{\sqrt{c_{*}}}\sqrt{q}\sigma\lambda).$ (113)
We compare the conditions and error bounds with those in Theorem 2.4 in Section 4.3.
8 Conclusion
In this paper, we show that the thresholding method is effective in variable selection and accurate in statistical estimation. It improves the ordinary Lasso in significant ways. For example, we allow very significant number of non-zero elements in the true parameter, for which the ordinary Lasso would have failed. On the theoretical side, we show that if $X$ obeys the RE condition and if the true parameter is sufficiently sparse, the Thresholded Lasso achieves the $\ell_{2}$ loss within a logarithmic factor of the ideal mean square error one would achieve with an oracle, while selecting a sufficiently sparse model $I$. This is accomplished when threshold level is at about $\sqrt{2\log p/n}\sigma$, assuming that columns of $X$ have $\ell_{2}$ norm $\sqrt{n}$. When the SNR is high, almost exact recovery of the non-zeros in $\beta$ is possible as shown in our theory; exact recovery of the support of $\beta$ is shown in our simulation study when $n$ is only linear in $s$ for several Gaussian and Bernoulli random ensembles. When the SNR is relatively low, the inference task is difficult for any estimator. In this case, we show that Thresholded Lasso tradeoffs Type I and II errors nicely: we recommend choosing the thresholding parameter conservatively. These findings not only validate our theoretical analysis excellently but also indicate that in practical applications, this approach could be made very effective and relevant.
Acknowledgement
I dedicate this work to Professor Peter Bühlmann, on the occasion of his 60th birthday. His unwavering support and mentorship during my career has been deeply appreciated. The author thanks the Editor, an AE and two anonymous referees for their valuable comments and suggestions. I thank Professor Bin Yu warmly for hosting my visit at UC Berkeley while I finished the first draft of this work in Spring 2010. The author thanks Peter Bühlmann, Emmanuel Candès, Sara van de Geer, John Lafferty, Po-Ling Loh, Richard Samworth, Xiaotong Shen, Martin Wainwright, Larry Wasserman, Dana Yang and Cun-Hui Zhang for helpful discussions. Part of this work has appeared in a conference paper with title: Thresholding Procedures for High Dimensional Variable Selection and Statistical Estimation, in Proceedings of Advances in Neural Information Processing Systems 22.
Appendix A Proof of the MSE lower bound
We show the proof of (10) for self-containment. Note that due to different normalization of columns of $X$, our expressions are slightly different from those by [9]. Hence we give a complete derivation here. Let $I$ be a fixed subset of indices and consider the OLS estimator $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}$ such that $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}_{I}=\hat{\beta}_{I}=(X_{I}^{T}X_{I})^{-1}X_{I}^{T}Y$ and $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}_{I^{c}}=\hat{\beta}_{I^{c}}=0$. Consider the least square estimator $\hat{\beta}_{I}=(X_{I}^{T}X_{I})^{-1}X_{I}^{T}Y$, where $|I|\leq s$ (and $\hat{\beta}_{I^{c}}=0$), and consider the ideal least-squares estimator $\beta^{\diamond}$ which minimizes the expected mean squared error
$\beta^{\diamond}=\operatorname{argmin}_{I\subset\{1,\ldots,p\},\;|I|\leq s}{\textbf{E}}\left\lVert\beta-\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)\right\rVert_{2}^{2}.$ (114)
Proposition A.1.
[9] If $\Lambda_{\max}(s)<\infty$, then
${\textbf{E}}\left\lVert\beta-\beta^{\diamond}\right\rVert_{2}^{2}\geq\min\left(1,1/\Lambda_{\max}(s)\right)\sum_{i=1}^{p}\min(\beta_{i}^{2},\sigma^{2}/n).$ (115)
Proof.
Here we denote by $\beta_{I}$ the restriction of $\beta$ to the set $I$, where $I\subset[p]$, and $\beta^{\mathop{\natural}\kern 0.60275pt}({I})$ its 0-extended version such that $\beta^{\mathop{\natural}\kern 0.60275pt}_{I^{c}}(I)=0$ and $\beta^{\mathop{\natural}\kern 0.60275pt}_{I}(I)=\beta_{I}$. The error of this OLS estimator, namely, $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}$ is given by
$\left\lVert\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)-\beta\right\rVert_{2}^{2}$ $=$ $\left\lVert\hat{\beta}_{I}-\beta_{I}\right\rVert_{2}^{2}+\left\lVert\beta^{\mathop{\natural}\kern 0.60275pt}({I})-\beta\right\rVert_{2}^{2}.$ (116)
The first term is:
$\hat{\beta}_{I}-\beta_{I}$ $:=$ $(X_{I}^{T}X_{I})^{-1}X_{I}^{T}Y-\beta_{I}$
$=$ $(X_{I}^{T}X_{I})^{-1}X_{I}^{T}(X_{I}\beta_{I}+X_{I^{c}}\beta_{I^{c}}+\epsilon)-\beta_{I}$
$=$ $(X_{I}^{T}X_{I})^{-1}X_{I}^{T}X_{I^{c}}\beta_{I^{c}}+(X_{I}^{T}X_{I})^{-1}X_{I}^{T}\epsilon,$
and its mean squared error is given by
${\textbf{E}}\left\lVert\hat{\beta}_{I}-\beta_{I}\right\rVert_{2}^{2}$ $=$ $\left\lVert(X_{I}^{T}X_{I})^{-1}X_{I}^{T}X_{I^{c}}\beta_{I^{c}}\right\rVert_{2}^{2}+{\textbf{E}}\left\lVert(X_{I}^{T}X_{I})^{-1}X_{I}^{T}\epsilon\right\rVert_{2}^{2},$
where
${\textbf{E}}\left\lVert(X_{I}^{T}X_{I})^{-1}X_{I}^{T}\epsilon\right\rVert_{2}^{2}$ $=$ $\frac{\sigma^{2}}{n}\textsf{Tr}\left(\left(\frac{X_{I}^{T}X_{I}}{n}\right)^{-1}\right)\geq\frac{\sigma^{2}}{n}\frac{|I|}{\Lambda_{\max}(|I|)},$
since eigenvalues of $\left(\frac{X_{I}^{T}X_{I}}{n}\right)^{-1}$ are in the range of $\left[\frac{1}{\Lambda_{\max}\left({X_{I}^{T}X_{I}}/{n}\right)},\frac{1}{\Lambda_{\min}\left({X_{I}^{T}X_{I}}/{n}\right)}\right]$. Thus
${\textbf{E}}\left\lVert\hat{\beta}_{I}-\beta_{I}\right\rVert_{2}^{2}$ $\geq$ $\frac{\sigma^{2}}{n}\frac{|I|}{\Lambda_{\max}(|I|)}.$ (117)
Thus for all sets $I$ such that $|I|\leq s$ and for $\Lambda_{\max}(s)<\infty$, we have for $\hat{\beta}_{I^{c}}=0$ and by (116),
${\textbf{E}}\left\lVert\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}({I})-\beta\right\rVert_{2}^{2}$ $=$ ${\textbf{E}}\left\lVert\hat{\beta}_{I}-\beta_{I}\right\rVert_{2}^{2}+\left\lVert\beta_{I^{c}}\right\rVert_{2}^{2}$
$\geq$ $\frac{\sigma^{2}}{n}\frac{|I|}{\Lambda_{\max}(s)}+\left\lVert\beta_{I^{c}}\right\rVert_{2}^{2}$
$\geq$ $\min\left(1,1/\Lambda_{\max}(s)\right)\left(\sum_{j\in I^{c}}\beta_{j}^{2}+\frac{\sigma^{2}}{n}|I|\right),$
which gives that the ideal mean squared error is bounded below by
${\textbf{E}}\left\lVert\beta-\beta^{\diamond}\right\rVert_{2}^{2}$ $\geq$ $\min\left(1,1/\Lambda_{\max}(s)\right)\min_{I}\left(\sum_{j\in I^{c}}\beta_{j}^{2}+\frac{\sigma^{2}}{n}|I|\right)$
$=$ $\min\left(1,1/\Lambda_{\max}(s)\right)\sum_{i=1}^{p}\min(\beta_{i}^{2},\sigma^{2}/n).$
Appendix B Proof of Lemma 2.8
It is sufficient to show that (118) holds for $\left\lVert c\right\rVert_{2}=\left\lVert c^{\prime}\right\rVert_{2}=1$.
$\frac{\left\lvert{\;\langle{\,X_{I}c,X_{S_{{\mathcal{D}}}}{c^{\prime}}\,}\rangle\;}\right\rvert}{n}\leq\frac{(\Lambda_{\max}(2s)-\Lambda_{\min}(2s))}{2}.$ (118)
Indeed, by (6) and (5), we have $2\Lambda_{\min}(2s)\leq\left\lVert X_{I}c+X_{S_{{\mathcal{D}}}}{c^{\prime}}\right\rVert_{2}^{2}/{n}\leq 2\Lambda_{\max}(2s)$ and $2\Lambda_{\min}(2s)\leq{\left\lVert X_{I}c-X_{S_{{\mathcal{D}}}}{c^{\prime}}\right\rVert_{2}^{2}}/{n}\leq 2\Lambda_{\max}(2s)$. Hence (118) follows from the parallelogram identity:
$\left\lvert{\;\langle{\,X_{I}c,X_{S_{{\mathcal{D}}}}{c^{\prime}}\,}\rangle\;}\right\rvert=\left\lvert\left\lVert X_{I}c+X_{S_{{\mathcal{D}}}}{c^{\prime}}\right\rVert_{2}^{2}-\left\lVert X_{I}c-X_{S_{{\mathcal{D}}}}{c^{\prime}}\right\rVert_{2}^{2}\right\rvert/{4}.$
Next, suppose $\Lambda_{\min}(2s)=0$. Then
$\left\lvert{\;\langle{\,X_{I}c,X_{S_{{\mathcal{D}}}}{c^{\prime}}\,}\rangle\;}\right\rvert$ $=$ $\left\lvert\left\lVert X_{I}c+X_{S_{{\mathcal{D}}}}{c^{\prime}}\right\rVert_{2}^{2}-\left\lVert X_{I}c-X_{S_{{\mathcal{D}}}}{c^{\prime}}\right\rVert_{2}^{2}\right\rvert/{4}$
$\leq$ $\frac{\left\lVert X_{I}c+X_{S_{{\mathcal{D}}}}{c^{\prime}}\right\rVert_{2}^{2}}{4}\vee\frac{\left\lVert X_{I}c-X_{S_{{\mathcal{D}}}}{c^{\prime}}\right\rVert_{2}^{2}}{4}$
$\leq$ $\Lambda_{\max}(2s)/2.$
Moreover, (34) follows from the arguments as in (30), using the Cauchy-Schwarz inequality. ∎
Appendix C Proof of Lemma 3.1
Decompose $h_{T_{01}^{c}}$ into $h_{T_{2}}$, …, $h_{T_{K}}$ such that $T_{2}$ corresponds to locations of the $s_{0}$ largest coefficients of $h_{T_{01}^{c}}$ in absolute values, and $T_{3}$ corresponds to locations of the next $s_{0}$ largest coefficients of $h_{T_{01}^{c}}$ in absolute values, and so on. Let $V$ be the span of columns of $X_{j}$, where $j\in T_{01}$, and $P_{V}$ be the orthogonal projection onto $V$. Decompose $P_{V}Xh$:
$P_{V}Xh$ $=$ $P_{V}Xh_{T_{01}}+\sum_{j\geq 2}P_{V}Xh_{T_{j}}=Xh_{T_{01}}+\sum_{j\geq 2}P_{V}Xh_{T_{j}},\text{ where }$
$\forall j\geq 2,\;\;\left\lVert P_{V}Xh_{T_{j}}\right\rVert_{2}$ $\leq$ $\frac{\sqrt{n}\theta_{s_{0},2s_{0}}}{\sqrt{\Lambda_{\min}(2s_{0})}}\left\lVert h_{T_{j}}\right\rVert_{2}\text{ and }\sum_{j\geq 2}\left\lVert h_{T_{j}}\right\rVert_{2}\leq\left\lVert h_{T_{0}^{c}}\right\rVert_{1}/\sqrt{s_{0}};$
see [9] for details. Thus we have
$\left\lVert Xh_{T_{01}}\right\rVert_{2}$ $=$ $\left\lVert P_{V}Xh-\sum_{j\geq 2}P_{V}Xh_{T_{j}}\right\rVert_{2}\leq\left\lVert P_{V}Xh\right\rVert_{2}+\left\lVert\sum_{j\geq 2}P_{V}Xh_{T_{j}}\right\rVert_{2}$
$\leq$ $\left\lVert Xh\right\rVert_{2}+\sum_{j\geq 2}\left\lVert P_{V}Xh_{T_{j}}\right\rVert_{2}\leq\left\lVert Xh\right\rVert_{2}+\frac{\sqrt{n}\theta_{s_{0},2s_{0}}}{\sqrt{\Lambda_{\min}(2s_{0})}\sqrt{s_{0}}}\left\lVert h_{T_{0}^{c}}\right\rVert_{1},$
where we used the fact that $\left\lVert P_{V}\right\rVert_{2}\leq 1$. Hence the lemma follows given
$\left\lVert h_{T_{01}}\right\rVert_{2}\leq\frac{1}{\sqrt{\Lambda_{\min}(2s_{0})}\sqrt{n}}\left\lVert Xh_{T_{01}}\right\rVert_{2}.$
For other bounds, the fact that the $k$th largest value of $h_{T_{0}^{c}}$ obeys $\left|h_{T_{0}^{c}}\right|_{(k)}\leq\left\lVert h_{T_{0}^{c}}\right\rVert_{1}/k$ has been used; see [9]. ∎
Appendix D Proof of Theorem 2.4
Let $\beta_{T_{0}}$ be the restriction of $\beta$ to the set $T_{0}$. Let $h=\beta_{\text{\rm init}}-\beta^{\mathop{\natural}\kern 0.60275pt}({T_{0}})$. Throughout this proof, as a shorthand, we write $h=\hat{\beta}-\beta_{0}$, where $\hat{\beta}=\beta_{\text{\rm init}}$ and $\beta_{0}=\beta^{\mathop{\natural}\kern 0.60275pt}({T_{0}})$ as defined in Lemma 2.3. We first show Lemma D.1, which gives us the prediction error using $\beta_{T_{0}}$. We do not focus on obtaining the best constants in the proof of Theorem 2.4. Recall we define a quantity $\lambda_{\sigma,a,p}$, which bounds the maximum correlation between the noise and covariates of $X$; For each $a\geq 0$, let
${{\mathcal{T}}_{a}}:=\biggl\{\epsilon:\left\lVert{X^{T}\epsilon}/{n}\right\rVert_{\infty}\leq\lambda_{\sigma,a,p},\text{ where }\lambda_{\sigma,a,p}=\sigma\sqrt{1+a}\sqrt{{2\log p}/{n}}\biggr\}.$ (119)
Then, we have ${\mathbb{P}}\left({\mathcal{T}}_{a}\right)\geq 1-(\sqrt{\pi\log p}p^{a})^{-1}$ when $X$ has column $\ell_{2}$ norms bounded by $\sqrt{n}$.
Lemma D.1.
Suppose $\beta$ is $s$-sparse. Let $T_{0}$ denote locations of the $s_{0}$ largest coefficients of $\beta$ in absolute values. Suppose (5) holds. We have for $\lambda=\sqrt{(2\log p)/n}$.
$\left\lVert X\beta-X\beta_{0}\right\rVert_{2}/\sqrt{n}$ $\leq$ $\sqrt{\Lambda_{\max}(s-s_{0})}\lambda\sigma\sqrt{s_{0}},$ (120)
where recall $\beta_{0}=\beta^{\mathop{\natural}\kern 0.60275pt}(T_{0})$, as defined in Lemma D.1.
Proof.
The lemma holds given that $\left\lVert\beta_{T_{0}^{c}}\right\rVert_{2}\leq\lambda\sigma\sqrt{s_{0}}$ by (13) and (14). Indeed, we have $\left\lvert\beta_{j}\right\rvert<\lambda\sigma$ for all $j\in T_{0}^{c}$ by definition of $T_{0}$ and hence
$\left\lVert\beta_{T_{0}^{c}}\right\rVert_{2}^{2}$ $\leq$ $\sum_{i=1}^{p}\min(\beta_{i}^{2},\lambda^{2}\sigma^{2})\leq s_{0}\lambda^{2}\sigma^{2}.$
Hence $\left\lVert X\beta-X\beta_{0}\right\rVert_{2}/\sqrt{n}\leq\sqrt{\Lambda_{\max}(s-s_{0})}\left\lVert\beta_{T_{0}^{c}}\right\rVert_{2}.$ ∎∎
Next, we state Lemma D.2. We prove Lemma D.2 in Section F.
Lemma D.2.
Suppose all conditions in Theorem 2.4 hold. Let $T_{1}$ be the $s_{0}$ largest positions of $h$ outside of $T_{0}$. Denote by $T_{01}=T_{0}\cup T_{1}$. Then
$\left\lVert Xh_{T_{01}}\right\rVert_{2}/\sqrt{n}$ $\leq$ $\left\lVert Xh\right\rVert_{2}/\sqrt{n}+{\sqrt{\Lambda_{\max}(s_{0})}}\left\lVert h_{T_{0}^{c}}\right\rVert_{1}/{\sqrt{s_{0}}}.$ (121)
Then by Lemma 3.1 and (121), we have
$\left\lVert h_{T_{01}}\right\rVert_{2}$ $\leq$ $\frac{1}{\sqrt{\Lambda_{\min}(2s_{0})}}\big(\left\lVert Xh\right\rVert_{n}+\ell(s_{0})\left\lVert h_{T_{0}^{c}}\right\rVert_{1}/\sqrt{s_{0}}\big)$
$\;\text{where}\;\ell(s_{0})$ $:=$ $\frac{\theta_{s_{0},2s_{0}}}{\sqrt{\Lambda_{\min}(2s_{0})}}\wedge\sqrt{\Lambda_{\max}(s_{0})}.$
D.1 Proof of Theorem 2.4
Throughout this proof, we assume that ${\mathcal{T}}_{a}$ holds.
Prelude. First recall the following elementary inequality. By the optimality of $\hat{\beta}=\beta_{\text{\rm init}}$, we have
$\;\;\;\;\;\frac{1}{2n}\left\lVert Y-X\hat{\beta}\right\rVert^{2}_{2}-\frac{1}{2n}\left\lVert Y-X\beta_{0}\right\rVert^{2}_{2}\leq\lambda_{n}\left\lVert\beta_{T_{0}}\right\rVert_{1}-\lambda_{n}\left\lVert\hat{\beta}\right\rVert_{1},$ (122)
where
$\left\lVert Y-X\hat{\beta}\right\rVert_{2}^{2}$ $=$ $\left\lVert X\beta-X\hat{\beta}+\epsilon\right\rVert_{2}^{2}$
$=$ $\left\lVert X\hat{\beta}-X\beta\right\rVert_{2}^{2}+2(\beta-\hat{\beta})^{T}X^{T}\epsilon+\left\lVert\epsilon\right\rVert^{2}_{2},\text{ and }$
$\left\lVert Y-X\beta_{0}\right\rVert_{2}^{2}$ $=$ $\left\lVert X\beta-X\beta_{0}+\epsilon\right\rVert_{2}^{2}$
$=$ $\left\lVert X\beta-X\beta_{0}\right\rVert_{2}^{2}+2(\beta-\beta_{0})^{T}X^{T}\epsilon+\left\lVert\epsilon\right\rVert^{2}_{2}.$
Thus by (122) and the triangle inequality, we have on ${\mathcal{T}}_{a}$ and $\lambda_{n}\geq 2\lambda_{\sigma,a,p}$ and $\left\lVert\frac{X^{T}\epsilon}{n}\right\rVert_{\infty}\leq\lambda_{\sigma,a,p}\leq\lambda_{n}/2$,
$\frac{\left\lVert X\hat{\beta}-X\beta\right\rVert_{2}^{2}}{n}$ $\leq$ $\frac{\left\lVert X\beta-X\beta_{0}\right\rVert_{2}^{2}}{n}+\frac{2h^{T}X^{T}\epsilon}{n}+2\lambda_{n}(\left\lVert\beta_{0}\right\rVert_{1}-\left\lVert h+\beta_{0}\right\rVert_{1})$
$\leq$ $\frac{\left\lVert X\beta-X\beta_{0}\right\rVert_{2}^{2}}{n}+2\left\lVert h\right\rVert_{1}\left\lVert\frac{X^{T}\epsilon}{n}\right\rVert_{\infty}+2\lambda_{n}(\left\lVert h_{T_{0}}\right\rVert_{1}-\left\lVert h_{T_{0}^{c}}\right\rVert_{1})$
$\leq$ $\frac{\left\lVert X\beta-X\beta_{0}\right\rVert_{2}^{2}}{n}+3\lambda_{n}\left\lVert h_{T_{0}}\right\rVert_{1}-\lambda_{n}\left\lVert h_{T_{0}^{c}}\right\rVert_{1},$
where we have used the fact that $\lambda_{n}\geq 2\lambda_{\sigma,a,p}$ for $a\geq 0$. Thus we have on ${\mathcal{T}}_{a}$,
$\left\lVert X\hat{\beta}-X\beta\right\rVert_{2}^{2}/n+\lambda_{n}\left\lVert h_{T_{0}^{c}}\right\rVert_{1}\leq\left\lVert X\beta-X\beta_{0}\right\rVert_{2}^{2}/n+3\lambda_{n}\left\lVert h_{T_{0}}\right\rVert_{1},$ (123)
This is the inequality we used in [48]. In the updated proof, we will apply the following Lemma D.3; cf. Lemma A.2 [2].
Lemma D.3.
[Lemma A.2] [2] Let $h:\mathbb{R}^{p}\longrightarrow\mathbb{R}$ be a convex function. Let $f,\xi\in\mathbb{R}^{n}$, $y=f+\xi$ and let $X$ be any $n\times p$ matrix. If $\hat{\beta}$ is a solution of the minimization problem $\min_{\beta\in\mathbb{R}^{p}}(\left\lVert X\beta-y\right\rVert^{2}_{n}+h(\beta)$, then $\hat{\beta}$ satisfies for all $\tilde{\beta}\in\mathbb{R}^{p}$
$\left\lVert X\hat{\beta}-f\right\rVert^{2}_{n}+\left\lVert X(\hat{\beta}-\tilde{\beta})\right\rVert^{2}_{n}$
[…]
We use $\hat{\beta}:=\beta_{\text{\rm init}}$ to represent the solution to the Lasso estimator in (2). Using Lemma D.3 with $y=f+\epsilon$, where $f=X\beta$, we obtain eq. (20) [11], where we set $\bar{\beta}=\beta_{0}:=\beta^{\mathop{\natural}\kern 0.60275pt}({T_{0}})$ and $\bar{\delta}=\hat{\beta}-\beta_{0}=h$,
$\left\lVert X(\hat{\beta}-\beta)\right\rVert^{2}_{n}+\left\lVert X(\hat{\beta}-\beta_{0})\right\rVert^{2}_{n}\leq\left\lVert X(\beta-\beta_{0})\right\rVert^{2}_{n}$ (125)
$+\frac{2}{n}\epsilon^{T}X(\hat{\beta}-\beta_{0})+2\lambda_{n}(\left\lVert\beta_{0}\right\rVert_{1}-\left\lVert\hat{\beta}\right\rVert_{1})$
Thus, we have the following updated inequality (42), replacing (D.1).
Denote by $\delta:=\hat{\beta}-\beta$ and $h=\hat{\beta}-\beta_{0}$. Thus we have by (125), on ${\mathcal{T}}_{a}$,
$\left\lVert X\delta\right\rVert^{2}_{n}+\left\lVert Xh\right\rVert^{2}_{n}$ $\leq$ $\left\lVert X\beta-X\beta_{0}\right\rVert^{2}_{n}+\frac{2h^{T}X^{T}\epsilon}{n}+2\lambda_{n}(\left\lVert\beta_{0}\right\rVert_{1}-\left\lVert h+\beta_{0}\right\rVert_{1})$ (126)
$\leq$ $\left\lVert X\beta-X\beta_{0}\right\rVert^{2}_{n}+2\left\lVert h\right\rVert_{1}\left\lVert\frac{X^{T}\epsilon}{n}\right\rVert_{\infty}+2\lambda_{n}(\left\lVert h_{T_{0}}\right\rVert_{1}-\left\lVert h_{T_{0}^{c}}\right\rVert_{1})$
$\leq$ $\left\lVert X\beta-X\beta_{0}\right\rVert^{2}_{n}+3\lambda_{n}\left\lVert h_{T_{0}}\right\rVert_{1}-\lambda_{n}\left\lVert h_{T_{0}^{c}}\right\rVert_{1},$
where we have used the fact that $\lambda_{n}\geq 2\lambda_{\sigma,a,p}$ for $a\geq 0$. (42) then follows immediately from (126). We consider three cases which do not need to be mutually exclusive.
Case 1. In the first case, suppose
$\left\lVert X\delta\right\rVert^{2}_{n}+\left\lVert Xh\right\rVert^{2}_{n}\geq\left\lVert X\beta-X\beta_{0}\right\rVert_{n}^{2}.$
Then by (42) and (43),
$\left\lVert h_{T_{0}^{c}}\right\rVert_{1}\leq 3\left\lVert h_{T_{0}}\right\rVert_{1},\text{ and hence }\;\;h\in{\mathcal{C}}(s_{0},3).$ (127)
Hence we can use the $\textnormal{{RE}}(s_{0},3,X)$ condition to bound $\left\lVert h_{T_{0}}\right\rVert_{2}$ with $K\left\lVert Xh\right\rVert_{n}$,
$\left\lVert X\delta\right\rVert^{2}_{n}+\left\lVert Xh\right\rVert^{2}_{n}-\left\lVert X(\beta-\beta_{0})\right\rVert_{n}^{2}+\lambda_{n}\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ (128)
$\leq$ $3\lambda_{n}\left\lVert h_{T_{0}}\right\rVert_{1}\leq 3\lambda_{n}\sqrt{s_{0}}\left\lVert h_{T_{0}}\right\rVert_{2}\leq 3\lambda_{n}\sqrt{s_{0}}K\left\lVert Xh\right\rVert_{n}$
$\leq$ $(3K\lambda_{n}\sqrt{s_{0}}/2)^{2}+\left\lVert Xh\right\rVert^{2}_{n}\;\text{where }\;K:=K(s_{0},3);$
Hence (45) follows immediately from (128), upon deleting the term $\left\lVert Xh\right\rVert^{2}_{n}$ from both sides. Moreover, we obtain the upper bound on $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ ( $\left\lVert h_{T_{0}}\right\rVert$) from (45) since
$\left\lVert X\delta\right\rVert^{2}_{n}+\lambda_{n}\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ $\leq$ $\left\lVert X(\beta-\beta_{0})\right\rVert_{n}^{2}+(3K\lambda_{n}\sqrt{s_{0}}/2)^{2},$
$\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ $\leq$ $\left\lVert X(\beta-\beta_{0})\right\rVert_{n}^{2}/\lambda_{n}+(3K/2)^{2}\lambda_{n}s_{0}$ (129)
$\leq$ $\big(\Lambda_{\max}(s-s_{0})/d_{0}+9d_{0}K(s_{0},3)^{2}/{4}\big)\lambda\sigma s_{0},$
following the proof of Lemma D.1, where $\lambda_{n}=d_{0}\lambda\sigma\geq 2\lambda_{\sigma,a,p}$. Moreover, we will show that by (128) and (45),
$\left\lVert h_{T_{0}}\right\rVert_{2}$ $\leq$ $K\left\lVert X\beta-X\beta_{0}\right\rVert_{n}+3K^{2}\lambda_{n}\sqrt{s_{0}}\leq D^{\prime}_{0}\lambda\sigma\sqrt{s_{0}},$
$\text{ where }\;D^{\prime}_{0}:=(K\sqrt{\Lambda_{\max}(s-s_{0})}+3d_{0}K^{2}).$
Then, we have by (127) and (D.1),
$\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ $\leq$ $3\left\lVert h_{T_{0}}\right\rVert_{1}\leq 3\sqrt{s_{0}}\left\lVert h_{T_{0}}\right\rVert_{2}\leq 3D^{\prime}_{0}\lambda\sigma s_{0}.$ (131)
Combining (131) and (D.1), we have
$\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ $\leq$ $D_{1,a}\lambda\sigma s_{0},\;\text{ where}$
$\;\;D_{1,a}$ $:=$ $\big(\Lambda_{\max}(s-s_{0})/d_{0}+9d_{0}K(s_{0},3)^{2}/{4}\big)\wedge 3D^{\prime}_{0}\;\;\;\text{ where}$
$3D^{\prime}_{0}$ $\leq$ $3K(s_{0},3)\sqrt{\Lambda_{\max}(s-s_{0})}+9d_{0}K^{2}(s_{0},3))$
$\leq$ $\Lambda_{\max}(s-s_{0})/d_{0}+9d_{0}K(s_{0},3)^{2}/{4}+9d_{0}K^{2}(s_{0},3).$
Thus (22) holds for Case 1, since for $\delta=X\hat{\beta}-X\beta$, we have by (45),
$\left\lVert X\delta\right\rVert_{n}$ $=$ $\left\lVert X\hat{\beta}-X\beta\right\rVert_{n}\leq\big(\left\lVert X\beta-X\beta_{0}\right\rVert^{2}_{n}+(\frac{3K\lambda_{n}\sqrt{s_{0}}}{2})^{2}\big)^{1/2}$
$\leq$ $\left\lVert X\beta-X\beta_{0}\right\rVert_{n}+\frac{3K(s_{0},3)\lambda_{n}\sqrt{s_{0}}}{2}.$
Upper bound on $\left\lVert h_{T_{0}}\right\rVert_{2}$ (44). Under the $\textnormal{{RE}}(s_{0},3,X)$ condition, we have by Definition (3) and (132), for $h=\hat{\beta}-\beta_{0}$, where $\beta_{0}=\beta^{\mathop{\natural}\kern 0.60275pt}({T_{0}})$,
$\left\lVert h_{T_{0}}\right\rVert_{2}^{2}$ $\leq$ $K(s_{0},3)^{2}\left\lVert Xh\right\rVert^{2}_{n}$
$\leq$ $K(s_{0},3)^{2}\left(\left\lVert X\beta-X\beta_{0}\right\rVert^{2}_{n}+3\lambda_{n}\sqrt{s_{0}}\left\lVert h_{T_{0}}\right\rVert_{2}\right),$
where recall we have on ${\mathcal{T}}_{a}$, by (42),
$\left\lVert Xh\right\rVert^{2}_{n}\leq\left\lVert X\beta-X\beta_{0}\right\rVert_{n}^{2}+3\lambda_{n}\left\lVert h_{T_{0}}\right\rVert_{1}\;\text{ where}\;\left\lVert h_{T_{0}}\right\rVert_{1}\leq\sqrt{s_{0}}\left\lVert h_{T_{0}}\right\rVert_{2}.$ (132)
Thus we have for $\lambda_{n}=d_{0}\lambda\sigma$ and $K:=K(s_{0},3)$,
$\left\lVert h_{T_{0}}\right\rVert_{2}^{2}-3K^{2}\lambda_{n}\sqrt{s_{0}}\left\lVert h_{T_{0}}\right\rVert_{2}\leq K^{2}\left\lVert X\beta-X\beta_{0}\right\rVert^{2}_{n},\;\text{and}$ (133)
$\big(\left\lVert h_{T_{0}}\right\rVert_{2}-\frac{3K^{2}}{2}\lambda_{n}\sqrt{s_{0}}\big)^{2}\leq K^{2}\left\lVert X\beta-X\beta_{0}\right\rVert^{2}_{n}+\big(\frac{3K^{2}}{2}\lambda_{n}\sqrt{s_{0}})^{2}.$
Thus we have (44) holds for Case 1 with $K:=K(s_{0},3)$, since
$\left\lvert\left\lVert h_{T_{0}}\right\rVert_{2}-\frac{3K^{2}}{2}\lambda_{n}\sqrt{s_{0}}\right\rvert \displaystyle\leq$ $K\left\lVert X\beta-X\beta_{0}\right\rVert_{n}+\frac{3K^{2}}{2}\lambda_{n}\sqrt{s_{0}},\text{ which implies that}$
$\left\lVert h_{T_{0}}\right\rVert_{2}$ $\leq$ $K\left\lVert X\beta-X\beta_{0}\right\rVert_{n}+3K^{2}\lambda_{n}\sqrt{s_{0}}.$
Thus we obtain
$\left\lVert h_{T_{0}}\right\rVert_{2}$ $\leq$ $D^{\prime}_{0,a}\lambda\sigma\sqrt{s_{0}},\;\text{ where }$
$D^{\prime}_{0,a}$ $:=$ $K(s_{0},3)\sqrt{\Lambda_{\max}(s-s_{0})}+3K(s_{0},3)^{2}d_{0}.$ (134)
Similarly, we can derive a bound on $\left\lVert h\right\rVert_{1}$ following (128) directly:
$\left\lVert X\delta\right\rVert^{2}_{n}+\left\lVert Xh\right\rVert^{2}_{n}+\lambda_{n}\left\lVert h_{T_{0}^{c}}\right\rVert_{1}+\lambda_{n}\left\lVert h_{T_{0}}\right\rVert_{1}-\left\lVert X\beta-X\beta_{0}\right\rVert^{2}_{n}$
$\leq$ $4\lambda_{n}\left\lVert h_{T_{0}}\right\rVert_{1}\leq 4\lambda_{n}\sqrt{s_{0}}\left\lVert h_{T_{0}}\right\rVert_{2}$
$\leq$ $4K\lambda_{n}\sqrt{s_{0}}\left\lVert Xh\right\rVert_{n}\leq\left\lVert Xh\right\rVert^{2}_{n}+(2K\lambda_{n}\sqrt{s_{0}})^{2},$
under $\textnormal{{RE}}(s_{0},3,X)$ condition, since $h\in{\mathcal{C}}(s_{0},3)$. Hence
$\left\lVert X\delta\right\rVert_{n}^{2}+\lambda_{n}\left\lVert h\right\rVert_{1}$ $\leq$ $\left\lVert X\beta-X\beta_{0}\right\rVert^{2}_{n}+(2K\lambda_{n}\sqrt{s_{0}})^{2}.$
Then
$\left\lVert h\right\rVert_{1}$ $=$ $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}+\left\lVert h_{T_{0}}\right\rVert_{1}\leq\left\lVert X\beta-X\beta_{0}\right\rVert_{n}^{2}/\lambda_{n}+4K^{2}\lambda_{n}s_{0}$ (135)
$\leq$ $D_{2,a}\lambda\sigma s_{0},\;\text{ where }\;D_{2,a}\leq\Lambda_{\max}(s-s_{0})/d_{0}+4K(s_{0},3)^{2}d_{0}.$
Case 2. Suppose
$\left\lVert X\delta\right\rVert^{2}_{n}+\left\lVert Xh\right\rVert^{2}_{n}\leq\lambda_{n}\left\lVert h_{T_{0}}\right\rVert_{1}.$ (136)
Let $T_{1}$ be the $s_{0}$ largest positions of $h$ outside of $T_{0}$. This is the easy case. First we show (138). By the triangle inequality,
$\left\lVert X\beta-X\beta_{0}\right\rVert_{n}\leq\left\lVert X(\beta-\hat{\beta})\right\rVert_{n}+\left\lVert X(\hat{\beta}-\beta_{0})\right\rVert_{n}=\left\lVert X\delta\right\rVert_{n}+\left\lVert Xh\right\rVert_{n}.$
Thus we have on ${\mathcal{T}}_{a}$, by (42),
$\left\lVert X\delta\right\rVert^{2}_{n}+\left\lVert Xh\right\rVert^{2}_{n}+\lambda_{n}\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ $\leq$ $\left\lVert X\beta-X\beta_{0}\right\rVert_{n}^{2}+3\lambda_{n}\left\lVert h_{T_{0}}\right\rVert_{1}$ (137)
$\leq$ $2(\left\lVert X\delta\right\rVert^{2}_{n}+\left\lVert Xh\right\rVert^{2}_{n})+3\lambda_{n}\left\lVert h_{T_{0}}\right\rVert_{1},$
since $(a+b)^{2}\leq 2a^{2}+2b^{2}$. Then we have by assumption (136) and (137),
$\lambda_{n}\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ $\leq$ $\left\lVert X\delta\right\rVert^{2}_{n}+\left\lVert Xh\right\rVert^{2}_{n}+3\lambda_{n}\left\lVert h_{T_{0}}\right\rVert_{1}\leq 4\lambda_{n}\left\lVert h_{T_{0}}\right\rVert_{1}.$ (138)
Then $h\in{\mathcal{C}}(s_{0},4)$. Hence, under $\textnormal{{RE}}(s_{0},4,X)$ condition, we have by assumption (136), (138) and Definition (3),
$\left\lVert X\delta\right\rVert^{2}_{n}+\left\lVert Xh\right\rVert^{2}_{n}$ $\leq$ $\lambda_{n}\left\lVert h_{T_{0}}\right\rVert_{1}\leq\lambda_{n}\sqrt{s_{0}}\left\lVert h_{T_{0}}\right\rVert_{2}$ (139)
$\leq$ $\lambda_{n}\sqrt{s_{0}}K(s_{0},4)\left\lVert Xh\right\rVert_{n}$
$\leq$ $\left\lVert Xh\right\rVert^{2}_{n}+(\lambda_{n}\sqrt{s_{0}}K(s_{0},4)/2)^{2},$ (140)
from which (48) and (143) immediately follow since
[…]
$\left\lVert h_{T_{0}}\right\rVert_{2}$ $\leq$ $K(s_{0},4)\left\lVert Xh\right\rVert_{n}\leq\lambda_{n}\sqrt{s_{0}}K(s_{0},4)^{2}.$ (143)
Moreover, we have for $h\in{\mathcal{C}}(s_{0},4)$,
$\left\lVert h_{T_{0}}\right\rVert_{1}$ $\leq$ $\sqrt{s_{0}}\left\lVert h_{T_{0}}\right\rVert_{2}\leq\lambda_{n}s_{0}K(s_{0},4)^{2},$
$\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ $\leq$ $4\lambda_{n}K(s_{0},4)^{2}s_{0}=:D_{1,b}\lambda\sigma s_{0},\;\text{ and }$ (144)
$\left\lVert h\right\rVert_{1}$ $\leq$ ${5}\lambda_{n}K(s_{0},4)^{2}s_{0}={5d_{0}}K(s_{0},4)^{2}\lambda\sigma s_{0}=:D_{2,b}\lambda\sigma s_{0},$ (145)
where $D_{1,b}=4d_{0}K(s_{0},4)^{2}$ and $D_{2,b}=5d_{0}K(s_{0},4)^{2}$. Case 3. Suppose
$\lambda_{n}\left\lVert h_{T_{0}}\right\rVert_{1}\leq\left\lVert X\delta\right\rVert^{2}_{n}+\left\lVert Xh\right\rVert^{2}_{n}\leq\left\lVert X\beta-X\beta_{0}\right\rVert_{n}^{2}.$ (146)
Thus we have on ${\mathcal{T}}_{a}$, by the triangle inequality, (146), and (150),
$\lambda_{n}\left\lVert h_{T_{0}^{c}}\right\rVert_{1}+\lambda_{n}\left\lVert h_{T_{0}}\right\rVert_{1}$ $\leq$ $\left\lVert X\delta\right\rVert^{2}_{n}+\left\lVert Xh\right\rVert^{2}_{n}+\lambda_{n}\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$
$\leq$ $\left\lVert X\beta-X\beta_{0}\right\rVert_{n}^{2}+3\lambda_{n}\left\lVert h_{T_{0}}\right\rVert_{1}.$
Then
$\lambda_{n}\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ $\leq$ $\left\lVert X\beta-X\beta_{0}\right\rVert_{n}^{2}+2\lambda_{n}\left\lVert h_{T_{0}}\right\rVert_{1},\;\text{ and}$ (147)
$\lambda_{n}\left\lVert h\right\rVert_{1}$ $\leq$ $\left\lVert X\beta-X\beta_{0}\right\rVert_{n}^{2}+3\lambda_{n}\left\lVert h_{T_{0}}\right\rVert_{1}\leq 4\left\lVert X\beta-X\beta_{0}\right\rVert_{n}^{2}.$ (148)
Now by (148), we have for $d_{0}\geq 2\sqrt{1+a}$ and $h=\hat{\beta}-\beta_{0}$,
$\left\lVert h\right\rVert_{1}$ $\leq$ $\left\lVert X\beta-X\beta_{0}\right\rVert^{2}_{n}/(\lambda_{n})+3\left\lVert h_{T_{0}}\right\rVert_{1}\leq 4\left\lVert X\beta-X\beta_{0}\right\rVert_{n}^{2}/(\lambda_{n})$
$=$ $4\Lambda_{\max}(s-s_{0})\lambda\sigma s_{0}/(d_{0}).$
Putting things together. We will show in the proof of Lemma 3.3 that for for Case 3, the following $\ell_{2}$-norm error bound for $h=\hat{\beta}-\beta_{0}$,
[…]
Combining the preceding bound with (143) for Case 2 and (134) for Case 1, we have the expression for (26) and the following $\ell_{2}$ error bound on $h_{T_{0}}$},
$\left\lVert h_{T_{0}}\right\rVert_{2}$ $\leq$ $D_{0}^{\prime}\lambda\sigma\sqrt{s_{0}}\;\;\text{ where }$
$D_{0}^{\prime}$ $=$ $\left\{D\vee[d_{0}K(s_{0},4)^{2}]\vee[(K(s_{0},3)\sqrt{\Lambda_{\max}(s-s_{0})}+3d_{0}K^{2}(s_{0},3))]\right\}.$
In summary, for Case 1, we have the following bounds in the $\ell_{1}$ norm:
$\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ $\leq$ $\left\lVert X(\beta-\beta_{0})\right\rVert_{n}^{2}/(\lambda_{n})+(3K(s_{0},3)/2)^{2}\lambda_{n}s_{0},$
$\leq$ $[\frac{\Lambda_{\max}(s-s_{0})}{d^{2}_{0}}+9K(s_{0},3)^{2}/4]d_{0}\lambda\sigma s_{0}.$
and for Case 2, we have by (49),
$\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ $\leq$ $4d_{0}K(s_{0},4)^{2}\lambda\sigma s_{0}=:D_{1,b}\lambda\sigma s_{0};$
Finally, for Case 3, we have by (51),
$\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ $\leq$ $3\left\lVert X\beta-X\beta_{0}\right\rVert^{2}_{n}/\lambda_{n}\leq D_{1,c}\lambda\sigma s_{0}\;$
$\text{ where}\;\;D_{1,c}$ $:=$ $3\Lambda_{\max}(s-s_{0})/d_{0}.$
Moreover, combining (46), (44), (49), and (51), we have for $D_{1}$ as in (27):
$\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ $\leq$ $\lambda\sigma s_{0}\left\{\big(D_{1,a}\wedge 3D^{\prime}_{0}]\vee 4d_{0}K(s_{0},4)^{2}\vee\frac{3\Lambda_{\max}(s-s_{0})}{d_{0}}\right\}$
$\leq$ $d_{0}\lambda\sigma s_{0}\left\{4K(s_{0},4)^{2}+\frac{3\Lambda_{\max}(s-s_{0})}{d^{2}_{0}}\right\}$
Similarly, combining (135), (145) for Case 2, and (D.1) for Case 3, we obtain the expression of $D_{2}$ in (29):
$\left\lVert h\right\rVert_{1}$ $\leq$ $\lambda\sigma s_{0}\left\{\big(\frac{\Lambda_{\max}(s-s_{0})}{d_{0}}+4d_{0}K^{2}(s_{0},3)\big)\vee 5d_{0}K(s_{0},4)^{2}\vee\frac{4\Lambda_{\max}(s-s_{0})}{d_{0}}\right\}.$
Appendix E Proof of Lemma 3.3
Let $h=\beta_{\text{\rm init}}-\beta^{\mathop{\natural}\kern 0.60275pt}({T_{0}})$. Throughout this proof, as a shorthand, we write $h=\hat{\beta}-\beta_{0}$, where $\hat{\beta}=\beta_{\text{\rm init}}$ and $\beta_{0}=\beta^{\mathop{\natural}\kern 0.60275pt}({T_{0}})$ as defined in Lemma 2.3. Under the settings of Theorem 2.4, we have on ${\mathcal{T}}_{a}$,
$\left\lVert X\delta\right\rVert^{2}_{n}+\left\lVert Xh\right\rVert^{2}_{n}+\lambda_{n}\left\lVert h_{T_{0}^{c}}\right\rVert_{1}\leq\left\lVert X(\beta-\beta_{0})\right\rVert^{2}_{n}+3\lambda_{n}\left\lVert h_{T_{0}}\right\rVert_{1}.$ (150)
The rest of the proof is devoted to the $\ell_{2}$ error bound on $h_{T_{01}}$ in view of Lemma 3.1. Let $T_{1}$ be the $s_{0}$ largest positions of $h$ outside of $T_{0}$. Denote by $J_{0}$ the locations of the $s_{0}$ largest coefficients of $h$ in absolute values. Then $J_{0}\subset T_{01}$. Case 1 Now by Proposition A.1 as derived by [46], (150), (44), and Lemma D.1, we have for $K=K(s_{0},3)$,
$\left\lVert h_{T_{01}}\right\rVert_{2}$ $\leq$ $\sqrt{2}K(s_{0},3)\left\lVert Xh\right\rVert_{n}$
$\leq$ $\sqrt{2}K\big(\left\lVert X\beta-X\beta_{0}\right\rVert_{n}^{2}+3\lambda_{n}\sqrt{s_{0}}\left\lVert h_{T_{0}}\right\rVert_{2}\big)^{1/2}$
$\leq$ $\sqrt{2}K(s_{0},3)\big(\left\lVert X\beta-X\beta_{0}\right\rVert^{2}_{n}+3\lambda_{n}\sqrt{s_{0}}K\left\lVert X\beta-X\beta_{0}\right\rVert_{n}+9K^{2}\lambda^{2}_{n}s_{0})\big)^{1/2}$
$\leq$ $\sqrt{2}K(s_{0},3)\big(\big(\left\lVert X\beta-X\beta_{0}\right\rVert_{n}+3\lambda_{n}\sqrt{s_{0}}K\big)^{2}\big)^{1/2}$
$\leq$ $D_{0,a}\lambda\sigma\sqrt{s_{0}},$
where
$D_{0,a}:=\sqrt{2}\big(\sqrt{\Lambda_{\max}(s-s_{0})}K(s_{0},3)+3d_{0}K^{2}(s_{0},3)\big).$
By Lemma 3.1,
$\left\lVert h\right\rVert_{2}^{2}$ $\leq$ $\left\lVert h_{T_{01}}\right\rVert_{2}^{2}+\left\lVert h_{T_{0}^{c}}\right\rVert^{2}_{1}/s_{0}\leq[D_{0,a}^{2}+D_{1}^{2}]\lambda^{2}\sigma^{2}s_{0}$
$\text{ and }\;\left\lVert h\right\rVert_{2}^{2}$ $\leq$ $(1+k_{0})\left\lVert h_{T_{01}}\right\rVert_{2}^{2},\;\text{ where}\;h\in{\mathcal{C}}(s_{0},k_{0}),$
when $\textnormal{{RE}}(s_{0},k_{0},X)$ holds. To see this, notice that we have by Lemma 3.2, where we set $k_{0}=3$,
$\left\lVert h\right\rVert_{2}$ $\leq$ $\sqrt{(1+3)}\left\lVert h_{J_{0}}\right\rVert_{2}\leq 2\left\lVert h_{T_{01}}\right\rVert_{2}\leq 2D_{0,a}\lambda\sigma\sqrt{s_{0}}.$
Case 2. Similar to Case 1, we have by (142),
$\left\lVert h_{T_{01}}\right\rVert_{2}$ $\leq$ $\sqrt{2}K(s_{0},4)\left\lVert Xh\right\rVert_{n}\leq\sqrt{2}K^{2}(s_{0},4)\lambda_{n}\sqrt{s_{0}}.$
since $h\in{\mathcal{C}}(s_{0},4)$. Hence, we have by Lemma 3.2, where we set $k_{0}=4$,
$\left\lVert h\right\rVert_{2}$ $\leq$ $\sqrt{(1+4)}\left\lVert h_{J_{0}}\right\rVert_{2}\leq\sqrt{(1+4)}\left\lVert h_{T_{01}}\right\rVert_{2}$
$\leq$ $\sqrt{5}\sqrt{2}K^{2}(s_{0},4)\lambda_{n}\sqrt{s_{0}}=\sqrt{10}d_{0}K(s_{0},4)^{2}\lambda\sigma\sqrt{s_{0}}.$
Case 3. Denote by
$D$ $:=$ $\frac{\sqrt{\Lambda_{\max}(s-s_{0})}}{\sqrt{\Lambda_{\min}(2s_{0})}}\left(1+\frac{3\ell(s_{0})\sqrt{\Lambda_{\max}(s-s_{0})}}{d_{0}}\right).$ (151)
By Lemmas 3.1,  D.2D.1, Assumption (146), and (51), we obtain for $h=\hat{\beta}-\beta_{0}$,
$\left\lVert h_{T_{01}}\right\rVert_{2}$ $\leq$ $\frac{1}{\sqrt{\Lambda_{\min}(2s_{0})}}\big(\left\lVert Xh\right\rVert_{n}+\ell(s_{0})\left\lVert h_{T_{0}^{c}}\right\rVert_{1}/\sqrt{s_{0}}\big)$
$\leq$ $\frac{1}{\sqrt{\Lambda_{\min}(2s_{0})}}\left\lVert X\beta-X\beta_{0}\right\rVert_{{n}}+\frac{\ell(s_{0})}{\sqrt{\Lambda_{\min}(2s_{0})}}\frac{3\left\lVert X\beta-X\beta_{0}\right\rVert^{2}_{n}}{\lambda_{n}\sqrt{s_{0}}}$
$\leq$ $\frac{\left\lVert X\beta-X\beta_{0}\right\rVert_{n}}{\sqrt{\Lambda_{\min}(2s_{0})}}\big(1+\frac{3\left\lVert X\beta-X\beta_{0}\right\rVert_{n}\ell(s_{0})}{\lambda\sigma d_{0}\sqrt{s_{0}}}\big)$
$\leq$ $\sqrt{\frac{\Lambda_{\max}(s-s_{0})}{\Lambda_{\min}(2s_{0})}}\left(1+\frac{3\ell(s_{0})\sqrt{\Lambda_{\max}(s-s_{0})}}{d_{0}}\right)\lambda\sigma\sqrt{s_{0}}=:D\lambda\sigma\sqrt{s_{0}}$
where recall for Case 3,
$\left\lVert h_{T_{0}^{c}}\right\rVert_{1}/\sqrt{s_{0}}$ $\leq$ $3\left\lVert X\beta-X\beta_{0}\right\rVert_{n}^{2}/(\lambda_{n}\sqrt{s_{0}})\leq D_{1,c}\lambda\sigma\sqrt{s_{0}},$
$=$ $3\Lambda_{\max}(s-s_{0})\lambda\sigma\sqrt{s_{0}}/d_{0}$
$\text{ and }\;\left\lVert Xh\right\rVert_{n}$ $\leq$ $\left\lVert X(\beta-\beta_{0})\right\rVert_{n}\leq\sqrt{\Lambda_{\max}(s-s_{0})}\lambda\sigma\sqrt{s_{0}}.$
Combining all three cases, we obtain the expression for (23):
$\left\lVert h_{T_{01}}\right\rVert_{2}$ $\leq$ $D_{0}\lambda\sigma\sqrt{s_{0}}\text{ where}\;\;D_{0}:=D_{0,a}\vee[d_{0}\sqrt{2}K^{2}(s_{0},4)]\vee D$
where $D$ is as defined in (151) and $D_{0,a}$ as in (52).
It remains to bound $\left\lVert h\right\rVert_{2}$ for Case 3. Thus we have on ${\mathcal{T}}_{a}$ for $\lambda_{n}=d_{0}\lambda\sigma$ and $\ell(s_{0})$ as defined in (D.2), for $D$ as in (151), by Lemma 3.1,
$\left\lVert h\right\rVert_{2}^{2}$ $\leq$ $\left\lVert h_{T_{01}}\right\rVert_{2}^{2}+\left\lVert h_{T_{0}^{c}}\right\rVert^{2}_{1}/s_{0}$
$\leq$ $\big(\frac{1}{\sqrt{\Lambda_{\min}(2s_{0})}}\big(\left\lVert Xh\right\rVert_{n}+\ell(s_{0})\left\lVert h_{T_{0}^{c}}\right\rVert_{1}/\sqrt{s_{0}}\big)\big)^{2}+\left\lVert h_{T_{0}^{c}}\right\rVert^{2}_{1}/s_{0}$
$\leq$ $(D^{2}+D_{1,c}^{2})\lambda^{2}\sigma^{2}s_{0},\;\text{where}\;D_{1,c}=\frac{3\Lambda_{\max}(s-s_{0})}{d_{0}},$
Finally, for $h=\hat{\beta}-\beta_{0}$, where $\beta_{0}=\beta^{\mathop{\natural}\kern 0.60275pt}({T_{0}})$, $D_{0}$ as defined in (23), and $D_{1}$ as defined in (27), we have by the triangle inequality and Lemma 3.1,
$\left\lVert\hat{\beta}-\beta\right\rVert_{2}$ $\leq$ $\left\lVert\hat{\beta}-\beta_{0}\right\rVert_{2}+\left\lVert\beta-\beta_{0}\right\rVert_{2}\leq\left\lVert h\right\rVert_{2}+\lambda\sigma\sqrt{s_{0}}$
$\leq$ $(\left\lVert h_{T_{01}}\right\rVert_{2}^{2}+\left\lVert h_{T_{0}^{c}}\right\rVert^{2}_{1}/s_{0})^{1/2}+\lambda\sigma\sqrt{s_{0}}$
$\leq$ $[\sqrt{D_{0}^{2}+D_{1}^{2}}+1]\lambda\sigma\sqrt{s_{0}},$
over all cases. Moreover, we have by Lemmas 3.1 and 3.2, for both Case 1 and Case 2, the following slightly stronger bound:
$\left\lVert\hat{\beta}-\beta\right\rVert_{2}$ $\leq$ $\left\lVert h\right\rVert_{2}+\lambda\sigma\sqrt{s_{0}}$
$\leq$ $(\left\lVert h_{T_{01}}\right\rVert_{2}^{2}+\left\lVert h_{T_{0}^{c}}\right\rVert^{2}_{1}/s_{0})^{1/2}\wedge\sqrt{1+k_{0}}\left\lVert h_{T_{01}}\right\rVert_{2}+\lambda\sigma\sqrt{s_{0}}$
$\leq$ $\lambda\sigma\sqrt{s_{0}}[(\sqrt{D_{0}^{2}+D_{1}^{2}}+1)\wedge(\sqrt{5}D_{0}+1)].$
Appendix F Proof of Lemma D.2
For $Xh_{T_{01}}=Xh-\sum_{j\geq 2}Xh_{T_{j}}$, we have by the triangle inequality and sparse eigenvalue condition,
$\left\lVert Xh_{T_{01}}\right\rVert_{2}/\sqrt{n}$ $\leq$ $\left\lVert Xh\right\rVert_{2}/\sqrt{n}+\sum_{j\geq 2}\left\lVert Xh_{T_{j}}\right\rVert_{2}/\sqrt{n}$
$\leq$ $\left\lVert Xh\right\rVert_{2}/\sqrt{n}+\sqrt{\Lambda_{\max}(s_{0})}\sum_{j\geq 2}\left\lVert h_{T_{j}}\right\rVert_{2}$
$\leq$ $\left\lVert Xh\right\rVert_{2}/\sqrt{n}+\sqrt{\Lambda_{\max}(s_{0})}\sum_{j\geq 1}\left\lVert h_{T_{j}}\right\rVert_{1}/{\sqrt{s_{0}}}$
$\leq$ $\left\lVert Xh\right\rVert_{2}/\sqrt{n}+\sqrt{\Lambda_{\max}(s_{0})}\left\lVert h_{T_{0}^{c}}\right\rVert_{1}/{\sqrt{s_{0}}}.$
Thus it follows from the proof Lemma 3.1 that
$\left\lVert h_{T_{01}}\right\rVert_{2}$ $\leq$ $\left\lVert Xh_{T_{01}}\right\rVert_{2}/\sqrt{n\Lambda_{\min}(2s_{0})}$ (152)
$\leq$ $\frac{1}{\sqrt{\Lambda_{\min}(2s_{0})}}\left(\left\lVert Xh\right\rVert_{2}/\sqrt{n}+\sqrt{\Lambda_{\max}(s_{0})}\left\lVert h_{T_{0}^{c}}\right\rVert_{1}/{\sqrt{s_{0}}}\right),$
where we replace $\frac{\theta_{s_{0},2s_{0}}}{\sqrt{\Lambda_{\min}(2s_{0})}}$ with $\sqrt{\Lambda_{\max}(s_{0})}$. The lemma thus holds in view of (38) and (152). ∎
Appendix G Nearly ideal model selections under the UUP
The Dantzig selector [9] is defined as follows: for some $\lambda_{n}\geq 0$,
$(DS)\;\;\arg\min_{\tilde{\beta}\in\mathbb{R}^{p}}\left\lVert\tilde{\beta}\right\rVert_{1}\;\;\text{subject to}\;\;\left\lVert X^{T}(Y-X\tilde{\beta})/n\right\rVert_{\infty}\leq\lambda_{n}.$ (153)
[47] shows that thresholding of an initial Dantzig selector $\beta_{\text{\rm init}}$ at the level of $\asymp\sqrt{2\log p/n}\sigma$ followed by OLS refitting, achieves the sparse oracle inequalities under a UUP.
section G.1.
(A Uniform Uncertainly Principle) For some integer $1\leq s<n/3$, assume $\delta_{2s}+\theta_{s,2s}<1-\tau$ for some $\tau>0$, which implies that $\Lambda_{\min}(2s)>\theta_{s,2s}$.
Remark G.2.
It is clear that $\delta_{2s}<1$ implies that the sparse eigenvalues condition (6) and (5) hold. Moreover, Assumption G.1 implies that $\textnormal{{RE}}(s_{0},k_{0},X)$ as in (3) hold for $s_{0}\leq s$ with
$K(s_{0},k_{0})\leq K(s,k_{0})=\frac{\sqrt{\Lambda_{\min}(2s)}}{\Lambda_{\min}(2s)-\theta_{s,2s}}\leq\frac{\sqrt{\Lambda_{\min}(2s)}}{1-\delta_{2s}-\theta_{s,2s}}\leq\frac{\sqrt{\Lambda_{\min}(2s)}}{\tau}$ (154)
for $k_{0}=1$, as $K(s,k_{0})$ is nondecreasing with respect to $s$ for the same $k_{0}$; see [3].
The Gauss-Dantzig Procedure: Assume $\delta_{2s}+\theta_{s,2s}<1-\tau$, where $\tau>0$:
  • Step 1
    First obtain an initial estimator $\beta_{\text{\rm init}}$ using the Dantzig selector in (153) with $\lambda_{n}=(\sqrt{1+a}+\tau^{-1})\sqrt{2\log p/n}\sigma=:\lambda_{p,\tau}\sigma$, where $a\geq 0$; then threshold $\beta_{\text{\rm init}}$ with $t_{0}$, chosen from the range $(C_{1}\lambda_{p,\tau}\sigma,C_{4}\lambda_{p,\tau}\sigma]$, for $C_{1}$ as defined in (157); set $I:=\left\{j\in\{1,\ldots,p\}:\beta_{j,\text{\rm init}}\geq t_{0}\right\}.$
  • Step 2
    Given a set $I$ as above, construct the estimator $\hat{\beta}_{I}=(X_{I}^{T}X_{I})^{-1}X_{I}^{T}Y$ and set $\hat{\beta}_{j}=0,\forall j\not\in I.$
Theorem G.3.
[…]
$|I|$ $\leq$ $2s_{0},\quad|I\setminus S|\leq s_{0}\leq s\quad\text{ and }\quad\lVert\hat{\beta}-\beta\rVert_{2}^{2}\leq C_{3}^{2}\lambda^{2}\sigma^{2}s_{0}$ (155)
with probability at least $1-(\sqrt{\pi\log p}p^{a})^{-1}$, where $C_{1}$ is defined in (157) and $C_{3}$ depends on $a,\tau$, $\delta_{2s}$, $\theta_{s,2s}$ and $C_{4}$; see (158).
Our analysis for Theorem G.3 builds upon Proposition G.4 [9], which shows the Dantzig selector achieves the oracle inequality as stated in (11) under Assumption G.1. We note that, in Assumption G.1, the sparsity level is fixed at $s$ rather than $s_{0}$. Hence it is stronger than the conditions we impose in Theorem 2.1 for the Thresholded Lasso estimator. We now show the oracle inequalities for the Dantzig selector. We then show in the supplementary Lemma H.1 that thresholding at the level of $\sigma\lambda$ as elaborated in Step 1 in the Gauss-Dantzig Procedure selects a set $I$ of at most $2s_{0}$ variables, among which at most $s_{0}$ are from the complement of the support set $S$ as required in (155).
Proposition G.4.
[9] Let $Y=X\beta+\epsilon$, for $\epsilon$ being i.i.d. $N(0,\sigma^{2})$ and $\left\lVert X_{j}\right\rVert_{2}^{2}=n$. Choose $\tau,a>0$ and set $\lambda_{n}=(\sqrt{1+a}+\tau^{-1})\sigma\sqrt{2\log p/n}$ in (153). Then if $\beta$ is $s$-sparse with $\delta_{2s}+\theta_{s,2s}<1-\tau$, the Dantzig selector obeys with probability at least $1-(\sqrt{\pi\log p}p^{a})^{-1}$,
$\lVert\hat{\beta}-\beta\rVert_{2}^{2}$ $\leq$ $C_{2}^{2}(\sqrt{1+a}+\tau^{-1})^{2}s_{0}\lambda^{2}\sigma^{2}$
From this point on we let $\delta:=\delta_{2s}$ and $\theta:=\theta_{s,2s}$; Analysis by [9] (Theorem 2) and the current paper yields the following constants,
$C_{2}$ $=$ $2C_{0}^{\prime}+\frac{1+\delta}{1-\delta-\theta}\;\text{ where }C^{\prime}_{0}=\frac{C_{0}}{1-\delta-\theta}+\frac{\theta(1+\delta)}{(1-\delta-\theta)^{2}},$ (156)
where $C_{0}=2\sqrt{2}\left(1+\frac{1-\delta^{2}}{1-\delta-\theta}\right)+(1+1/\sqrt{2})\frac{(1+\delta)^{2}}{1-\delta-\theta}$. We now define
$C_{1}$ $=$ $C_{0}^{\prime}+\frac{1+\delta}{1-\delta-\theta}\text{ and }$ (157)
$C_{3}^{2}$ $=$ $3(\sqrt{1+a}+\tau^{-1})^{2}((C_{0}^{\prime}+C_{4})^{2}+1)+{4(1+a)}/{\Lambda^{2}_{\min}(2s_{0})},$ (158)
where $C_{3}$ is used in (155) and has not been optimized in our analysis.
Appendix H Proof of Theorem G.3
Now similar to Lemma 2.3, Lemma H.1 bounds the size of $I$, as well as the bias that we introduce to model $I$ by thresholding. Theorem G.3 is an immediate corollary of Lemmas 2.5 and H.1. The proof follows from Lemma H.1 and Proposition G.4. We include its proof in Section H.1 for self-containment.
Lemma H.1.
Choose $\tau>0$ such that $\delta_{2s}+\theta_{s,2s}<1-\tau$. Let $\beta_{\text{\rm init}}$ be the solution to (153) with $\lambda_{n}=\lambda_{p,\tau}\sigma:=(\sqrt{1+a}+\tau^{-1})\sqrt{2\log p/n}\sigma$. Given some constant $C_{4}\geq C_{1}$, for $C_{1}$ as in (157), choose a thresholding parameter $t_{0}$ such that $C_{4}\lambda_{p,\tau}\sigma\geq t_{0}>C_{1}\lambda_{p,\tau}\sigma$ and set $I=\{j:\left|\beta_{j,\text{\rm init}}\right|\geq t_{0}\}.$ Then with probability at least $1-(\sqrt{\pi\log p}p^{a})^{-1}$, we have (155) and $\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}\leq\sqrt{(C_{0}^{\prime}+C_{4})^{2}+1}\lambda_{p,\tau}\sigma\sqrt{s_{0}}$, where ${\mathcal{D}}:=\{1,\ldots,p\}\setminus I$ and $C_{0}^{\prime}$ is defined in (156).
Proof of Theorem G.3. It holds by definition of $S_{{\mathcal{D}}}$ that $I\cap S_{{\mathcal{D}}}=\emptyset$. It is clear by Lemma H.1 that $|{S_{{\mathcal{D}}}}|<s$ and $|I|\leq 2s_{0}$ and $|I\cup S_{{\mathcal{D}}}|\leq|I\cup S|\leq s+s_{0}\leq 2s$; Thus for $\hat{\beta}_{I}=(X_{I}^{T}X_{I})^{-1}X_{I}^{T}Y$, $\hat{\beta}_{I^{c}}=0$ and $\lambda=\sqrt{2\log p/n}$, we have by Lemma 2.7
$\left\lVert\hat{\beta}-\beta\right\rVert_{2}^{2}\leq\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}^{2}\left(1+\frac{2\theta^{2}_{s,2s_{0}}}{\Lambda_{\min}^{2}(2s_{0})}\right)+\frac{4s_{0}}{\Lambda_{\min}^{2}(2s_{0})}\lambda_{\sigma,a,p}^{2}$
$\leq$ $\lambda^{2}\sigma^{2}s_{0}\left(\sqrt{1+a}+\tau^{-1})^{2}((C_{0}^{\prime}+C_{4})^{2}+1)\big(1+\frac{2\theta^{2}_{s,2s_{0}}}{\Lambda_{\min}^{2}(2s_{0})}\big)+\frac{4(1+a)}{\Lambda^{2}_{\min}(2s_{0})}\right)$
$\leq$ $C_{3}^{2}\lambda^{2}\sigma^{2}s_{0}$
with probability at least $1-(\sqrt{\pi\log p}p^{a})^{-1}-\exp(-3m/64)$ where $m=\left|I\right|$. Thus the theorem holds for $C_{3}$ as in (158), where it holds for $\tau>0$ that
$\frac{\theta_{s,2s_{0}}}{\Lambda_{\min}(2s_{0})}\leq\frac{\theta_{s,2s}}{\Lambda_{\min}(2s_{0})}\leq\frac{1-\delta_{2s}-\tau}{\Lambda_{\min}(2s)}<1$
given that $\theta_{s,2s}<1-\tau-\delta_{2s}<\Lambda_{\min}(2s)$ for $\tau>0$. ∎
H.1 Proof of Lemma H.1
Suppose ${\mathcal{T}}_{a}$ holds. Consider the set $I\cap T_{0}^{c}:=\{j\in T_{0}^{c}:\left|\beta_{j,\text{\rm init}}\right|>t_{0}\}$. It is clear by definition of $h=\beta_{\text{\rm init}}-\beta^{(1)}$ and (161) that
$\left|I\cap T_{0}^{c}\right|\leq\left\lVert\beta_{T_{0}^{c},\text{\rm init}}\right\rVert_{1}/t_{0}=\left\lVert h_{T_{0}^{c}}\right\rVert_{1}/t_{0}<s_{0},$ (159)
where $t_{0}\geq C_{1}\lambda_{p,\tau}\sigma$. Thus $|I|=|I\cap T_{0}|+|I\cap T_{0}^{c}|\leq 2s_{0}$; Now (155) holds given (159) and $|I\cup S|=|S|+|I\cap S^{c}|\leq s+|I\cap T_{0}^{c}|<s+s_{0}.$ We now bound $\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}^{2}$. By (160) and (56), where ${\mathcal{D}}_{11}\subset T_{0}$, we have for $\tau<C_{4}\lambda_{p,\tau}\sigma$, by the triangle inequality,
$\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}^{2}$ $=$ $\left\lVert\beta_{{\mathcal{D}}\cap T_{0}^{c}}\right\rVert_{2}^{2}+\left\lVert\beta_{{\mathcal{D}}\cap T_{0}}\right\rVert_{2}^{2}\leq\left\lVert\beta^{(2)}\right\rVert_{2}^{2}+\left\lVert\beta_{I^{c}\cap T_{0}}\right\rVert_{2}^{2}$
$\leq$ $s_{0}\lambda^{2}\sigma^{2}+(t_{0}\sqrt{s_{0}}+\left\lVert h_{T_{0}}\right\rVert_{2})^{2}\leq((C_{4}+C_{0}^{\prime})^{2}+1)\lambda_{p,\tau}^{2}\sigma^{2}s_{0}.$
The proof of Proposition G.4 [9] yields the following on ${\mathcal{T}}_{a}$,
[…]
$\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ $\leq$ $C_{1}\lambda_{p,\tau}\sigma s_{0},\;\text{ where }\;C_{1}=\left(C^{\prime}_{0}+\frac{1+\delta}{1-\delta-\theta}\right),\;\text{ and }$ (161)
[…]
The rest of the proof follows that of Lemma 2.3 and hence omitted. ∎
(a) (b)
Figure 5: Illustrative example: i.i.d. Gaussian ensemble; $p=256$, $n=72$, $s=8$, and $\sigma=\sqrt{s}/3$. (a) compare with the Lasso estimator $\tilde{\beta}$ which minimizes $\ell_{2}$ loss. Here $\tilde{\beta}$ has only 3 FPs, but $\rho^{2}$ is large with a value of $64.73$. (b) Compare with the $\beta_{\text{\rm init}}$ obtained using $\lambda_{n}$. The dotted lines show the thresholding level $t_{0}$. The $\beta_{\text{\rm init}}$ has 15 FPs, all of which were cut after the 2nd step; resulting $\rho^{2}=12.73$. After refitting with OLS in the 3rd step, for the $\hat{\beta}$, $\rho^{2}$ is further reduced to $0.51$.
Appendix I Numerical experiments
In this section, we present additional results from numerical simulations designed to validate the theoretical analysis presented in this paper.
The experiment set up here is the same as the one in Section  5.1. The main difference is that all non-zero entries in $\beta$ have large magnitudes around 1, in particular, they follow this distribution $\beta_{i}=\mu_{i}(1+|g_{i}|),$ where $\mu_{i}=\pm 1$ with probability 1/2 and $g_{i}\sim N(0,1)$. We use $\lambda_{n}=0.69\lambda\sigma$ throughout the experiments in this section to select a $\beta_{\text{\rm init}}$ as the initial estimator. We then threshold the $\beta_{\text{\rm init}}$ using a value $t_{0}$ typically chosen between $0.5\lambda\sigma$ and $\lambda\sigma$. See each experiment for the actual value used. Given that columns of $X$ being normalized to have $\ell_{2}$ norm $\sqrt{n}$, for each input $\beta$, we compute its SNR as follows: $SNR:=\left\lVert\beta\right\rVert_{2}^{2}/\sigma^{2}.$ To evaluate $\hat{\beta}$, we use metrics defined in Table 3; we also compute the ratio between squared $\ell_{2}$ error and the ideal mean squared error, known as the $\rho^{2}$; see Section I.2 for details.
I.1 Type I/II errors
We now evaluate the Thresholded Lasso estimator by comparing Type I/II errors under different values of $t_{0}$ and SNR. We consider Gaussian random matrices for the design $X$ with both diagonal and Toeplitz covariance. We refer to the former as i.i.d. Gaussian ensemble and the latter as Toeplitz ensemble. In the Toeplitz case, the covariance is given by $T(\gamma)_{i,j}=\gamma^{|i-j|}$ where $0<\gamma<1.$ We run under two noise levels: $\sigma=\sqrt{s}/3$ and $\sigma=\sqrt{s}$. For each $\sigma$, we vary the threshold $t_{0}$ from $0.01\lambda\sigma$ to $1.5\lambda\sigma$. For each $\sigma$ and $t_{0}$ combination, we run the experiment as described in Section 5.1 $200$ times with a new $\beta$ and $\epsilon$ generated each time and we count the number of Type I and II errors in $\hat{\beta}$. We compute the average at the end of 200 runs, which will correspond to one data point on the curves in Figure 6 (a) and (b).
For both types of designs, similar behaviors are observed. For $\sigma=\sqrt{s}/3$, FNs increase slowly; hence there is a wide range of values from which $t_{0}$ can be chosen such that FNs and FPs are both zero. In contrast, when $\sigma=\sqrt{s}$, FNs increase rather quickly as $t_{0}$ increases due to the low SNR. It is clear that the low SNR and high correlation combination makes it the most challenging situation for variable selection, as predicted by our theoretical analysis and others. In (c) and (d), we run additional experiments for the low SNR case for Toeplitz ensembles. The performance is improved by increasing the sample size or lowering the correlation factor.
Table 3: Metrics for evaluating $\hat{\beta}$
Metric Definition
Type I errors or False Positives (FPs) # of incorrectly selected non-zeros in $\hat{\beta}$
Type II errors or False Negatives (FNs) # of non-zeros in $\beta$ that are not selected in $\hat{\beta}$
True positives (TPs) # of correctly selected non-zeros
True Negatives (TNs) # of zeros in $\hat{\beta}$ that are also zero in $\beta$
False Positive Rate (FPR) FPR=FP/(FP+TN)=FP/(ps)FPR=FP/(FP+TN)=FP/(p-s)
True Positive Rate (TPR) $TPR=TP/(TP+FN)=TP/s$
I.2 $\ell_{2}$ loss
We now compare the performance of the Thresholded Lasso with the ordinary Lasso by examining the metric $\rho^{2}$ defined as follows: $\rho^{2}=\frac{\sum_{i=1}^{p}(\hat{\beta}_{i}-\beta_{i})^{2}}{\sum_{i=1}^{p}\min(\beta_{i}^{2},\sigma^{2}/n)}.$
(a) (b)
(c) (d)
(e) (f)
Figure 6: $p=256$ $s=8$. (a) (b) Type I/II errors for i.i.d. Gaussian and Toeplitz ensembles. Each vertical bar represents $\pm 1$ std. The unit of $x$-axis is in $\lambda\sigma$. For both types of design matrices, FPs decrease and FNs increase as the threshold increases. For Toeplitz ensembles, in (c) with fixed correlation $\gamma$, FNs decrease with more samples, and in (d) with fixed sample size, FNs decrease as the correlation $\gamma$ decreases. (e) (f) Histograms of $\rho^{2}$ under i.i.d Gaussian ensembles from 500 runs.
We first run the above experiment using i.i.d. Gaussian ensemble under the following thresholds: $t_{0}=\lambda\sigma$ for $\sigma=\sqrt{s}/3$, and $t_{0}=0.36\lambda\sigma$ for $\sigma=\sqrt{s}$. These are chosen based on the desire to have low errors of both types (as shown in Figure 6 (a)). Naturally, for low SNR cases, small $t_{0}$ will reduce Type II errors. In practice, we suggest using cross-validations to choose the exact constants in front of $\lambda\sigma$; See, for example, a subsequent work [50] for details. We plot the histograms of $\rho^{2}$ in Figure 6 (e) and (f). In (e), the mean and median are $1.45$ and $1.01$ for the Thresholded Lasso, and $46.97$ and $41.12$ for the Lasso. In (f), the corresponding values are $7.26$ and $6.60$ for the Thresholded Lasso and $10.50$ and $10.01$ for the Lasso. With high SNR, the Thresholded Lasso performs extremely well; with low SNR, the improvement of the Thresholded Lasso over the ordinary Lasso is less prominent; this is in close correspondence with the Gauss-Dantzig selector’s behavior as shown by [9].
Next we run the above experiment under different sparsity values of $s$. We again use i.i.d. Gaussian ensemble with $p=2000$, $n=400$, and $\sigma=\sqrt{s}/3$. The threshold is set at $t_{0}=\lambda\sigma$. The SNR for different $s$ is fixed at around $32.36$. Table 4 shows the mean of the $\rho^{2}$ for the Lasso and the Thresholded Lasso estimators. The Thresholded Lasso performs consistently better than the ordinary Lasso until about $s=80$, after which both break down. For the Lasso, we always choose from the full regularization path the optimal $\tilde{\beta}$ that has the minimum $\ell_{2}$ loss.
Table 4: $\rho^{2}$ under different sparsity and fixed SNR. Average over 100 runs for each $s$.
s 5 18 20 40 60 80 100
SNR 34.66 32.99 32.29 32.08 32.28 32.56 32.54
Lasso 17.42 22.01 44.89 52.68 31.88 29.40 47.63
Thresholded Lasso 1.02 0.96 1.11 1.54 10.32 29.38 53.81
I.3 Linear Sparsity
We next present results demonstrating that the Thresholded Lasso recovers a sparse model using a small number of samples per non-zero component in $\beta$ when $X$ is a subgaussian ensemble. We run under three cases of $p=256,512,1024$; for each $p$, we increase the sparsity $s$ by roughly equal steps from $s={0.2p}/{\log(0.2p)}$ to $p/4$. For each $p$ and $s$, we run with different sample size $n$. For each tuple $(n,p,s)$, we run an experiment similar to the one described in Section I.1 with an i.i.d. Gaussian ensemble $X$ being fixed while repeating Steps 232-3 100 times. In Step 2, each randomly selected non-zero coordinate of $\beta$ is assigned a value of $\pm 0.9$ with probability $1/2$. After each run, we compare $\hat{\beta}$ with the true $\beta$; if all components match in signs, we count this experiment as a success. At the end of the 100 runs, we compute the percentage of successful runs as the probability of success. We compare with the ordinary Lasso, for which we search over the full regularization path of LARS and choose the $\breve{\beta}$ that best matches $\beta$ in terms of support.
We experiment with $\sigma=1$ and $\sigma=\sqrt{s}/3$. The results are shown in Figure 7. We observe that under both noise levels, the Thresholded Lasso estimator requires much fewer samples than the ordinary lasso in order to conduct exact recovery of the sparsity pattern of the true linear model when all non-zero components are sufficiently large. When $\sigma$ is fixed as $s$ increases, the SNR is increasing; the experimental results illustrate the behavior of sparse recovery when it is close to the noiseless setting. Given the same sparsity, more samples are required for the low SNR case to reach the same level of success rate. Similar behavior was also observed for Toeplitz and Bernoulli ensembles with i.i.d. $\pm 1$ entries.
(a) (b)
(c) (d)
(e) (f)
Figure 7: (a) (b) Compare the probability of success for $p=256$ and $p=512$ under two noise levels. The Thresholded Lasso estimator requires much fewer samples than the ordinary Lasso. (c) (d) (e) show the probability of success of the Thresholded Lasso under different levels of sparsity and noise levels when $n$ increases for $p=512$ and $1024$. (f) The number of samples $n$ increases almost linearly with $s$ for p = 1024. More samples are required to achieve the same level of success when $\sigma=\sqrt{s}/3$ due to the relatively low SNR.
I.4 ROC comparison
We now compare the performance of the Thresholded Lasso estimator with the Lasso and the Adaptive Lasso by examining their ROC curves. Our parameters are $p=512$, $n=330$, $s=64$ and we run under two cases: $\sigma=\sqrt{s}/3$ and $\sigma=\sqrt{s}$. In the Thresholded Lasso, we vary the threshold level from $0.01\lambda\sigma$ to $1.5\lambda\sigma$. For each threshold, we run the experiment described in Section I.1 with an i.i.d. Gaussian ensemble $X$ being fixed while repeating Steps 232-3 100 times. After each run, we compute the FPR and TPR of the $\hat{\beta}$, and compute their averages after 100 runs as the FPR and TPR for this threshold. For the Lasso, we compute the FPR and TPR for each output vector along its entire regularization path. For the Adaptive Lasso, we use the optimal output $\tilde{\beta}$ in terms of $\ell_{2}$ loss from the initial Lasso penalization path as the input to its second step, that is, we set $\beta_{\text{\rm init}}:=\tilde{\beta}$ and use $w_{j}=1/\beta_{\text{\rm init},j}$ to compute the weights for penalizing those non-zero components in $\beta_{\text{\rm init}}$ in the second step, while all zero components of $\beta_{\text{\rm init}}$ are now removed. We then compute the FPR and TPR for each vector that we obtain from the second step’s LARS output. We implement the algorithms as given in  [52], the details of which are omitted here as its implementation has become standard. The ROC curves are plotted in Figure 8. The Thresholded Lasso performs better than both the ordinary Lasso and the Adaptive Lasso; its advantage is more apparent when the SNR is high.
(a) (b)
Figure 8: $p=512$ $n=330$ $s=64$. ROC for the Thresholded Lasso, ordinary Lasso and Adaptive Lasso. The Thresholded Lasso clearly outperforms the ordinary Lasso and the Adaptive Lasso for both high and low SNRs.